A Complete Bibliography of Publications in ACM SIGSOFT Software Engineering Notes: 2100–2109

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

26 December 2021 Version 1.04

Title word cross-reference

$105 [Rog10]. $39.99 [Saf10].

* [NS10b].

/ACM [TODM19, XPP19].

0-470-24211-6 [Tri10a]. 0-672-32877-1 [Saf10].

10th [RGBR14, MRJD17, SNGM19]. 12 [KS10]. 14th [CTD19, CEH19]. 1999 [Ber10a]. 1st [GKK^{+}19, LNG^{+}13, BCDE18a, BCDE18b, FS11, GGR10, GKMM18, GKS17, HMS16, KMR^{+}19, KKPJ10, KMTD17, RFD^{+}18b, RFD^{+}18a, TLG^{+}16].

2 [Sch16a]. 2009 [MMM10]. 2010 [CCM^{+}10, HRZN10, LAK10]. 2011 [FS11, OKNB11]. 2011/2012 [XZM13]. 2012 [GR12, HDDS12, LKM^{+}13, LIL13]. 2013 [Kat13, LNG^{+}13]. 2014 [REN^{+}14]. 2015 [EPBR16, Fra16]. 2016 [KOPR16]. 2017 [FM18c, FM18a, FM18d, FM18b, GKS17]. 2019 [SNGM19, SAHC19, TODM19, XPP19]. 21st [WC10]. 2nd [GR12, Gve13b, TSvD^{+}11, CHMW19, GK12, GKL18a, GKL18b, HDDS12, KKPJ12, LSM^{+}10, LMS11, OKNB11].

3 [Sch18, M.13]. 300 [BM10c]. 368pp. [Tri10a]. 3rd [Gve13c, Teo11, Tri10b, HKPS12].

404 [Sch16b]. 41st [SNGM19]. 4th [LRS11, BBU^{+}17, DJB17, GPW17, KNOF13, LLM^{+}12, LLM^{+}13].

510 [BM10a]. 580 [BM10b]. 5th [GKK^{+}19, CBK10, LAK10, UYG^{+}19].

6th [LIL13, LZK^{+}18].

7th [BBG^{+}13].

8 [BKP16].

978-0-387-21507-5 [BM10b]. 978-0-470-40129-3 [Sau10]. 978-0-521-81513-0 [Rog10]. 978-0-521-87546-2 [BM10c]. 978-0-521-88068-8 [Tri10b]. 978-0470-29455-0 [BM10a]. 9th [Gve13a, CD17].

Abbott [Gou12]. ABC [MBN13]. Abrial [Rus11]. Abstract [KPP12, JWB^{+}18, SSK13]. Abstraction [Wei18]. Abstracts [LMS11]. academic [Sch12d]. Academy [Sol19]. access [HDKB13, Jai11, Jai12]. Aceto [BM10c]. achieving [Gla12]. ACM [CEH19, KBRS17a, Not10, TODM19, XPP19, Del12a]. ACO [SS11]. Acquisition [HJ16, AJP13]. Across [ZS14, JS12]. Activities [KL18]. activity [CB12, GPC12, JDV12]. activity-based [JDV12]. ad [KSR12, JS12]. Ada [Aus11]. Adams [Sau13a]. adaptation [KB12]. Adaptive [HSS^{+}16, LPP^{+}19, MRJD17, HAJW13, RRSV13]. addressing [GD10]. Adequacy [PSJ18, BKMJ12]. Adi [Sch12a]. Adi-Tabatabai [Sch12a]. adjunct [Sch12c]. ADLs [Pan10]. Adriaan [Ngo12, Teo13c]. Advanced [Teo13a]. Advances [HRZN10]. advertisement [JS12]. Affairs [MWR19a, MWR19b]. Agenda [GKK^{+}19]. Agent [MWR19a, MWR19b, GSB11, GS12]. agents [MKK12b, MKK13b]. Aggregation [RR11]. Agile [APNT16, DK16, FCT^{+}17, Fra11, FM18c, FM18a, GR12, HMS16, Kay11, Miy11, Nie12, SM16, CSG13, DM13, HW13, Jan12, MM11a, MO11, Mor13, Jie16, Ban12b]. agility [WJ12, Whi11]. Ahead [MWR19a, MWR19b]. AHP [BA13]. AI [HWA12]. Ajaykumar [Wer10]. Akrivopoulou [Kie12]. al [BM10c]. Alfred [BM10b]. algebra [LGMM10]. algebraic [Rip10]. Algorithm [SNR17, Ban12a, GC12, JD13, MKP12, MM10a, MNB13, MT13, Men13, RFS10, RRN13, SAM13b]. Ali [Sch12a, HMB18]. Ali-Reza [Sch12a]. Aliasing [Wei18]. Aligning [MS19]. Alistair [Ber10a]. allocation [Ban11]. Alloy [DR18]. Alon [Jah12]. alternative [Ban11]. Amber [Cha13b]. among [Yu11]. Analysis [BS17, Bul18, GB13a, HAM^{+}19, KS11c, KBRS17b, LPP^{+}19, Lee18, SM16, BK11, CN11, CMGV13, DC13, Gre12b, HDKB13, KK12b, KS13b, Lan11, MG12, RT10, SBB12, TJ12, VS11b, dSAVP10]. Analytical [GS12, KSR12]. Analytics [Noo18, Ban12b]. analyze [PASS13]. analyzer [MBN13]. Analyzing [NUK13, SS11]. Andrea [Gve13b]. Andrew [Teo13h]. Android [MMP^{+}12, SGS12a, vdMvdMV12]. Anhai [Jah12]. Ann [Sch13a]. anonymous [VS11b]. Ant [BDJ10, SKS10]. Anton [Sau13b]. anywhere [Dek10]. API [JY12, Gla11]. app [Ngo12]. applicability [SS13]. Application [BK16, BL10, Ban12a, BB11b, KB11a, MKS10, SK10b, AJP13, Ban10a, Ban11, CSKB11, Del13, Gre12b, Jai12, PA10, RB10, RPB12, SM12a, WJ12, Yu11]. Applications [Arr18, GGR10, HRZN10, VA17, Aus11, BD11, JS12, KB12, RA13, Whi11, vdMvdMV12, BM10b, Del11c, Pai13b]. Applied [HMS16]. Applying [MKP12, SK12a, HW13]. Approach [Jai11, MD12, NGD14a, NGD14b, NP16, RK16, BKMJ12, BK11, CSKB11, CSKB12, CSKB13, CJ10, Dah10, DBK^{+}13, Eis12b, eAMO10, GB11, GDF13, GKK11, GC12, JG13, JRX12, JDV12, JZY12, Loc12, Mac10, MKP12, MKB11, MM10a, MBN13, MVGM10, Mor13, NUK13, NB10, RMFO13, RVR12, RVB12, SNS10a, SNS10b, SBS11, SV13, UDA10, YAS11, YO11, dCBS13]. Approaches [KMTD17, GD10, JG12, PASS13, GR12]. apps [MMP^{+}12, Ngo12]. Arbon [Tan12]. Arch [Ber13, Swa12b]. architect [Cho10]. Architecting [GTK17, TCB^{+}12]. Architectural [Pan10, KJ10, MBC10]. Architecture [AFF^{+}16, HMB18, AJP13, Ban10b, Ber11a, Del13, GAWM11, GMCH^{+}13, GSB11, KS13b, LAK10, LGMM10, MRN13, Miy11, SKJ^{+}13, SNS10c, Wer10, Bes13b, Del12c]. Architectures [RFD^{+}18b, RFD^{+}18a]. Archive [ASN19]. area [RPB12]. areas [HBM13]. arising [CS12b]. Art [Tri10b, Vu11]. Artefacts [Sin19]. Artifact [Kri13]. Artificial [HdCH^{+}12, MSS19, Sch19]. ASDM [Jan12]. Asia [WL13]. Asia-Pacific [WL13]. Aspect [SPKM16, CBdRS10, CC13, MKB11, NKS10, SBK13, Tek12b]. Aspect-Oriented [SPKM16, CC13, NKS10, SBK13, Tek12b]. Aspects [BWSF18a, BWSF18b, DST^{+}10, MRJD17, PDS^{+}13, Tra11b]. Assembly [SP13]. assertion [BRD^{+}12]. assertion-based [BRD^{+}12]. assess [JDV12]. Assessing [Mun19, PSJ18]. Assessment [BK16, DR10, DR11a, KS12b, NKS10, kP16, eAMO10, HBM13, HPO^{+}13, Nie12]. assignment [RRSV13]. assurance [Yaz10]. AST [CEH19]. ASTD [MGLF12]. Athanasios [Kie12]. Atomic [HNT16]. Attack [SKE^{+}18, SEK^{+}19, RRSV13, SGM12]. Attacks [KK14, BRS13, BSS12, BSS13c, VS11a, VS11b, YAS11]. attending [Kat13]. Attention [HNT16]. attributes [CPG^{+}12, GD10]. automata [BSS13b, MB12]. automated [CJ10, RA13]. Automatic [RMFO13, SS10a, ZLNP18, dSAVP10]. Automation [BCDE18a, BCDE18b, CBK10, CEH19, Bas10]. Autonomous [GKS17, GKL18a, GKL18b, RFD^{+}18b, RFD^{+}18a]. Availability [CK11a]. avoid [Ber12a]. Avoidance [SGS12a]. AVR [HB10]. Aware [DRO^{+}17, HB10, RFD^{+}18b, RFD^{+}18a]. awareness [BP10].

B [GB10, Rus11, dSAVP10]. bad [SK11]. Balanced [WZ12]. balancing [KAZS14]. Balasubramanian [Ebe13]. Bang [Sch16b]. Bar [WCG^{+}18]. Barcodes [Bel11]. Barnum [Del11a]. Barry [Teo13e]. Based [KS12a, KS13a, KAZS14, LIL13, LAX17, Mot19, Mun19, NGD14a, NGD14b, SPKM16, Ban11, BRD^{+}12, BMMR12, BD11, BSS13b, BMRB10, BDJ10, BK11, BZC^{+}18, CV13, Cat13, CSKB12, DBK^{+}13, Eis12b, Fra11, GT10, HWA12, HB10, JM13, JRX12, JDV12, KS11b, KB12, KSR12, KK13, KB11b, LPP^{+}19, Lon10a, Lon10b, MKP12, MKB11, MM10a, NS10a, PGP13, PM10, PCR12, RMFO13, RRSV13, RB10, RK16, SNS10a, SNS10b, SBS11, SAM13b, SAS11, SK12a, SK13, SV13, Swa12a, Tee10b, TG11, TS11, VM13, VS11b, WJ12, YAS11, dCBS13, Men13, SS10b, YA12, Sau10, Del11c]. Bashan [Teo12d]. Bashar [Ber11d]. basics [Win11b]. basis [DD11, JZY12, SBS11]. Bayesian [JRX12, JDV12]. Be [Ost16a, Ost17, Ost18a]. become [Tra10a]. BeginToReason [FS18]. Behavior [Sun18]. Behavioral [CSKB13, BS12, SK13]. behaviour [SK12c]. Bellagio [Teo12d]. Ben [Teo12a]. benchmark [Gre12b]. Benchmarking [CKS18, Loc12]. benchmarking-inspired [Loc12]. Benefits [HNT16, Swa12a]. Benjamin [Ebe13]. Bente [Bes13b]. better [Bes13a]. Between [Sol19, BBF13, BKP16, CSKB12, GBSL16, KD11]. Beyond [Jin18]. Bibliometric [KBRS17a]. bidirectional [Jai11]. Big [Arr18, Tra10a]. Bill [Mit11]. Binary [Asi18]. Bio [RT13]. Bio-inspired [RT13]. bipartite [GDF13]. BIRT [Teo11]. BIXSAN [VS11a]. Black [Del12a]. Blockchain [TODM19]. blocks [CCM12]. BMCLua [MJCdLF17]. body [LAK10]. Bombosch [Bes13b]. Bondurand [Teo11]. Book [Act11, Aus11, Ban12b, Bel11, BM10a, BM10c, BM10b, Ber10a, Ber11d, Ber13, Bes13a, Bes13b, Cha13b, Cha13a, Coo12, Del11b, Del11a, Del11c, Del12a, Del12d, Del12c, Del12b, Del13, Ebe13, Epp11, Fro12a, Fro12b, Fro13a, Fro13b, Gla12, Gla11, Gou12, Gve13a, Gve13b, Hag11, Hat12, Jah13, Kie13a, Kie13b, Kim13, M.13, Men12, Mit11, Mor13, Ngo11, Ngo12, Pai13a, Pai13b, Pay13, Rog10, Rus11, Saf10, Sam13a, Sau10, Sau13a, Sch12a, Sch13a, Sch13b, St.12, Sto13, Swa12a, Swa12b, Tan12, Teo11, Teo12a, Teo12b, Teo12c, Teo12d, Teo13a, Teo13b, Teo13c, Teo13d, Teo13e, Teo13g, Teo13f, Teo13h, Tra10c, Tra10a, Tra10b, Tra11a, Tri10a, Tri10b, Vu11, Wer10, Whi11, Cho10, Cha13b]. Boundary [BWSF18a, BWSF18b]. boxing [Mir11]. BRASS [HSS^{+}16]. Brave [Del11b]. breach [SGS12a]. Brendan [Coo12]. Brian [Fro12a, Mit11, Tri10b]. Bridge [Sol19]. Bridging [BB11a]. brief [MM13b]. Broadening [AENK16]. Brooks [Tra10c]. browser [VS11a]. buffered [Mir11]. bug [KCS11, Tra12]. bugs [Akb18]. building [Act11, Ebe13, Eis12a, Ngo12, HSS^{+}16, Aus11]. built [SGM12]. built-in [SGM12]. Bullet [FM18d, FM18b]. Burt [Win10a]. Business [Sch13c, JZGH13, Mit11]. Business-efficient [Sch13c]. buy [Kim13]. buy-in [Kim13]. buyers [SGM12].

C [Fro13a, Gla11, SM12a, Teo12c, JWB^{+}18, KB11a]. C# [SSK13]. C/C [SM12a]. caching [WZ12]. Cambridge [BM10c, Rog10, Tri10b]. CAME [SP13]. candidates [CBdRS10]. capability [Act11]. Care [SNR17]. Cares [SNR17]. Carlo [MP17]. Carlyle [Ebe13]. Carol [Del11a]. Carollo [Tan12]. CASA [RFD^{+}18b, RFD^{+}18a]. Case [HJ16, Jie16, LPP^{+}19, GKK11, JK11, MT13, PM10, RB10, SK13, SKS10, SS11, Yaz10]. CASE-tools [Yaz10]. cases [BBF13, CBdRS10, CJ10, GDF13, GC12, MM10a]. Casoni [Bat11]. Catch [Sch18]. cause [DC13]. Cay [Gve13a]. Cellular [BSS13b, BS13]. Cellular-genetic [BS13]. center [Zag13]. Centralized [VA17]. Centric [HJ16]. century [WC10]. Cerroni [Bat11]. Certification [HMS16]. Cesare [Ebe13]. CESI [DJB17]. CH [GR12]. Chai [Bel11]. Chair [Ros12b, Ros12a]. Challenges [BWS^{+}17, DK16, SWMV17, Wei18, BMRB10, DM13, GAWM11, LKM^{+}13]. Change [KBRS17b, PC14, Eis12a, MPR12]. Change-History [PC14]. Change-Proneness [KBRS17b]. Change-Readiness [PC14]. Channel [Bul18]. Characterizing [Ber10a]. CHASE [DST^{+}10, PDS^{+}13]. Chatalbasheva [Teo11]. check [CS12a]. Checking [FSK12, MP17, MJCdLF17]. China [Mei17]. chip [Sch13a]. Chris [Eis12a, Sto13]. Christiansen [Fro12a]. Christina [Kie12]. churn [MW12]. Cities [VA17]. Class [NP16, CSKB12, SM12b, Tee10a, YA12]. Classes [PC14, CS12a, GS10, SK11, SK12d]. Clayberg [Teo12b]. ClearQuest [Teo12d]. client [BSS12, BSS13c]. clients [SKJ^{+}13]. Clif [Teo13j]. climate [MBC10]. Clinical [RRK13]. clones [BSS13a]. Cloud [KAZS14, LIL13, ACS13, BD11, Fro13a, Fro13b, HBM13, PCR12]. Cloud-Based [LIL13, PCR12]. clustered [NUK13]. clustering [MKK^{+}12a, MKK^{+}13a].

Clymer [Sau10]. Co [KOPR16, MRJD17]. Co-located [KOPR16, MRJD17]. coalition [Sau11]. cochlear [RRK13]. Cockburn [Ber10a]. Code [FS18, ANCM12, CMGV13, MW12, MKK^{+}12a, MKK^{+}13a, RMFO13, SK11, SSK13, Tai13, Win10c]. Coevolution [Yu11]. Coffin [Teo12a]. cognitive [BK11]. Cohesion [RC17, Tee10a, TJ12]. Collaboration [RGBR14, MG12]. Collaborations [MSS19]. Collaborative [GKMM18, Roy19, CV13, Fra11, Bes13b]. collections [Lon10a, Lon10b]. Collier [Ban12b]. Collingbourne [Del12d]. colony [BDJ10, SKS10]. colored [ACK12]. Colt [BKP16]. combating [RRSV13]. Combinatorial [SA14, SNR17]. COMETS [CMGV13]. commands [Ban11, Cha13a]. commercialized [Swa12a]. COMMitMDE [GKMM18]. common [Jan12]. communication [Saf10, Sch13a]. Community [Xie16a]. COMP [CKS18]. Comparative [HDKB13, KB11a, JDV12, MM13b, RT13]. Comparison [BKP16, DD11, VB13]. compatibility [CCM12]. Competition [CKS18]. competitive [Act11, JD13]. compiled [TMVB13]. compiler [Hat12]. Complete [NNT^{+}19]. Complex [KAZS14, Sau10, SA16]. Complexity [GWG^{+}17, CSKB11, GS10, KCS11, KB11b, SK10a, SK12a]. compliant [Lan11]. Component [KS12a, KS13a, NGD14a, NGD14b, BDJ10, CSKB12, GT10, KS11b, KB11b, MBN13, PGP13, RT10, TS11, UDA10, SP13]. Component-Based [KS12a, KS13a, NGD14a, NGD14b, GT10, KS11b, KB11b]. components [Ber10a, Gve13c, KS10, MNB13, SNS10a, SNS10b]. composite [KB11b]. composition [Rip10]. comprehensive [JG12]. computation [Swa12b]. Computational [BDM12, Sch13a]. Computer [BM10a, BMRB10, Lev13, kP16, Ber12b, SAM13b, Tra10c]. Computers [Gve13c]. Computing [GL18, MPR12, SM17, Fro13a, Fro13b, Gve13c, KS12b, SNS10a, SNS10b, Teo13e, Tri10b]. Concept [Kra18, TJ12, Yaz10]. conceptual [DBK^{+}13, Tai13]. Concern [Noo18]. Concern-Driven [Noo18]. Concerns [KD11, MKB11]. Concurrency [KL18, SK12b]. Concurrent [Wei18, ACK12]. conditional [Teo12c]. Conducting [DJB17, Geo10]. Conference [CTD19, KOPR16, KOH^{+}18a, KOH^{+}18b, MRJD17, SS16, SNGM19, SAHC19, TDWV17, GJ13]. Conferences [SAS16, Kat13, Kri13, MMM10]. Conferernce [LMS11]. Confidence [XZM13]. Configuration [SA16, GD12]. Configurations [PSJ18]. Configuring [JZGH13]. Connected [GKS17, Teo13e]. considered [Sch13d]. considering [SKT10]. Consistency [DRO^{+}17, TLG^{+}16, CCM12, CS12a, dSAVP10]. Consistency-Aware [DRO^{+}17]. Consistent [DRO^{+}17, BM12]. consolidation [Fro13b]. constant [HK12]. constrained [MB12]. Constraints [ACG^{+}19, Ebe13]. Construction [BZC^{+}18]. consumption [JS18a]. content [JS18a]. contest [MFF^{+}10]. Context [Arr18, RFD^{+}18b, RFD^{+}18a, VA17]. Context-aware [RFD^{+}18b, RFD^{+}18a]. Continuous [GKK^{+}19, Yam18, Teo12a]. Contract [Kra18]. Contributors [Lee18]. Control [GKS17, GKL18a, GKL18b, KS12a, KS13a, HDKB13, KS11b, SRS12, SRS13, Whi11]. Controlled [Net19, Tei18]. controller [HB10]. Controlling [SA16]. Cooper [Hat12]. Cooperative [DST^{+}10, PDS^{+}13]. Coordinates [BM10b]. Core [Gve13a]. Cornell [Gve13a]. Corpus [TMVB13]. correct [Lan11]. cost [BDM12]. COTS [KS10]. Counting [Bul18]. coupling [Gup11, SBB12]. Course [Kra18, Cat13]. Courses [EH19a, EH19b, HPO^{+}13].

Coverage [MSM18, SNR17, JZY12]. Cp [CN11]. CPSWeek [GKS17]. crafting [Bes13a]. Craig [Saf10]. crawler [RFS10]. CRAYSE [RFS10]. CreaRE [HDDS12]. creating [Kie13b]. creative [Sau13b]. Creativity [HDDS12]. crisper [KNOF13]. Critical [APNT16, DK16, MNB13, kP16, BKMJ12, HMS16, MBN13, RT13, SBK13, TS11]. Criticality [MBN13]. Cross [Ber12a, KBRS17b, Ber11a]. Cross-Project [KBRS17b]. Crosscutting [KD11]. Crowd [HJ16, Wan18]. Crowd-Centric [HJ16]. crowdsourced [JY12]. Crowley [Teo13e]. cryptography [Lan11]. CSP [YKF^{+}12]. CUDA^{T M} [SM12a]. Culture [FM18c, FM18a, Sch11b]. Curated [ASN19]. Current [Wel18, GAWM11, Tra13]. curriculum [MM11a, MM11b, MM13a]. curse [Gre12a]. cursor [Jai11]. Customer [HKPS12]. customization [GS12, Teo12d]. CVS [BR10]. Cyber [BZC^{+}18, BWS^{+}17, BWSF18a, BWSF18b, BWS^{+}19, Tur19]. Cyber-Physical [BZC^{+}18, BWS^{+}17, BWSF18a, BWSF18b, BWS^{+}19]. Cybersecurity [vdLR18a, vdLR18b]. Cycle [Sha16, KK13]. cycles [And13]. cyclic [Ban10b].

D [Fro12a, Hat12, Mit11, Swa12b, Teo12b]. Daigneau [Del12b]. Dani [Teo13b]. Daniel [Del11b]. Data [Arr18, GKK^{+}19, Lee18, RRSV13, SA14, VA17, Wei18, ZLNP18, BB11a, BSS13b, BS13, HB10, Jah12, JG12, JZY12, Kie12, KGS11, MKP12, MKK12b, MKK13b, NMPS19, RRK13, SK12c, TG11, VM13, Jah13, Jah13, Ngo11]. Data-driven [GKK^{+}19]. database [Gre12a]. dataset [CMGV13]. David [Del12a, Kie13a, Teo12d, Teo13i]. days [BR10]. deadlocks [Gre12a]. Dealing [Mat19]. Debt [Ald19, AENK16, FCT^{+}17, IOSS18, Eis12a, Eis12b, KNOV12, KNOF13, OKNB11, KNOV12, KNOF13, OKNB11]. debugging [SKT10]. Decision [Ald19, Asi18, RK16, BA13, KK11]. decision-making [BA13]. dedicated [GPC12]. defect [NS10a]. Defects [KD11, SBB12]. definition [KNOF13]. degradations [Lev13]. denial [RRSV13]. Dennis [Del11b]. dependability [GD10]. Dependence [RT10]. dependency [GS12, SKT10, SBB12]. DEPLOY [Rom12]. Deployment [Rom12]. derived [JS18a]. Description [HMB18, Pan10]. descriptions [CBdRS10]. Design [Asi18, Dro16, GWG^{+}17, Kra18, Kre19, NAS10, Pas19, Ber11a, Ber12d, Ber11d, CN11, Del12b, Gla11, Gve13c, KK12a, Lon10a, Lon10b, O’S11, RFS10, RRN13, Sch12b, Sch13a, SS10a, SS10b, Tra10c, Wer10, YO11, Sau13a]. Design-by-Contract [Kra18]. Designers [Teo13b]. Designing [NOFK18b, NOFK18a, O’S11, SAS11, Del11b, Teo13e]. designs [BRD^{+}12]. desktop [SKJ^{+}13, SKJ^{+}13]. Despite [HNT16]. detect [Lan11]. detection [JC13, Men13, NMPS19, SK12c]. Determination [KS11a]. determine [BKMJ12, Loc12]. Develop [REN^{+}14]. developer [Teo13j]. Developers [EH19a, EH19b]. Developing [CSG13, GK12, UDA10]. Development [APNT16, BS17, BM18, BR16, DK16, FCT^{+}17, HMS16, HKPS12, JS18b, Jin18, KMTD17, MT13, Mun19, NGD14a, Roy19, Sha16, Wel18, Ber12e, DD11, Del11c, DM13, Fra11, Geo10, GT10, Jan12, KNOV12, KK13, MG12, Mac10, MSK^{+}10, MM11a, MM13a, MO11, Mor13, OKNB11, PGP13, Rup10, SM12a, Sch13c, SK12a, SV13, Swa12a, Tek12b, WC10, dCBS13, KMTD17, Ber10a, Sch12a]. Developmental [MRJD17]. developments [GJ13]. devices [Bel11]. DevOps [Pay13]. DG [SS10b]. DG-metrics [SS10b]. DGML [SS10b]. DGML-based [SS10b]. Diagram [Asi18, BMMR12, GC12]. diagrams [NK13, SK12b, YA12]. diary [Tra12]. DIAS [MRJD17]. different [VB13]. differentiation [BSS12, BSS13c]. digital [Kie13b]. digraph [UDA10]. Dimensions [GTK17]. Dinesh [Sau11]. dining [BL10]. direction [CK11b]. Directional [JS18a]. Discovering [RVR12]. dispatch [HK12]. distortions [Mat19]. Distributed [KAZS14, MMM10, MMM16, SPKM16, ZS14, Ban11, BM10c, DBK^{+}13, MMM11, MMM13, SAM13b, WZ12, ZS13, Fro13a]. distribution [SRS12, SRS13]. diversity [CA10]. Do [Sch19, CA10, Sch16a]. Doan [Jah12]. document [SK10a]. documentation [JY12]. doesn’t [Win10a]. Dojo [HW13]. Domain [ANCM12, Pas19]. Domain-driven [ANCM12]. Dongarra [Fro13a]. Don’t [SNR17]. Doug [Ban13]. Douglas [Bel11]. Driven [Noo18, ANCM12, GKK^{+}19, GPW17, Mac10, SNS10c]. Drupal [Teo13b]. DTrace [Coo12]. Dumbill [M.13]. Dutson [Pai13a, Teo13c]. Dwells [Ost18b]. DWEVOLVE [TG11]. Dynamic [MKK12b, MKK13b, SM17, SPKM16, SMP19, BRD^{+}12, Bas10, Coo12, Gup11, Jai12, SGM12, Sch13a]. dynamics [DBA13].

e-governance [RB10]. e-market [SGM12]. e-nursing [NS10b]. each [BR16]. Eadline [Ban13]. Early [Adr19, BR10, SK10a, MVGM10]. Easy [SA17, Ber12a]. easy-to-use [Ber12a]. Eclipse [Teo12b]. economic [Swa12a]. Ecosystem [KS11c]. ecosystems [Yu11]. ECSA [RFD^{+}18b, RFD^{+}18a]. Edd [M.13]. Edie [Sau13a]. editing [Teo12b]. Edition [Sau10, SBP19, Tri10b, Cha13b, Cha13a, CEH19, Del12c, Fro12a, Gve13a, Gve13b, Gve13c, Hat12, Ngo11, Teo11, Teo13a, Teo13g, Teo13h, Teo13j, Mit11]. editors [Cha13a]. Education [AH10b, AH11b, AH11c, AH12b, SSJM12, Ard10, AH10a, AH11a, AH12a, AH12c, AH12e, AH12d, AH13a, AH13b, AH13c, AH13d, SFTS18].

Effect [JK11, SS13, vdLR18a, vdLR18b, PM12]. Effective [Ald19, MBN13, RVB12, Kie13a]. Effectiveness [SK11, SK12d, HPO^{+}13, NB10]. effects [KS13b]. Efficiency [FS11, MKP12, MBC10, RRN13, Sch11b, Tee10b]. efficient [HK12, RFS10, Sch13c, Swa12b]. Effort [BM18, DD11, MSK^{+}10, MKS10, NUK13, SK12a, SV13, TG13]. Efforts [SW18]. Egypt [ED12]. Eibe [Ngo11]. Eighth [IOSS18]. elements [Epp11]. Eleven [SAS16]. Elizabeth [Gve13b]. Embedded [Jie16, Aus11, Gve13c, Teo13e, Whi11]. Emerging [CCM^{+}10, SFTS18, TODM19, BDM12]. EMF [BGKS12]. Emotion [Sau13a]. Empirical [CN11, DC13, DJB17, JM13, KKPJ10, KKPJ12, MG12, PCR12, CMGV13, GB11, JK12, LRS11, PASS13, Tai13]. employing [eAMO10, VB13]. Enabled [LNG^{+}13]. enabling [Jai11]. encapsulation [SK11]. encoding [SM12a]. End [Sch16b]. Energy [HB10]. enforcer [BRS13]. enforcing [BRS13]. Engine [SMP19, Ngo12, Ngo12]. engineer [VCPR12]. Engineering [AH10b, AH11b, AH11c, AH12b, Arr18, ASN19, BCKS12a, BCKS12b, BCKS13, BBU^{+}17, BWS^{+}17, BWSF18a, BWSF18b, BWS^{+}19, CTD19, Car18, CD17, CHMW19, Doe10a, Doe10b, Doe10c, Doe10d, Doe10e, Doe11a, Doe11b, Doe11c, Doe11d, Doe12a, Doe12b, Doe12c, Doe12d, Doe12e, Doe12f, Doe13a, Doe13b, Doe13c, Doe13d, Doe13e, Doe13f, Doe14, Doe16a, Doe16b, Doe16c, Doe16d, Doe18, DBK^{+}13, EPBR16, Fra16, FS11, GGR10, GKK^{+}19, GPW17, GR12, HdCH^{+}12, HDDS12, JRG^{+}13, KMR^{+}19, KKPJ10, Kra18, KKPJ12, Kre19, KBRS17a, LLM^{+}12, LLM^{+}13, LSM^{+}10, LNG^{+}13, LMS11, MS19, MSS19, MWR19a, MWR19b, Mei17, MRJD17, Mot19, Net19, Pas19, RJJ13, REN^{+}14, Sau10, Sch16a, Sch18, SAS16, SS16, SSJM12, SFTS18, SNGM19, Tei18, Tek12a, TDWV17, TODM19, UYG^{+}19]. engineering [Ard10, AH10a, AH11a, AH12a, AH12c, AH12e, AH12d, AH13a, AH13b, AH13c, AH13d, Ber10b, BÁ10a, BÁ10b, BÁ10c, BÁ10d, Ber11b, Ber11a, BÁ11, Ber11c, Ber12c, Ber12d, BMRB10, CA10, CV13, Cat13, Dah10, DST^{+}10, ED12, GMCH^{+}13, GKK11, GT10, GD12, HPO^{+}13, JD13, KGS11, LKM^{+}13, LRS11, MFF^{+}10, MM11b, PDS^{+}13, SK10a, Sol12, TCB^{+}12, Tra11b, Tra13, YO11, Zag13, TDWV17, TSEvD10, TSvD^{+}11, Rus11, Hat12]. Engineers [Ost16b, Tra10b]. enhancing [NS10a]. Ensemble [BM18]. ensuring [St.12]. enterprise [Bes13b, Ebe13, Men12]. entire [Cat13]. enumeration [YA12]. Environment [BWSF18a, BWSF18b, Roy19, ACS13, ZS13, Teo13a]. Environments [MMM10, MMM16, Wel18, Fra11, MMM11, MMM13]. Envisioning [Roy19]. Eoin [Del12c]. Equivalence [NP16]. era [Kie12]. Eric [M.13]. Erik [Kie13b]. Erl [Ebe13]. Error [Sch16b, SK11, SK12d, SD11]. Errors [GL18]. ESEC [Kat13]. ESEC/FSE [Kat13]. ESEM [LRS11]. essays [Tra10c, Tra11b]. Essence [CC13, PMM16]. Essential [Ngo12]. essentials [Del11a]. estimate [SNS10a, SNS10b]. estimates [Mat19]. Estimation [BM18, DR11b, GB11, ZS14, BDM12, DD11, NUK13, NAS10, SK10a, SK12a, SV13]. ETDSOA [SNS10c]. Eternal [Ber11b, Ber11a]. Ethical [Ost16b]. ETL [MKK12b, MKK13b]. Evaluating [NB10, Gre12b]. Evaluation [HSS^{+}16, kP16, eAMO10, Kri13, PCR12]. Event [Rus11, SNS10c]. Event-B [Rus11]. Ever [Sha16]. Ever-growing [Sha16]. Everything [Win11a]. Evidence [Kam19, Mot19, CV13, Cat13, JM13]. Evidence-Based [Mot19, CV13, Cat13].

Evolution [KK19, LIL13, RB10, SA16, VCPR12, AJP13, CSKB13, CMGV13, JG12, JK11, SS13, TG11, WJ12]. examples [Del13, Teo13c]. exchange [Sau11]. exclusion [Ban12a]. Excursion [Sha16]. Execution [Bul18, NNTK17, NNT^{+}19, SM17, SWMV17, ACS13, MMP^{+}12, ZS13]. Executions [WCG^{+}18]. Exercise [Asi18]. Experience [SFTS18, CS12b, ED12, Kat13, Kim13, RR11, Sam13a, Gve13b, St.12]. experiences [RR11]. Experiment [BKP16, HAJW13]. Experimentation [Yam18]. Experiments [Net19, Tei18]. Exploration [ACG^{+}19]. Exploratory [BKP16, BBF13]. Explore [Sin19]. Exploring [LKM^{+}13]. extended [LGMM10, NK13, YKF^{+}12]. extending [Bas10, Teo11]. extension [PdMG12]. extensions [MPR12]. Extraction [SP13]. Eye [SBP19].

F# [Tri10a]. Facing [JY12]. factories [RR11]. Failure [Adr19, DC13]. Fairley [BM10a]. FAQs [Not10]. Fault [BB11b, CK11b, HWA12, JM13, MJ11, SKT10]. fault-proneness [JM13]. Faults [RK16, CN11, GB13b]. faulty [Tra11a]. Feasibility [KBRS17b]. Feature [SM17, Ber11a, Jai12]. feature-access [Jai12]. feature/architecture [Ber11a]. features [JS18a, MKK^{+}12a, MKK^{+}13a]. Fedora [Men12]. feed [DD11, SK10b]. feed-forward [DD11, SK10b]. feedback [HAJW13]. Figures [Not10]. File [Dro16, SM17]. files [Tra11b]. finger [Jai13]. finger-tip [Jai13]. Finite [YAS11, MB12, RMFO13]. firmware [Mor13]. first [Ber10a, LKM^{+}13, MFF^{+}10, HdCH^{+}12, RJJ13, Tek12a, TSEvD10]. first-order [Ber10a]. Fisher [Gou12]. Five [SW18]. Flannery [Tri10b]. flexibility [Whi11]. FLOSS [Lee18, RGBR14]. Flow [SA14, MKP12, PMTP12]. flow-based [MKP12]. fluctuation [SKT10]. Flynt [Teo13j]. FMEA [Geo10]. focused [GJ13]. focusing [SGS12b]. Food [Win12a]. Ford [Bes13a]. Formal [GP12, GR12, YKF^{+}12, BS12, JM13, Rom12, SGS12b]. Formalization [Wel18]. Formalizing [ACK12, CCM12, BGKS12]. format [YAS11]. formation [Bas10]. FormSERA [GR12]. formulization [SS10b]. forward [DD11, SK10b]. foster [SM12a]. foundation [BS12, Saf10]. Fourth [XPP19, Fro12a, Teo13g, WL13]. Fox [Fro13a]. Foy [Fro12a]. FP7 [Rom12]. Framework [BS17, GKK11, HJ16, JS18b, Mot19, SA16, AJP13, Lon10a, Lon10b, MT13, Teo12b, TG11]. Frank [Aus11, Ngo11]. Frederick [Tra10c]. Free [KS11c, MG12]. FreeBSD [Coo12]. Frequent [RC17]. FSE [Kat13]. Fulghm [Mor13]. Function [Jon13, SBS11]. functional [GB10, NAS10]. functionality [ANCM12]. functions [SKT10]. fundamental [Del12b]. Fundamentals [Ban13, Teo13d, Gve13a, Sch12a]. Future [LWT^{+}19, NOFK18b, NOFK18a, CK11b]. FutureSmart [Mor13]. Fuzzing [LPP^{+}19]. Fuzzy [BK16, BK11, BA13, BSS13a, JDV12, KKK11, NKS10, SV13].

G [Cha13a, Dek10, Men12]. game [RMFO13]. Games [BCKS12a, BCKS12b, BCKS13, Teo12c]. gap [BB11a]. Gary [Gve13a, Mor13]. GAS [BCKS12a, BCKS12b, BCKS13]. GEF [Teo12b]. General [EPBR16, JRG^{+}13, RJJ13, REN^{+}14]. generalized [Ban11]. generate [CS12a, GC12]. generated [JS18a, NMVS11]. generating [SSK13]. generation [AHS12, BSS13b, BS13, BDJ10, CJ10, Gre12b, JZY12, Mac10, MKP12, Pha18, RMFO13, Rim12, SK13, SD11, SS10a, VM13]. Generic [KK19, SA16, GB11, KK12b, SD11].

Genetic [LWT^{+}19, BS13, GC12, MKP12, MM10a, MNB13, MT13, RRN13, SAM13b]. Geoffry [Fro13a]. Geometric [Rog10]. Geometry [BM10b]. gesture [Del11b]. get [Tra10a]. getting [Kim13]. Giri [Rog10]. Glitch [Tra11a]. Global [CTD19, JS18b, MG12, TCB^{+}12]. Good [MS19, Gre12a, dCBS13]. Goodman [Gve13b]. Google [Ngo12, Tan12]. Gorp [Sau13a]. governance [RB10]. Government [VA17]. Gracious [Ost18a]. graduate [MM11b]. gram [KPA10]. Grammar [LPP^{+}19, AHS12]. Grammar-based [LPP^{+}19]. Graner [Cha13b]. Graph [SPKM16, Sin19, BGKS12, DBK^{+}13, GDF13, MKB11]. Graph-Based [SPKM16, DBK^{+}13]. graphical [CSKB11, LLS12, Teo12b]. graphs [Tee11, YA12]. great [Tra10a]. green [LKM^{+}13]. GREENS [LKM^{+}13]. Gregg [Coo12]. Grey [Kam19, Sol19]. Grounded [Kie13b]. growing [Sha16]. growth [SKT10, Yu11]. Gruber [M.13]. Gruver [Mor13]. GTSE [JRG^{+}13, RJJ13, EPBR16, REN^{+}14]. GUI [ÖZ16a, ÖZ16b]. guidance [BMRB10]. guide [Cha13a, Dek10, Gve13b, Men12, O’S11, Pai13b, Teo13d, Teo13j, Sto13, Teo12d]. guided [Tra12]. guidelines [St.12].

H [Del11c, Ngo11, Tri10b]. Habanero [NMPS19]. Hadoop [Ban13]. Halevy [Jah12]. Han [Whi11]. Han-Michale [Whi11]. handbook [Mit11]. Hands [Bat11, Saf10, Teo13d]. Hands-on [Bat11, Saf10, Teo13d]. Hanisch [Whi11]. Hardback [Rog10, BM10a, BM10c, Tri10a, Tri10b]. hardcover [Sau10]. hardware [Yu11]. harmful [Sch13d]. harness [Kie13a]. Harnessing [Wan18]. Harrop [Tri10a]. Hartson [St.12]. having [Sch12c]. held [OKNB11]. Hell [Ngo11]. Helmke [Cha13b, Teo13h]. Herlihy [Vu11]. Herman [Kie13a]. Heterogeneous [SA17, JS12]. Heuristics [SKE^{+}18]. hidden [Tra11a]. Hierarchical [BK16, YKF^{+}12]. High [XZM13, BSS13a, Bas10, CN11, GB13b, Ngo12, ZS13]. high-level [BSS13a, Bas10]. high-performance [Ngo12]. Higher [SNR17]. Hiroko [Bel11]. History [GL18, PC14, Sch11b]. HJ [HAM^{+}19]. hoc [KSR12]. Holmquist [Kie13b]. Holtsnider [Mit11]. Horstmann [Gve13a]. hours [Pai13a]. House [Ost18b]. HP [Mor13]. Hsuing [Sch13a]. Http [Sch16b]. HTTPS [BRS13]. Hudson [Teo13h, Teo13h]. Hugues [Aus11]. human [DST^{+}10, PDS^{+}13]. Humans [Ber10a]. Humphrey [Act11]. Hunter [Jah13, Tra12]. Huw [Del12d]. Hwang [Fro13a]. Hybrid [KMTD17, GKK11, MM10a, MNB13, PdMG12].

Ian [Ngo11]. IBM [Teo12d]. ICGSE [CTD19]. ICPE [LMS11]. ICSE [CCM^{+}10, Fra16, GGR10, HRZN10, LNG^{+}13, SNGM19, Elb16, Elb17, Elb19, KOPR16, LAK10, LKM^{+}13, OKNB11]. ICSE’2012 [GR12]. ICSSP [KOPR16, KOH^{+}18a, KOH^{+}18b, SAHC19]. ICT [Rom12]. idea [NS10b]. Identification [CBdRS10, ANCM12, MBN13]. identify [SK11, SK12d]. Identifying [SA14, Tee11, MKK12b, MKK13b]. Idle [And13]. IEEE [BM10a, BCDE18a, BCDE18b, Fra16, LIL13, TODM19, XPP19, CEH19]. IEEE/ [CEH19]. If [Sch16a]. illustrative [RVB12]. imbalance [WZ12]. Impact [BR16, Mun19, GB13b, MPR12, MD12, Tra11a]. Impactful [Xie16b]. Imperative [ACG^{+}19]. imperialist [JD13]. implant [RRK13]. implementation [BB11a, RFS10, Wer10]. Implications [MSS19, Swa12a]. important [Kat13]. improve [SS10a]. Improved [KJ10, ÖZ16a, ÖZ16b, KS10].

Improvement [LWT^{+}19, RC17, GJ13, MM10a]. Improving [SM12a, Tei18, HWA12, YO11]. incorrect [Ban12a]. increase [MKP12, MBC10]. Incremental [SEK^{+}19]. independent [VS11a]. Index [KS11a, UDA10]. India [KMR^{+}19, MRJD17, SS16]. Industrial [Fra16, Jie16]. Industries [Jie16]. Industry [DJB17, MS19, MM11b, Sol19, MM10b, Rom12]. inequality [JY12]. inevitable [Eis12a]. inference [GS12]. informal [CJ10]. Information [MMM10, MMM11, MMM13, MMM16, WC10, Cho10, KK11, KS13b, PMTP12, Sau11, Yu11]. infrastructure [Fro13b]. initial [LKM^{+}13]. Injection [KK14]. innovation [Kie13b]. Innovations [MRJD17, Sch13a]. innovative [RVR12]. Input [NP16, Pha18]. Ins [GK12]. Inselberg [BM10b]. Insights [SAS16, Jah13, Sch13b]. inspecting [CBdRS10]. inspection [DBA13, NS10a]. inspired [Loc12, RT13]. Install [Dek10]. intangible [CPG^{+}12]. Integrated [Roy19, Wel18, CSKB13, Rom12]. integrating [SKJ^{+}13, Teo11]. Integration [MSM18, BB11a, Jah12]. Integrity [KK12a]. Intelligence [HdCH^{+}12, MSS19, Sch19, BDM12]. Intelligent [MRJD17, Roy19]. intensive [HBM13]. Inter [Sin19, GS12, MO11]. inter-agent [GS12]. Inter-Related [Sin19]. inter-relationship [MO11]. interaction [Sau11]. Interactions [SA14]. Interactive [MW12]. intercomparison [MBC10]. interface [O’S11]. interfaces [Del11b]. International [BCKS12a, BCKS12b, BCKS13, BBG^{+}13, BBU^{+}17, CTD19, CBK10, CHMW19, DJB17, FS11, GK12, GKMM18, GKK^{+}19, GKS17, GKL18a, GKL18b, GFBE10, GPW17, HRZN10, HdCH^{+}12, HDDS12, HKPS12, KKPJ10, KKPJ12, KNOV12, KNOF13, KMTD17, LLM^{+}12, LLM^{+}13, LZK^{+}18, LSM^{+}10, LIL13, LNG^{+}13, LMS11, OKNB11, RFD^{+}18b, RFD^{+}18a, RGBR14, SBP19, SNGM19, SAHC19, TODM19, TLG^{+}16, TSvD^{+}11, UYG^{+}19, XPP19, LRS11, KOPR16, KOH^{+}18a, KOH^{+}18b, TDWV17, XZM13]. Internet [Fro13a]. Internetware [WL13]. Interoperability [MB12]. interpretation [JWB^{+}18]. Interscience [Tri10a]. Introducing [Kay11]. Introduction [GP12, Sau13b]. intrusion [SK12c]. intuitive [BKMJ12]. Investigating [GB13b]. investigation [Tai13]. Involvement [Yam18]. Ioan [Tra10b]. IoT [Mot19]. IP [Wer10]. irregular [SKT10]. ISBN [BM10a, BM10c, BM10b, Rog10, Saf10, Sau10, Tri10a, Tri10b]. ISDE [MMM10, MMM11, MMM13]. ISEC [MRJD17]. ISHCS [XZM13]. isomorphic [Tee11]. Issue [CB12]. Issues [CS12b, Ost16b, Saa19]. item [Teo12d]. Ives [Jah12]. IWSPM’09 [GFBE10].

J [Fro13a, M.13, Teo11, Teo12b]. Jack [Fro13a]. Jackson [Ber11d]. Jaffe [Mit11]. Jaipur [MRJD17]. James [Act11, Ber13, Sam13a, Swa12b, Tan12, Tra10b]. Jason [Tan12]. Java [Gve13a, Ngo12, BKP16, CS12a, CYWD19, CKS18, KB11a, Lon10a, Lon10b, MPR12, MKK^{+}12a, MKK^{+}13a, NNTK17, NNT^{+}19, PdMG12, SM12b, SWMV17, SMP19, vdMvdMV12, Del13]. JavaScript [Kie13a, Teo13d, Kie13a, Teo12a, Teo13d]. Jazz [Teo12d]. Jean [Rus11]. Jean-Raymond [Rus11]. Jeff [O’S11, Sam13a, Tan12, Tra11a]. Jerome [Aus11]. Jih [Sch13a]. Jih-Sheng [Sch13a]. Jim [Coo12]. job [Sch12c, Tra10a]. John [Aus11, BM10a, Hag11, Sau10]. Johnson [Epp11, O’S11]. Joint [GKK^{+}19]. Jon [Fro12a, Tri10a]. Jonge [Teo13c, Ngo12]. JPF [BA19, HAM^{+}19, WKG17, WCG^{+}18]. JPF-HJ [HAM^{+}19]. JPR [WKG17]. jQuery [Pai13a, Teo13c, Teo13c]. Jr.

[Tra10c]. June [GR12]. JustInMind

[FV11]. JVM [WKG17].

K-Medoids [BB11b]. Kai [Fro13a]. Kato [Bel11]. Kd [BB11b]. Kd-Tree [BB11b]. Keith [Hat12]. Ken [Ban12b]. Keng [Bel11]. Key [BR16, MM10b]. keywords [SS10a]. Khalgui [Whi11]. Kirk [Del13]. Klein [Tra12]. Kliem [Gla12]. Knoernschild [Del13]. Knowledge [Car18, AJP13, LAK10, dCBS13]. knowledge-based [dCBS13]. Korat [ACG^{+}19]. Kuniavsky [Gve13b]. Kyle [Pay13].

L [BM10c, Gla12, Gou12]. Lack [HNT16]. ladder [Sch13d]. Lag [SKT10]. Lahman [Del11c]. Landscape [Wel18]. Langade [Bes13b]. Language [Mun19, Sin19, Mac10, HMB18]. languages [KK12b, Pan10, Sch11a]. Large [KS11c, Ber12d, CB12, DM13, MM13a, Mor13]. large-scale [DM13, Mor13]. Larry [Fro12a]. Lars [Kie13b]. LaserJet [Mor13]. lattices [TJ12]. Laurent [M.13]. laws [SS13]. Layered [GSB11]. layout [Jai12]. Leadership [Act11]. Leading [BM10a]. learned [Ber12e]. Learning [KBRS17b, PSJ18, CV13, Del12a, MKS10, Ngo11, RPB12, M.13, Teo13d]. lecture [Cat13]. legacy [Rom12]. Lehman [Sch13e, SS13]. Letter [Ros12b, Ros12a]. Level [BCDE18a, BCDE18b, BSS13a, Bas10, Sol12, VS11b]. Lewis [Sam13a]. Library [ÖZ16a, ÖZ16b, Jai13]. license [Men13]. Licenses [SRS12, SRS13]. Lichty [Sch13b]. lie [Win11a]. Life [Sha16, KK13, Sch12d]. lifecycle [Rup10]. Lightstone [Tra10a]. like [JS12, Sau13b]. Limitations [Sch12b]. limits [Sch10, Sch11a]. Linda [Hat12]. Line [Tek12a, Rip12]. linear [Ber10a]. lines [dCMMdA12]. lingering [Akb18]. link [VS11b]. linked [MM11b]. Linux [Cha13a, Men12, Wer10]. Lion [Sto13, Teo13i]. LISISAP [VS11b]. Literature [Kam19, Sol19, MJ11]. Load [KAZS14, SAM13b, WZ12]. Load-balancing [KAZS14]. loaders [SM12b]. localization [HWA12]. located [KOPR16, MRJD17]. Logic [ZLNP18, BMMR12, NKS10, Pha18, Sch13d]. logic-based [BMMR12]. logs [RVB12]. long [Cat13]. Loop [Tur19]. Lord [Tra11b]. Low [Bas10, CN11, GB13b]. Low-level [Bas10]. LR [SD11]. Lua [MJCdLF17]. Luisa [Bat11].

M [Del11a, Wer10]. Mac [Coo12]. machine [MKS10, Ngo11, RPB12, YAS11]. machines [ACK12, BS12, RMFO13]. Maintainability [DRD12, KS11a, DR11a, JDV12, NAS10, UDA10]. Maintenance [LIL13, CPPC12, JG12, KPA10, PA10]. make [Jah13]. Making [Ald19, BA13, Ber12a, Sch10, Win10b, Tra10a]. Mallows [CN11]. Malware [JC13]. manage [RRK13, Win10a]. Management [GFBE10, HNT16, JS18b, SA16, BGS^{+}13, Fra11, GSB11, HB10, Lee10, MM10b, MM13b, NS10a, Teo12d]. manager [Mit11]. Managing [AENK16, Ber12b, Eis12a, FCT^{+}17, IOSS18, KNOV12, KNOF13, OKNB11, CSG13, Sch13b, BM10a, Gla12, Sch13b]. Mantle [Sch13b]. manual [Teo13i]. map [BK11]. Mapping [NK13, CV13, JS18a, NAS10]. Marc [Saf10]. Maria [Bat11]. Marilyn [Gve13c]. Mark [Cha13a, Men12, Ngo11]. market [SGM12]. MARTE [GPC12]. Martin [Gla11, Gou12]. Master [EH19a, EH19b, Cat13]. masters [Sol12]. matching [KPA10, PA10]. mathematical [MSK^{+}10]. MATLAB [Epp11]. matrix [UDA10, YKF^{+}12]. Matt [Teo13g]. matters [Sch13e]. Matthew [Bes13a, Cha13b, Teo13h]. Maurice [Vu11]. Maurizio [Bat11]. Mauro [Coo12]. May [Fra16, SNGM19]. McCool [Ber13, Swa12b].

McCormick [Aus11]. McCullough [Bes13a]. McMurtry [Saf10]. MDE [GKMM18]. Mean [Sch19]. means [NUK13]. measure [CSKB11, CPPC12, KB11b]. Measurement [NGD14a, LRS11, VB13]. Measuring [KCS11, KKK11, PC14, SKT10, Tee10a, Sin13]. mechanism [KS11b, WJ12, YO11]. media [Teo13e]. media-rich [Teo13e]. mediocracy [Sch13e]. medium [CN11]. Medoids [BB11b]. Meeting [KMR^{+}19]. Memory [JWB^{+}18]. menu [Jai11, Jai12]. Merani [Bat11]. Mercuri [Saf10]. MESOCA [LIL13]. messages [Ban10b]. Meta [SKE^{+}18]. Meta-Heuristics [SKE^{+}18]. Metamorphic [XPP19]. metaphor [KNOV12]. Method [BA19, GD12, Tee10b, TJ12, eAMO10, HK12, JS18a, NAS10, Rip12, dSAVP10]. Methodological [Saa19]. methodologies [MM11a]. methodology [RRN13]. Methods [GP12, GR12, HMS16, MKS10, Rom12, Tee10a, Tee11]. Metric [Kay11, MSM18, Jon13, NS10a, PGP13]. Metrics [BR16, DR11b, HNT16, KB11a, SP13, Sin13, CN11, CMGV13, DR10, DR11a, GB13a, Gup11, JM13, JK11, JK12, Loc12, MW12, MO11, NAS10, NMVS11, PASS13, PM12, SGS12b, SK11, SK12d, SS13, SSK13, Yaz10, CCM^{+}10, SS10b]. Michael [Ber11d, Ber13, Gou12, Rog10, Swa12b]. Michal [Pai13b]. Michale [Whi11]. Micky [Sch13b]. micro [HB10]. Microsoft [Teo13f, Saf10]. Microtasking [Adr19]. Mike [Gve13b, Mor13]. military [Sau11]. Mills [Hag11]. mind [O’S11]. Mining [Mun19, PC14, SAS16, Sin19, MKB11, Ngo11, RVB12]. MiSE [CD17]. missing [Teo13i]. Mitch [Teo13f]. Mitigating [BGS^{+}13]. Mitigation [KK14]. MITM [BRS13]. mixed [eAMO10]. Mobile [Hal13, LNG^{+}13, Teo13c, BD11, Bel11, MKK12b, MKK13b, Pai13a]. Mobile-Enabled [LNG^{+}13]. MOBS [LNG^{+}13]. Model [BZC^{+}18, Bul18, FSK12, GPW17, KAZS14, LLS12, LAX17, Mac10, MP17, MJCdLF17, PM10, Tur19, ACS13, AHS12, BSS13a, CS12a, CSKB12, CK11a, DD11, DBA13, DRD12, GS12, GKK11, GS10, GB10, Jan12, JDV12, KS10, KS11b, KK12a, KJ10, KKK11, MSK^{+}10, MBC10, MKK12b, MKK13b, SNS10c, SAS11, SK13, SK12d, SK12c, Tai13, Tee10b, Jie16, Del11c]. Model-Based [LAX17, BZC^{+}18, PM10, Del11c]. Model-driven [GPW17]. Model-in-the-Loop [Tur19]. Modeling [Ban10a, GWG^{+}17, KK14, KK11, SM12b, ZS13, Ban11, BGKS12, CSG13, Rip12, SAM13b, SVT13, Rus11]. Modelling [BM10c, CD17, GKMM18]. Models [BBG^{+}13, BWSF18a, BWSF18b, Jin18, Sha16, BMMR12, JZGH13, KSR12, KK13, Rup10]. modern [Pai13b, Teo13d, Teo13e]. Modified [GT10, KS12a, KS13a, MKK12b, MKK13b]. modularity [Del13]. Moed [Gve13b]. Mohamed [Whi11]. monitoring [WJ12]. Monte [MP17]. Monte-Carlo [MP17]. MOOCs [AH12e]. Moose [Nie12]. Moral [Ost16b]. MORSE [GPW17]. Moscow [Mir11]. Mountain [Sto13, Teo13i]. movement [Jai11]. Movements [SBP19]. MTD [AENK16, FCT^{+}17, IOSS18]. Multi [BA13, JS18a, MWR19a, MWR19b, BMMR12, BGS^{+}13, GSB11, HK12, KPA10]. Multi-Agent [MWR19a, MWR19b, GSB11]. multi-diagram [BMMR12]. Multi-Directional [JS18a]. multi-method [HK12]. multi-patterns [KPA10]. Multi-person [BA13]. multi-threats [BGS^{+}13]. multicore [Sch12a]. Multidimensional [BM10b]. multilayer [DRD12]. multiple [JS12]. multiprocessor [Vu11]. mutant [Tee10a]. mutual [Ban12a].

Nam [Fro13b]. Narasimhan [Rog10]. Nathaniel [Bes13a]. Natural [Sin19, Del11b, Mac10]. navigation [JS18a]. Neal [Bes13a]. necessary [MBC10]. need [CA10]. negotiation [Mat19]. nested [Jai11]. net [Doe10a, Doe10b, Doe10c, Doe10d, Doe10e, Doe11a, Doe11b, Doe11c, Doe11d, Doe12a, Doe12b, Doe12c, Doe12d, Doe12e, Doe12f, Doe13a, Doe13b, Doe13c, Doe13d, Doe13e, Doe13f, Doe14, Doe16a, Doe16b, Doe16c, Doe16d, Doe18]. nets [ACK12]. Network [GGR10, KAZS14, Sau11, DD11, JRX12, SBS11, Sch13a, ZS13]. network-on-chip [Sch13a]. networking [Bat11]. Networks [ZS14, KSR12, SK10b, Rog10]. neural [DD11, SBS11, SK10b]. news [Not12]. Next [BCDE18a, BCDE18b]. NEXTA [BCDE18a, BCDE18b]. Nick [Del12c]. Nigel [Saf10]. Nine [SS16]. Ninth [FCT^{+}17]. Nir [Vu11]. NLP [SKJ^{+}13]. No [FM18d, FM18b, Gre12a]. non [ACK12, Ber10a, HK12, Lan11]. non-compliant [Lan11]. non-concurrent [ACK12]. non-constant [HK12]. non-linear [Ber10a]. Nordin [Teo13b]. Notation [DR18, LLS12]. note [LG12]. Notes [APNT16, KBRS17a, Doe10a, Doe10b, Doe10c, Doe10d, Doe10e, Doe11a, Doe11b, Doe11c, Doe11d, Doe12a, Doe12b, Doe12c, Doe12d, Doe12e, Doe12f, Doe13a, Doe13b, Doe13c, Doe13d, Doe13e, Doe13f, Doe14, Doe16a, Doe16b, Doe16c, Doe16d, Doe18]. Novel [NGD14b, NP16, GDF13, JG13, SV13]. NUI [Del11b]. Number [RK16]. Numerical [NP16, Tri10b]. nursing [NS10b]. Nuseibeh [Ber11d].

Object

[DR11b, KS11a, KBRS17b, KB11a, MSM18, RC17, CN11, DR10, DR11a, DRD12, GB11, GS10, GB13a, GB13b, Gup11, HK12, JG13, JK12, KK12a, KCS11, KK12b, MJ11, PM12, SGS12b, SK11, Sin13, Teo13g, YA12].

Object-Oriented [DR11b, KBRS17b, KB11a, MSM18, RC17, CN11, DR10, DR11a, DRD12, GB11, GB13a, GB13b, Gup11, SGS12b, SK11, Sin13, Teo13g]. objective [HPO^{+}13]. Objectives [HSS^{+}16]. Observational [Saa19]. Observing [Gve13b]. OCL [JM13]. Oct [Ber10a]. off [Hag11]. off-the-shelf [Hag11]. Office [Teo13f]. official [Cha13b]. OLAP [NK13]. Old [Sha16]. One [Lee18]. One-Time [Lee18]. only [Sch16a]. OO [CSKB13]. OODPM [Dro16]. OOPLs [KK19]. Open [Jin18, KS11c, RGBR14, VA17, CV13, JK11, JK12, MG12, SK12d, Tai13]. open-source [Tai13]. OpenSym [RGBR14]. operating [SGS12a]. operations [BGKS12, Sau11]. Opportunities [DK16]. Optimal [PGP13]. optimally [BGS^{+}13]. optimization [Jai13, MM10a, RT13, SAM13b, SKS10]. optimize [GC12]. optimized [JS18a]. Optimizing [Jai12, Lee10, RRN13]. Oracle [Coo12, SBS11]. oracles [SD11]. Order [SNR17, SMP19, Ber10a]. ordering [Ban10b]. org [Tra10a]. Organizational [GTK17]. organizations [Sch10]. Organizing [LAK10]. Orientation [HRZN10]. Oriented [BS17, DR11b, HKPS12, KS11a, KBRS17b, KB11a, LLM^{+}12, LLM^{+}13, LSM^{+}10, LIL13, MSM18, RC17, SM17, SPKM16, CC13, CN11, DR10, DR11a, DRD12, Fro13b, GB11, GS10, GB13a, GB13b, Gup11, HK12, JG13, JK12, KK12a, KCS11, KB12, KGS11, KK12b, MJ11, NB10, NKS10, PM12, SKJ^{+}13, SNS10c, SGS12b, SK11, Sin13, SBK13, Tek12b, Teo13g]. Orwant [Fro12a]. OSGi [GD10, Del13]. OSS [Mun19]. other [BR16, MO11]. OTM [MMM10]. Our [Xie16a, Kim13]. outlawed [Sch16a]. outlaws [Sch16a]. Outward [Xie16a]. Ovans [Tra11b]. ownership [CB12].

P [Tri10b]. Pacific [WL13]. Pamela [Ber11d]. Panel [FM18d, FM18b]. Panels [FM18c, FM18a]. Pankratius [Sch12a]. Pao [Sch13a]. Pao-Ann [Sch13a]. Paper [LMS11, SAS16]. paperback [Saf10]. papers [Sch12d]. Papows [Tra11a]. paradigm [BD11, BDM12, NS10a]. Parallel [HAM^{+}19, MP17, ÖZ16a, ÖZ16b, Aus11, Ber13, Fro13a, SM12a, Swa12b, BKP16, BM10b]. Parameterized [PA10, KPA10]. Pardha [St.12]. parser [SD11]. Part [Sch19, Sch16a, Sch18]. Partial [SMP19, BM12]. Participation [KS11c]. partition [WZ12]. Partitioning [NP16]. Passages [Gro13a, Gro13b, Gro14, Gro16a, Gro16b, Gro16c, Gro16d, Gro17a, Gro17b, Gro17c, Gro18a, Gro18b, Gro18c, Gro18d, Gro19c, Gro19a, Gro19b]. Pat [Mor13]. Path [NGD14a, JZY12, Del12a]. Path-Based [NGD14a]. pathfinder [KPP12, CS12a, MPR12, NNTK17, NNT^{+}19, PdMG12, SM12b, ZCW12, vdMvdMV12, SMP19]. patient [RRK13]. Patrick [Teo13e]. Pattern [KB12]. Patterns [LZK^{+}18, MRN13, RC17, Bes13a, Del12b, Del13, Ebe13, KPA10, MGLF12, RR11, RVR12, Swa12b]. Paul [Teo13h]. Pautasso [Ebe13]. peaks [GMCH^{+}13]. People [MM10b, Ber10a, Sch13b]. perceptron [DRD12]. Performance [GBSL16, LMS11, ZS14, GKK11, Ngo12, WC10]. perl [Fro12a]. permissions [SGS12a]. persistence [KJ10]. person [BA13]. Personal [Kie12]. Personality [GBSL16, CA10, VCPR12]. perspective [Mat19, VM13]. Perspectives [AENK16, Sch11b]. pervasive [Teo13e]. PESOS [LLM^{+}12, LLM^{+}13, LSM^{+}10]. Peter [Teo13e]. Petri [ACK12]. PhD [HAJW13, Pas19, SW18]. Phil [Pai13a, Teo13c]. philosophers [BL10]. Physical [BZC^{+}18, BWS^{+}17, BWSF18a, BWSF18b, BWS^{+}19, Tur19]. Pithy [Ber10b, BÁ10a, BÁ10b, BÁ10c, BÁ10d, BÁ11, Ber11c, Ber12c]. planning [Mir11]. plate [Men13]. Platform [VA17, BRD^{+}12, GD10, MD12]. pliability [PGP13]. Plug [GK12]. Plug-Ins [GK12]. Pocket [Sto13]. Poetry [Win10c]. Pogue [Teo13i]. points [Jon13]. Post [DJB17, TLG^{+}16]. Post-workshop [DJB17]. power [Kie13a]. Powered [PMM16]. pp [BM10a, BM10c, BM10b, Rog10, Saf10]. pp. [Sau10, Tri10b]. PPCA [Men13]. PPCA-based [Men13]. practical [Cha13a, Dah10, Men12, Mor13, Ngo11, RVR12, Sam13a]. Practice [Car18, Fra16, MS19, Xie16b, Bat11, CBK10, ED12, GAWM11, Tra13]. Practice-Impactful [Xie16b]. Practices [MS19, Kie12, dCBS13]. practitioner [Gve13b]. Practitioners [Kam19]. Pragmatic [HBM13]. PRAT [Kay11]. Precise [KK12b]. predict [JDV12]. predicting [CN11]. Prediction [BB11b, KBRS17b, RK16, DRD12, GS12, GKK11, HPO^{+}13, JRX12, MKS10, MJ11, MVGM10, RPB12, SK10b, SVT13]. predictor [JM13]. Prepared [Ost16a]. Preprocessing [RVB12]. Presentation [Bes13a]. presentations [Bes13a]. Press [BM10c, Rog10, Tri10b]. prevent [YAS11]. preventing [BRS13]. Prevention [BSS12, BSS13c, VS11a, VS11b]. preventive [CPPC12]. Preview [Ost16b]. Principles [EPBR16, Jah12, LLM^{+}12, LLM^{+}13, LSM^{+}10, Ber12d, Ebe13, Gou12, Gve13c, HW13]. prioritisation [GDF13]. Prioritization [CPG^{+}12, HKPS12, SNR17, BA13, MT13, PM10, SKS10, SS11, VB13]. Prioritizing [Akb18]. privacy [Kie12, Teo13f]. Private [Fro13b]. proactive [BGS^{+}13]. Probability [LAX17]. Problem [HRZN10, BL10, Ban11, Gre12a, Sau13b]. problems [BP10, WZ12]. Proceedings [HMS16]. Process [Rip10, DBA13, GKK11, Geo10, GT10, GJ13, JZGH13, Lee10, MKK12b, MKK13b, NS10a, St.12, YO11, Hag11, Teo13g].

Processes [KOPR16, KOH^{+}18a, KOH^{+}18b, SAHC19, MKK12b, MKK13b, Sch13c]. Processing [Sin19, Fro13a, Sch13a]. Product [GFBE10, Tek12a, GJ13, dCMMdA12, Rip12]. product-focused [GJ13]. Productivity [Mun19, YO11]. Products [SA16, dCMMdA12, Kie13b]. PROFES [GJ13]. Professional [EH19a, EH19b]. profile [VCPR12]. Program [HAM^{+}19, WCG^{+}18, Gre12b, Sol12, Tai13]. Programmer [GBSL16, Sau13b]. Programming [KK19, Mun19, RRK13, Sch11a, Swa12b, Teo13a, SBP19, Ber13, Cha13a, Vu11, Fro12a]. Programs [CYWD19, DRO^{+}17, MJCdLF17, MSM18, PSJ18, SM17, Wei18, JG13, KK12b, NMPS19]. Progress [WCG^{+}18]. Project [JS18b, KBRS17b, MM13b, Rom12, WC10]. Projects [BR16, KS11c, CB12, Fra11, Gla12, GB13b, MG12, BM10a, Kim13]. prolific [CJ10]. Promising [BWS^{+}17]. PROMPT [EH19a, EH19b]. prone [SK11, SK12d]. Proneness [KBRS17b, JM13]. Properties [CYWD19, GPC12]. property [NAS10]. Proposal [eAMO10, MSM18]. proposed [SK12c]. prospects [BP10]. protection [Kie12]. protocol [MD12, VS11b]. protocols [HDKB13, KSR12]. Psygkas [Kie12]. Public [Neu11b, Neu12a, Neu13e, Neu14, Neu16a, Neu16b, Neu16c, Neu16d, Neu17a, Neu17b, Neu17c, Neu18a, Neu18b, Neu18c, Neu18d, Neu18e, Neu19a, Neu19b, Neu19c, Neu19d, Neu10a, Neu10b, Neu10c, Neu10d, Neu10e, Neu11a, Neu11c, Neu11d, Neu12b, Neu12c, Neu12d, Neu12e, Neu12f, Neu13a, Neu13b, Neu13c, Neu13d, Neu13f]. Publications [SAS16]. published [BM10a, BM10c, BM10b, Ber10a, Rog10, Saf10, Sau10, Tri10a, Tri10b]. Publishing [Saf10]. Purpose [FS18]. Pursuit [Xie16b]. Pyla [St.12]. q [KPA10]. q-gram [KPA10]. Q’Facto [KS10]. Qualitas [TMVB13]. Qualitas.class [TMVB13]. qualitative [KS13b]. Qualities [AFF^{+}16]. Quality [BK16, GWG^{+}17, HNT16, Kay11, KGS11, MM10a, Tei18, BK11, JDV12, KS10, RPB12, Sin13, St.12, Yaz10]. Quality-oriented [KGS11]. quantification [KK12a]. Quantifying [ACG^{+}19, MSK^{+}10, Sam13a]. Quantitative [kP16, CPPC12, HPO^{+}13, PMTP12]. queries [NK13]. querying [LLS12]. questions [Ber11b]. quotes [Ber10b, BÁ10a, BÁ10b, BÁ10c, BÁ10d, BÁ11, Ber11c, Ber12c]. QVT [LLS12].

R [Ber10a, Fro13b, Sam13a, Sau10, Tra10c]. race [NMPS19]. Radial [SBS11, DD11]. Rady [Teo12a]. Rago [Teo13a]. Rails [M.13, Teo12a]. RAISE [HdCH^{+}12]. Raj [Ebe13]. Rajlich [Tra13]. Ralph [Gla12]. Random [Gre12b]. Ranged [FSK12]. ranked [MG12]. Rankin [Pay13]. Rapid [GKK^{+}19]. rate [RRSV13]. rates [Yu11]. rational [Teo12d]. rationale [NK13]. Raymond [Rus11]. RCoSE [GKK^{+}19]. RE [SAS16, JD13]. re-engineering [JD13]. Reactive [BM10c]. readability [SM12a]. Readiness [PC14]. ready [AH12e, Del11a]. real [Aus11, HDKB13, MKK12b, MKK13b]. real-time [Aus11]. Realizing [HdCH^{+}12]. Reasoning [KL18, Sun18, Wei18]. Recipes [Tri10b, Teo13c]. recommendation [KK13]. Recommendations [Cho10]. reconfigurable [Sch13a, Whi11]. recovery [SD11]. Red [Men12]. Reddy [Gla11]. reducing [Jai11, TG13]. Reduction [LAX17, SMP19, MSK^{+}10]. REEW [FS11]. refactor [SK11]. refactoring [CS12b, SK12d]. Reference [Sun18, Dek10]. Refinement [MGLF12]. Reflection [Tek12b]. refute [Ban12a]. registration [RB10]. Regression [RK16, DD11, NUK13, PM10, RA13].

Regular [CYWD19, AHS12]. Reinders [Ber13, Swa12b]. Related [Sin19]. relational [LGMM10]. relations [LLS12, Yu11]. Relationship [GBSL16, KD11, BBF13, MO11, NAS10]. Relay [Sch13d]. relevancy [Tee10b]. Reliability [NGD14a, SVT13, TS11, KKK11, MVGM10, SK10b, SKT10]. Reloaded [FM18d, FM18b]. Remarkable [Tra10b]. RePa’16 [LZK^{+}18]. Repair [ZLNP18]. Replaying [WKG17]. replicability [LG12]. Replication [KKPJ10, KKPJ12]. Replications [Net19]. Report [AENK16, BCKS12a, BCKS12b, BCKS13, BBG^{+}13, Ber10a, BWS^{+}19, CD17, CHMW19, DJB17, EPBR16, FCT^{+}17, Fra16, FM18c, FM18a, FM18d, FM18b, GK12, GGR10, GKMM18, GKK^{+}19, GKS17, GKL18a, GKL18b, GPW17, HdCH^{+}12, HKPS12, HAJW13, IOSS18, JRG^{+}13, KKPJ10, KKPJ12, LLM^{+}12, LLM^{+}13, LZK^{+}18, LSM^{+}10, LIL13, LNG^{+}13, LRS11, RFD^{+}18b, RFD^{+}18a, RJJ13, REN^{+}14, SFTS18, WL13, XZM13, CBK10, CS12b, GJ13, KMR^{+}19, KNOV12, KNOF13, MMM10, OKNB11, SSJM12, TLG^{+}16, TSEvD10, TSvD^{+}11]. Repositories [Mun19]. RePriCo’12 [HKPS12]. reputation [SGM12]. Requirement [HJ16, SK10a, SK12a, TG11]. Requirements [Arr18, BS17, BBU^{+}17, FS11, HDDS12, HKPS12, HNT16, LZK^{+}18, Noo18, SAS16, Sin19, UYG^{+}19, BA13, Ber11d, CJ10, DBK^{+}13, GMCH^{+}13, GS12, GB10, KK11, KGS11, VB13, dSAVP10]. Research [ASN19, BP10, DM13, Fra16, GKK^{+}19, IOSS18, KMR^{+}19, KKPJ10, KKPJ12, MS19, MSS19, Mei17, Pas19, RGBR14, SW18, TDWV17, Xie16a, Xie16b, CBK10, CMGV13, Gve13b, HBM13, HW13, Kat13, Kim13, Zag13, HW13, Sam13a]. RESER [KKPJ10, KKPJ12]. resilience [SGM12]. Resolution [Adr19]. RESOLVE [Kra18, SW18, Sun18]. Resource [HSS^{+}16, Ban11, HDKB13]. Responsible [Ost17]. REST [Ebe13, Ebe13]. RESTful [Del12b]. Restructuring [RC17]. Results [Adr19, SS10a]. Resurgence [MSS19]. RET [BBU^{+}17, UYG^{+}19]. retrieval [BDJ10]. Reusability [PM12, Tai13, CC13, eAMO10, GB11, GB13a, JRX12, MSK^{+}10, NKS10, SNS10a, SNS10b]. reusable [BDJ10]. Reuse [TG13]. revelation [MRN13]. Review [Act11, Aus11, Ban12b, Bel11, BM10a, BM10c, BM10b, Ber10a, Ber11d, Ber13, Bes13a, Bes13b, Cha13b, Cha13a, Coo12, Del11b, Del11a, Del11c, Del12a, Del12d, Del12c, Del12b, Del13, Ebe13, Epp11, Fro12a, Fro12b, Fro13a, Fro13b, Gla12, Gla11, Gou12, Gve13a, Gve13b, Hag11, Hat12, Jah13, Kam19, Kie13a, Kie13b, Kim13, M.13, Mei17, Men12, Mit11, Mor13, Ngo11, Ngo12, Pai13a, Pai13b, Pay13, Rog10, Rus11, Saf10, Sam13a, Sau10, Sau13a, Sch12a, Sch13a, Sch13b, St.12, Sto13, Swa12a, Swa12b, Tan12, Teo11, Teo12a, Teo12b, Teo12c, Teo12d, Teo13a, Teo13b, Teo13c, Teo13d, Teo13e, Teo13g, Teo13f, Teo13h, Tra10c, Tra10a, Tra10b, Tra11a, Tri10a, Tri10b, Vu11, Wer10, Whi11, HBM13, MJ11, Pan10, PCR12, SBK13]. Rex [St.12]. Reza [Sch12a]. rich [Teo13e]. Richard [BM10a, Epp11, Teo13a]. Rigorous [GR12]. rigour [LG12]. Riquet [Tra10b]. risk [BK11]. Risks [Neu10a, Neu10b, Neu10c, Neu10d, Neu10e, Neu11a, Neu11b, Neu11c, Neu11d, Neu12a, Neu12b, Neu12c, Neu12d, Neu12e, Neu12f, Neu13a, Neu13b, Neu13c, Neu13d, Neu13e, Neu13f, Neu14, Neu16a, Neu16b, Neu16c, Neu16d, Neu17a, Neu17b, Neu17c, Neu18a, Neu18b, Neu18c, Neu18d, Neu18e, Neu19a, Neu19b, Neu19c, Neu19d, KS13b]. Road [MWR19a, MWR19b]. Roadmap [IOSS18]. Robert [Del12b]. Robison [Swa12b, Ber13]. Robot [GPW17]. Robotics [CHMW19]. Rod [Teo12a]. Ron [Sch13b]. root [DC13]. RoSE’19 [CHMW19]. routing [KSR12].

Rozanski [Del12c]. Rubel [Teo12b]. Ruby [Del12a, Teo12a, Del12d]. rule [BDJ10, KK13]. rule-based [KK13]. Rules [TLG^{+}16, Ban10a, Gou12, Mir11, O’S11, Sch13b]. Runtime [BBG^{+}13]. Russel [Tra11b].

S [Act11, Del11c, Gve13a, St.12, Jai11]. S2ERC [Zag13]. S2PF [ZCW12]. Safe [GKS17, GKL18a, GKL18b, Sol19]. safeguard [SGM12]. Safety [APNT16, DK16, HMS16, SM16, kP16, BKMJ12, JWB^{+}18]. Safety-Critical [APNT16, kP16, HMS16, BKMJ12]. Saffron [LPP^{+}19]. Sam [Tra10a]. Sameer [Wer10]. Sams [Saf10, Pai13a]. sanitizer [VS11a]. Saul [Tri10b]. Sauro [Sam13a]. SCAG [CSKB11]. Scalability [Gou12]. Scalable [MP17]. Scale [KS11c, Ber12d, DM13, MM13a, Mor13]. scales [VB13]. scaling [Gou12]. scattered [MKB11]. SCAV’17 [GKS17]. SCAV’18 [GKL18a, GKL18b]. Scenarios [KL18, BM12]. Scheduling [DRO^{+}17]. Scheme [KAZS14, WZ12]. Scholarly [SAS16]. Schutta [Bes13a]. Science [Pas19, Sch16a, Sch18, Sau11]. Scientific [Tri10b]. scientist [Tra10c]. Scientists [Tri10a]. SCORE [MFF^{+}10]. SCRUM [Jie16, Kay11, PMM16]. SE [Win11a]. Search [MP17, VM13, WCG^{+}18, RFS10, SS10a]. Second [Sau10, TODM19, Del12c, Hat12, Sch12c, BCKS12a, BCKS12b, BCKS13, JRG^{+}13]. Secondary [Car18]. secret [Sch12d]. secure [VS11b]. securing [Pai13b]. Security [Kre19, BP10, CPG^{+}12, GSB11, KK11, KS13b, Lev13, SGS12a, Tra12, Zag13, Teo13f]. SEEd [Ard10, AH10a, AH10b, AH11a, AH11b, AH11c, AH12a, AH12b, AH12c, AH12e, AH12d, AH13a, AH13b, AH13c, AH13d].

SEET [BM18]. SEGarage [ASN19]. Seibold [Sto13]. Selection [dCBS13, GDF13, JS12, KK13, MD12, PGP13, SS11]. selective [SGS12a]. Self [NS10b, CSG13]. Self-* [NS10b]. Semantic [BK16, DR18, JS18a, MKK^{+}12a, MKK^{+}13a]. Semantics [AHS12, BMMR12, YKF^{+}12]. SEMAT [JRG^{+}13, RJJ13]. semester [Cat13]. semester-long [Cat13]. sense [Jah13]. Sensor [GGR10]. Separability [EPBR16]. Separation [MKB11, ZLNP18, Pha18]. SERI [KMR^{+}19]. Series [SAS16, CMGV13, Tek12b]. Server [BSS12, BSS13c, RVB12]. Servers [KAZS14]. Service [BK16, BS17, LLM^{+}12, LLM^{+}13, LSM^{+}10, LIL13, ACS13, CSKB12, Fro13b, GD10, KB12, Lev13, MD12, RRSV13, Rip10, SKJ^{+}13, SNS10c, WJ12, Del12b]. service-based [WJ12]. Service-Oriented [LLM^{+}12, LLM^{+}13, LSM^{+}10, LIL13, Fro13b]. services [SKJ^{+}13, Del12b]. SEsCPS [BWS^{+}19]. SESENA [GGR10]. Session [FM18d, FM18b]. set [Lon10a, Lon10b]. set-based [Lon10a, Lon10b]. Seth [Wer10]. settest [Del11a]. seventh [Cha13b, AENK16]. severity [CN11]. Shadow [NNTK17, NNT^{+}19]. Shailendra [Bes13b]. Shannon [Tra10b]. SHARK [LAK10]. Sharon [Kim13]. Shavit [Vu11]. shelf [Ber12a, Cho10, Hag11]. shelf-ware [Ber12a]. shell [Cha13a]. shell-programming [Cha13a]. Shen [Sch13a]. Sheng [Sch13a]. Shmuel [Teo12d]. SHS [BRS13]. SHS-HTTPS [BRS13]. Side [Bul18, BSS12, BSS13c]. Side-Channel [Bul18]. signature [VS11b]. signing [JC13]. SIGSOFT [KBRS17a]. Silver [FM18d, FM18b]. similarity [Tee11]. Simon [Hag11, M.13]. simple [O’S11, RB10]. simplifications [Ber11a]. Simulation [ZS14, DBA13, Rim12, Sau10]. Simulation-Based [Sau10]. SIMUROSOT [RMFO13]. Singh [Fro12b]. Singhoff [Aus11]. single [Cat13]. Singular [HNT16]. sites [Gou12]. situational [GD12]. six [MG12]. Sixth [SBP19]. size [BBF13]. Slices [SM17]. Slicing [SPKM16, JG13]. SlowPOST [RRSV13]. small [dCBS13]. Smart [BWS^{+}17, BWSF18a, BWSF18b, BWS^{+}19, RFD^{+}18b, RFD^{+}18a, VA17]. smells [SK11]. smelly [SK12d]. Smid [Rog10]. Smoot [Fro13b]. SMT [YA12]. SMT-based [YA12]. SOA [CSKB11, CSKB13, KJ10, RB10, SAS11, Ebe13]. SOAP [Del12b]. SOAP/WSDL [Del12b]. Sobell [Cha13a, Men12]. Social [BWSF18a, BWSF18b, GTK17, Tra11b]. Society [BM10a, Sch16a, Sch18]. socio [MW12]. socio-technical [MW12]. socioeconomic [YO11]. soft [KS12b, SNS10a, SNS10b]. Software [Adr19, Ald19, AFF^{+}16, Ard10, AH10a, AH10b, AH11a, AH11b, AH11c, AH12a, AH12b, AH12c, AH12e, AH12d, AH13a, AH13b, AH13c, AH13d, ASN19, BCKS12a, BCKS12b, BCKS13, BM10a, BM18, Ber12d, BR16, BB11b, BWS^{+}17, BWSF18a, BWSF18b, BWS^{+}19, CTD19, CCM^{+}10, Car18, CBK10, CD17, CEH19, CHMW19, DR11b, EH19a, EH19b, EPBR16, GTK17, GWG^{+}17, GGR10, GKK^{+}19, GFBE10, GPW17, GR12, GBSL16, HdCH^{+}12, HKPS12, HSS^{+}16, JS18b, JD13, Jie16, Jin18, JRG^{+}13, KMR^{+}19, KKPJ10, Kra18, KKPJ12, Kre19, KS11c, KOPR16, KMTD17, KOH^{+}18a, KOH^{+}18b, KPA10, KBRS17a, KBRS17b, MJ11, MS19, MSS19, Mei17, MM13b, MRJD17, Mot19, Mun19, NGD14b, Net19, NOFK18b, NOFK18a, Ost16b, Ost18b, Pas19, RJJ13, REN^{+}14, RC17, RK16, Rup10, SA17, Sch11b, Sch16a, Sch18, Sha16, SRS12, SRS13, SS16]. Software [SPKM16, SFTS18, Sin19, SA16, SNGM19, SAHC19, Tei18, Tek12a, TDWV17, TODM19, Tra13, TSEvD10, TSvD^{+}11, XZM13, vdLR18a, vdLR18b, Act11, AJP13, BKMJ12, BSS13a, BP10, BD11, BDM12, Ber10a, Ber10b, BÁ10a, BÁ10b, BÁ10c, BÁ10d, Ber11b, Ber11a, BÁ11, Ber11c, Ber12c, Ber12e, BMRB10, BDJ10, BK11, CB12, CA10, CV13, Cat13, CSG13, CPPC12, CK11b, CPG^{+}12, CMGV13, Dah10, DC13, DD11, DBA13, DM13, DST^{+}10, DR10, DR11a, DRD12, Eis12a, ED12, eAMO10, Fra11, GAWM11, GMCH^{+}13, GB11, Geo10, GT10, GJ13, Gup11, Hag11, HPO^{+}13, Jan12, JRX12, JDV12, JK11, JK12, Jon13, KCS11, Kri13, KNOV12, KKK11, KS12b, KK13, LAK10, LKM^{+}13, LGMM10, LRS11, MG12, dCMMdA12, MKS10, MFF^{+}10, MRN13, Mat19, MM10b, MM11a, MM11b, MM13a, Miy11]. software [MVGM10, NUK13, Nie12, NMVS11, OKNB11, PGP13, PASS13, PA10, PDS^{+}13, RRK13, RRN13, RPB12, RT10, RR11, Rip12, Rus11, SNS10a, SNS10b, Sch12a, Sch13c, Sch13b, SK10a, SK12a, SK10b, SKT10, SBB12, SK12d, SS13, Sin13, SBK13, SV13, Sol12, SS10a, SS10b, TCB^{+}12, Tek12b, Tra10a, Tra11b, Tra12, UDA10, VCPR12, VM13, Wan18, YO11, Yu11, Zag13, dCBS13, CKS18, Fra16, HMS16, SSJM12, Tan12, Tra11a, Ber11d, Del12c, Doe10a, Doe10b, Doe10c, Doe10d, Doe10e, Doe11a, Doe11b, Doe11c, Doe11d, Doe12a, Doe12b, Doe12c, Doe12d, Doe12e, Doe12f, Doe13a, Doe13b, Doe13c, Doe13d, Doe13e, Doe13f, Doe14, Doe16a, Doe16b, Doe16c, Doe16d, Doe18, Fro12b]. Solaris [Coo12]. solution [Gre12a, RRSV13]. Solutions [BWS^{+}17, Del12b, Ebe13]. solve [WZ12]. Solver [ACG^{+}19]. solving [Sau13b]. Sons [BM10a, Sau10]. Sorting [BKP16]. Sound [SMP19]. Source [Jin18, KS11c, ANCM12, CV13, CMGV13, JK11, JK12, MG12, MKK^{+}12a, MKK^{+}13a, SK12d, SSK13, Tai13]. source-code [MKK^{+}12a, MKK^{+}13a]. sourceforge.net [MG12]. sourcing [JS12]. Space [HK12, Swa12a]. Spanner [Rog10]. Spatial [Ban10b]. specific [Kie13a]. Specification [BM10c, JM13, SGS12b]. specification-based [JM13]. specifications [GB10, SS10a]. spectrum [HWA12]. spectrum-based [HWA12]. speculative [ZCW12]. speed [ZS13]. SPF’s [PSJ18]. SPL [SA16]. Spraul [Sau13b]. Springer [BM10b]. SQL [KK14]. SST’19 [SNGM19]. St [M.13]. Stakeholder [Noo18, Kim13]. Standard [WKG17]. Startups [Ald19]. State [Elb16, Elb17, Elb19, LWT^{+}19, MWR19a, MWR19b, ACK12, Ban10a, BS12, CK11a, ED12, GC12, MB12, RMFO13, YAS11, YKF^{+}12]. state-machines [BS12]. Static [ZS14, KK12b, Lan11]. statistical [Yu11]. statistics [Sam13a]. STCD [BSS12, BSS13c]. Stefan [Bes13b]. Stella [Swa12a]. Step [ZS14]. Stephen [Fro13b, Teo13a]. Sterling [Eis12a]. Stevens [Teo13a]. Stirling [Teo12c]. Story [SS16]. Strategies [dCMMdA12, Kie13b, RMFO13]. Strategy [Net19]. string [KPA10, PA10, YAS11]. Strings [SKE^{+}18]. Strongly [BM12]. Structural [MSM18, CSKB13, Tee11, VM13]. Structure [ZLNP18]. Structured [Ber13, Swa12b]. student [MFF^{+}10]. students [Cat13]. Studies [Car18, DJB17, Saa19, CS12b, MG12]. Study [GBSL16, Jie16, KBRS17a, KB11a, PASS13, BBF13, CV13, DC13, JG12, JDV12, JK11, JK12, RT13, RB10, Yu11]. style [Epp11]. success [MM10b, Gla12]. Suite [LAX17]. suites [BBF13]. summaries [BA19]. Summary [BBU^{+}17, BCDE18a, BCDE18b, CTD19, CEH19, KOPR16, KMTD17, KOH^{+}18a, KOH^{+}18b, SBP19, SNGM19, SAHC19, TODM19, UYG^{+}19, XPP19, HW13, Kat13, LAK10, LKM^{+}13, Tek12a]. suppliers [Ber12b]. Support [Net19, VA17, Bas10, Rip10, WJ12]. Supporting [Car18, Mot19]. Surfing [Doe10a, Doe10b, Doe10c, Doe10d, Doe10e, Doe11a, Doe11b, Doe11c, Doe11d, Doe12a, Doe12b, Doe12c, Doe12d, Doe12e, Doe12f, Doe13a, Doe13b, Doe13c, Doe13d, Doe13e, Doe13f, Doe14, Doe16a, Doe16b, Doe16c, Doe16d, Doe18]. surveillance [Kie12]. Survey [DK16, MO11, SGS12b, GD10, RPB12, SK13, TS11]. Surveys [Lee18]. SV [CKS18]. SV-COMP [CKS18]. Symbolic [Bul18, CYWD19, NMPS19, NNTK17, NNT^{+}19, PMTP12, SWMV17, MMP^{+}12, Rim12, ZCW12]. Symposium [RGBR14, WL13, XZM13, HAJW13, LRS11]. Synergies [HdCH^{+}12]. Synopsis [SW18]. syntactic [MKK^{+}12a, MKK^{+}13a]. syntax [SSK13]. Synthesis [SKE^{+}18, SEK^{+}19]. Synthetic [PSJ18]. SysML [CCM12]. System [BK16, BWSF18a, BWSF18b, DBA13, Dro16, DR11b, HSS^{+}16, KOPR16, KOH^{+}18a, KOH^{+}18b, SAHC19, Bas10, DR10, DR11a, DRD12, DBK^{+}13, GSB11, GPC12, GS10, Gve13c, JC13, JK12, KCS11, KS13b, KK13, RB10, Rus11, SGS12a, SGM12, SK12c, SVT13, SD11, Yu11]. System-Environment [BWSF18a, BWSF18b]. systematic [HBM13, PCR12]. Systems [BS17, BM10c, BZC^{+}18, BWS^{+}17, BWSF18a, BWSF18b, BWS^{+}19, DK16, HSS^{+}16, KS11a, KS12a, KS13a, KMTD17, KBRS17b, LLM^{+}12, LLM^{+}13, LSM^{+}10, LIL13, LNG^{+}13, MWR19a, MWR19b, MMM10, MMM16, MRJD17, Mot19, NOFK18b, NOFK18a, SA17, SNGM19, Tur19, kP16, Ber12a, Ber12d, CC13, CSG13, Del12c, GB13a, HK12, HB10, KS11b, KK11, KB11b, Mac10, MJ11, MMM11, MMM13, NKS10, PdMG12, PM12, SAM13b, Sch10, SBK13, TS11, WZ12, Whi11, WC10, Sau10, Teo13e]. systems-making [Sch10].

T [Bel11, Gou12, Teo11, Tri10b]. Tabatabai [Sch12a]. tabular [Rip12]. tagging [BSS12, BSS13c]. Talk [Sch19, Win11b]. taming [Ber12e]. Tan [Bel11, Fro13b]. tangled [Pai13b]. target [Jai13]. Task [HAM^{+}19, ÖZ16a, ÖZ16b]. tasks [ZS13]. Tcl [Teo13j]. Tcl/Tk [Teo13j]. TCP [Wer10]. TCP/IP [Wer10]. teach [Pai13a]. Teaching [CV13, Cat13, Dah10, Kra18]. Team [Dek10, YO11]. teams [dCBS13, Sch13b]. teamwork [Act11, HPO^{+}13]. Technical [Ald19, AENK16, Ber10a, FCT^{+}17, IOSS18, KNOV12, KNOF13, Eis12b, MW12, OKNB11]. technique [ANCM12, JS12, Jai13, SSK13, Ngo11]. Techniques [BM18, SA14, Bes13a, CK11b, KS12b, PCR12, RT13, SBK13, VB13]. technologies [Kie12, Swa12a]. Technology [Ber10a]. teleteaching [DBK^{+}13]. Temperament [GBSL16]. Template [Dro16]. Test [BCDE18a, BCDE18b, CEH19, JZY12, LAX17, MSM18, Pha18, Rim12, SNR17, SKS10, BKMJ12, BBF13, BSS13b, BS13, CJ10, GDF13, GC12, MKP12, MM10a, MT13, PM10, SBS11, SK13, SS11, TG13, VM13, CBK10]. tester [MBN13]. Testing [BBU^{+}17, CK11b, GWG^{+}17, GB10, Hal13, Jie16, Kay11, MMP^{+}12, NGD14b, ÖZ16a, ÖZ16b, SA14, SK12b, Tur19, UYG^{+}19, XPP19, BD11, CBK10, Del11a, Gre12b, Hag11, KS12b, dCMMdA12, MNB13, PM12, PCR12, RA13, RRN13, SK12a, SBK13, Teo12a, VM13, Fro12b, Hag11]. tests [Tan12]. Teukolsky [Tri10b]. text [RFS10]. Thank [Win12b]. their [MO11]. theories [Mat19]. Theory [EPBR16, JRG^{+}13, RJJ13, REN^{+}14, Bat11, BB11a, KNOV12, Teo12c]. Things [Ber12e, Fro13a]. Think [Sau13b]. Thinking [Xie16a]. third [Cha13a, Ngo11, Teo13a, Teo13j, GFBE10, KNOV12]. Thomas [Ebe13]. thought [Teo13g, Win12a]. thread [RRSV13]. Threat [KK14, BGS^{+}13, GSB11]. threats [BGS^{+}13]. threshold [Eis12b, Loc12]. Tichy [Sch12a]. Tim [Teo13d]. Time [GPC12, Lee18, Mir11, Aus11, CMGV13, HK12, HDKB13, Jai11, Jai12, KS12b, MKK12b, MKK13b, SNS10c, SKT10]. Times [BKP16]. tip [Jai13]. Tk [Teo13j]. Tkatchova [Swa12a]. TLRO [BL10, Ban11, Ban12a]. Tobias [Tra12]. tokens [SD11]. tolerance [CK11b]. Tom [Fro12a]. Tomer [Kim13]. Tomorrow [Ost18b]. Too [Ost18b]. Tool [HAM^{+}19, KS12a, KS13a, SSK13, ZS14, RA13, SRS12, SRS13]. Tools [ASN19, CS12a, CKS18, GK12, Gre12b, Lan11, MM13b, Ngo11, Sch13b, Yaz10]. top [MG12]. TOPI [GK12]. Topology [ZS14]. Torczon [Hat12]. TOSEM [Not10, Not12]. Touch [Jai13, Del11b]. tour [Tra12]. Trace [SM17]. Traceability [CSKB12, SNGM19]. Traces [WKG17]. tracing [Coo12]. Track [RGBR14]. tradeoffs [MBC10]. traditional [MO11]. traffic [VS11b]. Transfer [Car18, KBRS17b]. Transformation [DR18, AHS12, BM12, GPC12]. transformations [BGKS12]. transformed [Mor13]. Transition [Jie16, Ban10a, CK11a, GPC12, YKF^{+}12]. Translator [MJCdLF17]. Tree [BB11b, MP17, RK16, SSK13]. Trends [CCM^{+}10, TODM19]. Trevor [Sau13a]. trouble [Gla12]. Troubleshooting [Pay13]. trust [Act11, KSR12]. trustworthiness [Wan18]. Tulloch [Teo13f]. Turkish [Tek12a, Tek12b]. turnaround [Gla12]. tutorial [Dek10]. Twenty [SW18]. twin [GMCH^{+}13]. two [CB12]. types [BS12].

UAVs [BMRB10]. Ubuntu [Cha13b, Teo13h]. UCFrame [HJ16]. UI [Teo13c]. ultra [Ber12d]. ultra-large-scale [Ber12d]. UML [TLG^{+}16, ACK12, AHS12, BRD^{+}12, BMMR12, Bas10, BS12, GPC12, GP12, GC12, HDKB13, NMVS11, Pan10, SAM13b, SP13, SK12b, SK13, YA12]. UML-MARTE [GPC12]. Understanding [FS18, GTK17, MBC10, O’S11]. unified [Rip12]. Unifying [MKK^{+}12a, MKK^{+}13a]. universal [Jon13]. University [BM10c, Rog10, Tri10b]. UNIX [Teo13a]. Unknownness [RD13]. unleashed [Teo13h]. unmanageable [Sch13b]. Update [KMR^{+}19]. updates [Jai12]. UPnP [Bas10]. Usability [DR11b, DR10, Del11a]. Usage [RC17, RVR12, RVB12]. Use [HJ16, Kam19, BBF13, Ber12a, CBdRS10, GKK11]. User [Yam18, Del11b, Gve13b, JS18a, Kim13, O’S11, Sam13a, SK12c, St.12]. users [Teo13f]. Using [BM18, Dro16, Kra18, Lan11, RC17, SM17, Sin19, WKG17, WCG^{+}18, ACK12, BA13, BSS12, BSS13c, CSG13, CN11, CMGV13, Del13, DR11b, GSB11, GB13a, GC12, GB10, HDKB13, Jai13, JDV12, JD13, JK12, KK14, KS12b, MBN13, MNB13, MT13, MKK12b, MKK13b, NUK13, NAS10, NGD14a, NKS10, ÖZ16a, ÖZ16b, PGP13, Pha18, RRN13, RVR12, SA14, SKE^{+}18, SP13, SK10a, SKS10, SK11, SK12c, SSK13, SS11, Tee10a, Tee11, YO11, ZLNP18, vdMvdMV12]. Uwe [Bes13b]. UX [St.12]. UXSOM [NMVS11].

V [Sau13b, Jie16]. V-Model [Jie16].

Vaclav [Tra13]. Validation

[Gup11, JK12, BRD^{+}12, SGS12b, ZS13]. value [Tee10b]. Values [SNR17, Loc12]. Variability [AFF^{+}16, GAWM11, GWG^{+}17]. variants [MW12, Rip12]. various [SBK13]. Vastu [MRN13]. vehicle [RB10]. Vehicles [GKS17, GKL18a, GKL18b]. Venkatesulu [Wer10]. Verification [BZC^{+}18, CYWD19, CKS18, BRD^{+}12, BMMR12, MBN13, BM10c]. Verifying [vdMvdMV12, CCM12]. Veritesting [SWMV17]. Verma [Sau11]. Version [Dro16, KS12a, KS13a, SM12b, TMVB13]. versioning [KS11b]. Vetterling [Tri10b]. via [Bul18, MP17, Rim12]. Victor [Sch12a]. view [ED12, JG12, MM13b]. Views [GMCH^{+}13]. virtual [Jai13, NAS10]. virtualization [Fro13b]. vision [BMRB10, NB10]. Visual [BM10b, Sch11a]. visualize [Jah13]. visualizing [MPR12]. volume [Gve13a]. Volunteer [KS11c, MG12]. VOSDM [NB10]. vs [Pan10].

W [Act11, Aus11, Sch13b, Teo13a]. WAAPO [HRZN10]. WAAPO-2010 [HRZN10]. Wall [Fro12a]. Walter [Sch12a, Bat11]. ware [Ber12a]. warehouse [KGS11, SK12c, TG11]. warehousing [JG12]. Warranties [vdLR18a, vdLR18b]. Watkins [Hag11]. Watling [Saf10]. Watts [Act11]. ways [Jah13, Kie13a]. Weakly [DRO^{+}17]. Weathersby [Teo11]. Web [TSEvD10, TSvD^{+}11, Del12b, Gou12, MD12, Pai13b, RFS10, RA13, RVR12, RVB12, Rip10, Sch12b, ACS13, BK16]. Web2SE [TSEvD10, TSvD^{+}11]. Weisfeld [Teo13g]. WETSEB [TODM19]. Whitney [Jah13]. Whittaker [Tan12]. Wigdor [Del11b]. wilds [Tra12]. Wiley [BM10a, Sau10, Tri10a]. Wiley-Interscience [Tri10a]. Will [Sch16b, Sch16a]. William [Tri10b]. win [HWA12]. Windows [Saf10]. wireless [KSR12]. wisdom [Wan18]. within [Cat13, Jai11]. Witten [Ngo11]. Wixon [Del11b]. Wolf [Gve13c]. Woods [Del12c]. work [Ber11d, Tra10a, Teo12d]. Works [GKK^{+}19]. Workshop [AENK16, BCKS12a, BCKS12b, BCKS13, BBG^{+}13, BBU^{+}17, BCDE18a, BCDE18b, BWS^{+}19, CCM^{+}10, CBK10, CD17, CEH19, CHMW19, DJB17, EPBR16, FCT^{+}17, Fra16, FS11, GK12, GGR10, GKMM18, GKK^{+}19, GKS17, GKL18a, GKL18b, GFBE10, GPW17, GR12, HRZN10, HMS16, HdCH^{+}12, HDDS12, HKPS12, IOSS18, JRG^{+}13, KKPJ10, KKPJ12, KNOV12, KNOF13, KMTD17, LLM^{+}12, LLM^{+}13, LZK^{+}18, LSM^{+}10, LIL13, LNG^{+}13, OKNB11, RFD^{+}18b, RFD^{+}18a, RJJ13, REN^{+}14, SBP19, SNGM19, TODM19, TLG^{+}16, TSEvD10, TSvD^{+}11, UYG^{+}19, XPP19, HW13, LAK10, LKM^{+}13, Tek12b, MMM10, MRJD17, SSJM12, SFTS18, Tek12a]. World [Sch16b, Del11b]. Worst [LPP^{+}19]. Worst-Case [LPP^{+}19]. Wren [Teo12b]. Wright [Teo13d]. Write [SA17]. WSDL [Del12b]. WUCOR [TLG^{+}16]. Wynn [Teo12c].

X [Coo12, Sto13, Teo13i]. XML [NMVS11]. XP [FM18c, FM18a, FM18d, FM18b]. XP2013 [HW13, HAJW13]. XSS [BSS12, BSS13c, VS11a].

Year [SS16]. Years [SAS16, SW18]. Yogesh [Fro12b]. Young [Mor13]. yourself [Pai13a].

Zachary [Jah12]. Zalewski [Pai13b]. Zave [Ber11d]. Zero [Dek10]. Zürich [GR12].

References

Almaawi:2019:QEK

[ACG^{+}19] Alyas Almaawi, Hayes Converse, Milos Gligoric, Sasa Misailovic, and Sarfraz Khurshid. Quantifying the exploration of the Korat solver for imperative constraints. ACM SIGSOFT Software Engineering Notes, 44(4):15, December 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

                                                     //dl.acm.org/doi/10.1145/
                                                     3364452.3364456.

Andre:2012:FNC

[ACK12] Étienne André, Christine Choppy, and Kais Klai. Formalizing non-concurrent UML state machines using colored Petri nets. ACM SIGSOFT Software Engineering Notes, 37(4):1–8, July 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ahmad:2013:WSE

[ACS13] Faisal Ahmad, Suvamoy Changder, and Anirban Sarkar. Web service execution model for cloud environment. ACM SIGSOFT Software Engineering Notes, 38(6):1–13, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Acton:2011:BRL

[Act11] Dorothy Acton. Book review:

Leadership, teamwork and trust: building a competitive software capability by Watts S. Humphrey and James W. Over. ACM SIGSOFT Software Engineering Notes, 36 (6):28, November 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Adriano:2019:MSF

[Adr19] Christian Adriano. Microtasking software failure resolution: Early results. ACM SIGSOFT Software Engineering Notes, 44(1):36, January 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Avgeriou:2016:TDB

[AENK16] Paris Avgeriou, Neil A. Ernst, Robert L. Nord, and Philippe Kruchten. Technical debt:

Broadening perspectives report on the Seventh Workshop on Managing Technical Debt (MTD 2015). ACM SIGSOFT Software Engineering Notes, 41(2):38–41, March 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Alebrahim:2016:VQS

[AFF^{+}16] Azadeh Alebrahim, Stephan Faßbender, Martin Filipczyk, Michael Goedicke, Maritta Heisel, and Uwe Zdun. Variability for qualities in software architecture. ACM SIG-SOFT Software Engineering Notes, 41(1):32–35, January 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2010:SEEb

[AH10a] Mark A. Ardis and Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 35(4):6–7, July 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2010:SEEc

[AH10b] Mark A. Ardis and Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 35(6):7–8, November 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2011:SEEa

[AH11a] Mark A. Ardis and Peter B.

Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 36(2):8–9, March 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2011:SEEb

[AH11b] Mark A. Ardis and Peter B.

Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 36(3):7–8, May 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2011:SEEc

[AH11c] Mark A. Ardis and Peter B.

Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 36(6):7–8, November 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2012:SEEa

[AH12a] Mark A. Ardis and Peter B.

Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 37(1):10–11, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2012:SEEb

[AH12b] Mark A. Ardis and Peter B.

Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 37(2):8–10, March 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2012:SEEc

[AH12c] Mark A. Ardis and Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 37(3):8–9, May 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2012:SEEe

[AH12d] Mark A. Ardis and Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 37(6):8–9, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2012:SEEd

[AH12e] Mark A. Ardis and Peter B. Henderson. Software engineering education (SEEd): is software engineering ready for MOOCs? ACM SIG-SOFT Software Engineering Notes, 37(5):14, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2013:SEEa

[AH13a] Mark A. Ardis and Peter B. Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 38(2):9–10, March 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2013:SEEb

[AH13b] Mark A. Ardis and Peter B.

Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 38(3):5–6, May 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2013:SEEc

[AH13c] Mark A. Ardis and Peter B.

Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 38(5):18, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2013:SEEd

[AH13d] Mark A. Ardis and Peter B.

Henderson. Software engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 38(6):9–10, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Arora:2012:SUM

[AHS12] Deepak Arora, Bramah Hazela, and Vipin Saxena. Semantics for UML model transformation and generation of regular grammar. ACM SIGSOFT Software Engineering Notes, 37(3):1–5, May 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ahmad:2013:FAA

[AJP13] Aakash Ahmad, Pooyan Jamshidi, and Claus Pahl. A framework for acquisition and application of software architecture evolution knowledge: 14. ACM SIGSOFT Software Engineering Notes, 38(5):1–7, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Akbarinasaji:2018:PLB

[Akb18] Shirin Akbarinasaji. Prioritizing lingering bugs. ACM SIGSOFT Software Engineering Notes, 43(1):1–6, January 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Aldaeej:2019:TET

[Ald19] Abdullah Aldaeej. Towards effective technical debt decision making in software startups. ACM SIGSOFT Software Engineering Notes, 44(3):22, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3356773.3356793.

Anwikar:2012:DDT

[ANCM12] Vallabh Anwikar, Ravindra Naik, Adnan Contractor, and Hemanth Makkapati. Domaindriven technique for functionality identification in source code. ACM SIGSOFT Software Engineering Notes, 37 (3):1–8, May 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anderson:2013:IC

[And13] Glyn Anderson. Idle cycles. ACM SIGSOFT Software Engineering Notes, 38(4):6–9, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Axelsson:2016:NAS

[APNT16] Jakob Axelsson, Efi Papatheocharous, Jaana Nyfjord, and Martin Törngren. Notes on agile and safety-critical development. ACM SIGSOFT Software Engineering Notes, 41(2):23–26, March 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ardis:2010:SEEa

[Ard10]  Mark A. Ardis.  Soft-

ware engineering education (SEEd). ACM SIGSOFT Software Engineering Notes, 35(2):4, March 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Arruda:2018:REC

[Arr18] Darlan Arruda. Requirements engineering in the context of big data applications. ACM SIGSOFT Software Engineering Notes, 43(1):1–6, January 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Asim:2018:EDB

[Asi18] Saad F. Asim. An exercise in design: The binary decision diagram. ACM SIGSOFT Software Engineering Notes, 43(3):19, July 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Arumugam:2019:SCA

[ASN19] Lakshmanan Arumugam, Vikram N. Subramanian, and Meiyappan Nagappan. SEGarage: a curated archive for software engineering research tools. ACM SIGSOFT Software Engineering Notes, 44(3):13, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3356773.3356777.

Ausden:2011:BRB

[Aus11] Howard Ausden. Book review: Building parallel, embedded, and real-time applications with Ada by John W. McCormick, Frank Singhoff and Jerome Hugues. ACM SIGSOFT Software Engineering Notes, 36(6):28–29, November 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2010:PSEb

[BÁ10a] Larry Bernstein and Hákon Ágústsson. Pithy software engineering quotes. ACM SIG-SOFT Software Engineering Notes, 35(3):6–7, May 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2010:PSEc

[BÁ10b] Larry Bernstein and Hákon Ágústsson. Pithy software engineering quotes. ACM SIG-SOFT Software Engineering Notes, 35(4):3–4, July 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2010:PSEd

[BÁ10c] Larry Bernstein and Hákon Ágústsson. Pithy software engineering quotes. ACM SIGSOFT Software Engineering Notes, 35(5):4–5, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2010:PSEe

[BÁ10d] Larry Bernstein and Hákon Ágústsson. Pithy software engineering quotes. ACM SIGSOFT Software Engineering Notes, 35(6):5, November 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2011:PSEa

[BÁ11] Larry Bernstein and Hákon Ágústsson. Pithy software engineering quotes. ACM SIG-SOFT Software Engineering Notes, 36(2):6–7, March 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bajaj:2013:MPD

[BA13] Punam Bajaj and Vineet Arora. Multi-person decisionmaking for requirements prioritization using fuzzy AHP. ACM SIGSOFT Software Engineering Notes, 38(5):1– 6, September 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Berglund:2019:MSJ

[BA19] Lasse Berglund and Cyrille Artho. Method summaries for JPF. ACM SIGSOFT Software Engineering Notes, 44(4):16, December 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3364452.33644560.

Bandyopadhyay:2010:MST

[Ban10a] Anup Kumar Bandyopadhyay. Modeling of state transition rules and its application. ACM SIGSOFT Software Engineering Notes, 35(2):1–7, March 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bandyopadhyay:2010:SOM

[Ban10b] Anup Kumar Bandyopadhyay.

Spatial ordering of messages in a cyclic architecture. ACM SIGSOFT Software Engineering Notes, 35(6):1–6, November 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bandyopadhyay:2011:TBM

[Ban11] Anup Kumar Bandyopadhyay.

TLRO based modeling of alternative commands and its application to generalized distributed resource allocation problem. ACM SIGSOFT Software Engineering Notes, 36(6):1–5, November 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bandyopadhyay:2012:ATR

[Ban12a] Anup Kumar Bandyopadhyay.

Application of TLRO to refute an incorrect mutual exclusion algorithm. ACM SIG-SOFT Software Engineering Notes, 37(5):1–5, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bangalore:2012:BRA

[Ban12b] Srinivasan M. Bangalore.

Book review: Agile Analytics by Ken Collier. ACM SIG-SOFT Software Engineering Notes, 37(6):43–44, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bangalore:2013:HFD

[Ban13]  Bangalore.  Hadoop fun-

damentals by doug eadline. ACM SIGSOFT Software Engineering Notes, 38(6):43– 44, November 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bastani:2010:LLD

[Bas10] Behzad Bastani. Low-level dynamic system formation with high-level automation: extending UML in support of UPnP. ACM SIGSOFT Software Engineering Notes, 35 (3):1–10, May 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bateman:2011:HNT

[Bat11]  Anita Bateman.  Hands-

on networking: from theory to practice by Maria luisa merani, maurizio casoni, and Walter cerroni. ACM SIG-SOFT Software Engineering Notes, 36(2):38, March 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bennett:2011:BDI

[BB11a] Travis A. Bennett and Coskun Bayrak. Bridging the data integration gap: from theory to implementation. ACM SIG-SOFT Software Engineering Notes, 36(3):1–8, May 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bishnu:2011:AKM

[BB11b] P. S. Bishnu and V. Bhattacherjee. Application of K-Medoids with Kd-tree for software fault prediction. ACM SIGSOFT Software Engineering Notes, 36(2):1–6, March 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Badri:2013:RBU

[BBF13] Mourad Badri, Linda Badri, and William Flageol. On the relationship between use cases and test suites size: an exploratory study. ACM SIG-SOFT Software Engineering Notes, 38(4):1–5, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bencomo:2013:RIW

[BBG^{+}13] Nelly Bencomo, Gordon Blair, Sebastian Götz, Brice Morin, and Bernhard Rumpe. Report on the 7th International Workshop on Models@Runtime. ACM SIG-SOFT Software Engineering Notes, 38(1):27–30, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Borg:2017:SIW

[BBU^{+}17] Markus Borg, Elizabeth Bjarnason, Michael Unterkalmsteiner, Tingting Yu, Gregory Gay, and Michael Felderer. Summary of the 4th International Workshop on Requirements Engineering and Testing (RET 2017). ACM SIG-SOFT Software Engineering Notes, 42(4):28–31, October 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Borg:2018:SIWa

[BCDE18a] Markus Borg, Adnan Causevic, Serge Demeyer, and Sigrid Eldh. Summary of the 1st IEEE Workshop on the Next Level of Test Automation (NEXTA 2018). ACM SIGSOFT Software Engineering Notes, 43(4):36–38, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Borg:2018:SIWb

[BCDE18b] Markus Borg, Adnan Causevic, Serge Demeyer, and Sigrid Eldh. Summary of the 1st IEEE Workshop on the Next Level of Test Automation (NEXTA 2018). ACM SIGSOFT Software Engineering Notes, 43(4):53, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bell:2012:RSIa

[BCKS12a] Jonathan Bell, Kendra M. L. Cooper, Gail Kaiser, and Swapneel Sheth. Report from the Second International Workshop on Games and Software Engineering (GAS 2012). ACM SIGSOFT Software Engineering Notes, 37 (6):1–6, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bell:2012:RSIb

[BCKS12b] Jonathan Bell, Kendra M. L.

Cooper, Gail Kaiser, and Swapneel Sheth. Report from the Second International Workshop on Games and Software Engineering (GAS 2012). ACM SIGSOFT Software Engineering Notes, 37(6):26– 27, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bell:2013:RSI

[BCKS13] Jonathan Bell, Kendra M. L.

Cooper, Gail Kaiser, and Swapneel Sheth. Report from the Second International Workshop on Games and Software Engineering (GAS 2012). ACM SIGSOFT Software Engineering Notes, 38 (1):34–35, January 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Baride:2011:CBS

[BD11] Srikanth Baride and Kamlesh Dutta. A cloud based software testing paradigm for mobile applications. ACM SIGSOFT Software Engineering Notes, 36(3):1–4, May 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhatia:2010:ACB

[BDJ10] Rajesh K. Bhatia, Mayank Dave, and R. C. Joshi. Ant colony based rule generation for reusable software component retrieval. ACM SIG-SOFT Software Engineering Notes, 35(2):1–5, March 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Benala:2012:CIS

[BDM12] Tirimula Rao Benala, Satchidananda Dehuri, and Rajib Mall. Computational intelligence in software cost estimation: an emerging paradigm. ACM SIGSOFT Software Engineering Notes, 37(3):1–7, May 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Belani:2011:BRB

[Bel11]  Hrvoje Belani.  Book re-

view: Barcodes for mobile devices by Hiroko Kato, Keng T. Tan and Douglas Chai. ACM SIGSOFT Software Engineering Notes, 36(3):32–33, May 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2010:BRC

[Ber10a] Larry Bernstein. Book review: Characterizing people as non-linear, first-order components in software development, by Alistair A. R. Cockburn and published in Humans and Technology, HaT Technical Report 1999.03, Oct 21, 1999. ACM SIGSOFT Software Engineering Notes, 35 (4):33–34, July 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2010:PSEa

[Ber10b] Larry Bernstein. Pithy software engineering quotes. ACM SIGSOFT Software Engineering Notes, 35(2):3–4, March 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2011:ESEb

[Ber11a] Larry Bernstein. Eternal software engineering cross feature/architecture design simplifications. ACM SIGSOFT Software Engineering Notes, 36(6):6, November 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2011:ESEa

[Ber11b]  Larry Bernstein.  Eternal

software engineering questions. ACM SIGSOFT Software Engineering Notes, 36 (5):7, September 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2011:PSEb

[Ber11c] Larry Bernstein. Pithy software engineering quotes. ACM SIGSOFT Software Engineering Notes, 36(3):4, May 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Berry:2011:BRS

[Ber11d] Daniel M. Berry. Book review: Software requirements and design: the work of Michael Jackson by Bashar Nuseibeh and Pamela Zave. ACM SIGSOFT Software Engineering Notes, 36(2):39–40, March 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2012:CAS

[Ber12a] Larry Bernstein. Cross avoid shelf-ware by making your systems easy-to-use. ACM SIGSOFT Software Engineering Notes, 37(1):8–9, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2012:MCS

[Ber12b]  Larry Bernstein.  Man-

aging computer suppliers. ACM SIGSOFT Software Engineering Notes, 37(5):12– 13, September 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2012:PSE

[Ber12c] Larry Bernstein. Pithy software engineering quotes. ACM SIGSOFT Software Engineering Notes, 37(2):6–7, March 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2012:SED

[Ber12d] Larry Bernstein. Software engineering design principles for ultra-large-scale systems. ACM SIGSOFT Software Engineering Notes, 37(4):8–9, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bernstein:2012:TLT

[Ber12e] Larry Bernstein. Things i learned from taming software development. ACM SIG-SOFT Software Engineering Notes, 37(6):5–6, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Berzal:2013:BRS

[Ber13] Fernando Berzal. Book review: Structured parallel programming by Michael McCool, James Reinders & Arch Robison. ACM SIGSOFT Software Engineering Notes, 38 (2):35–39, March 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Beschastnikh:2013:BRP

[Bes13a] Ivan Beschastnikh. Book review: Presentation patterns: techniques for crafting better presentations by Neal Ford, Matthew McCullough and Nathaniel Schutta. ACM SIGSOFT Software Engineering Notes, 38(5):67, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Beshilas:2013:BRC

[Bes13b] Bill Beshilas. Book review: Collaborative enterprise architecture by Stefan Bente, Uwe Bombosch, and Shailendra Langade. ACM SIGSOFT Software Engineering Notes, 38(1):54, January 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brosch:2012:FEM

[BGKS12] Petra Brosch, Sebastian Gabmeyer, Gerti Kappel, and Martina Seidl. On formalizing EMF modeling operations with graph transformations. ACM SIGSOFT Software Engineering Notes, 37 (4):1–8, July 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bedi:2013:MMT

[BGS^{+}13] Punam Bedi, Vandana Gandotra, Archana Singhal, Himanshi Narang, and Sumit Sharma. Mitigating multithreats optimally in proactive threat management. ACM SIGSOFT Software Engineering Notes, 38(1):1–7, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhatia:2011:FCM

[BK11] Nitin Bhatia and Namarta Kapoor. Fuzzy cognitive map based approach for software quality risk analysis. ACM SIGSOFT Software Engineering Notes, 36(6):1–9, November 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Baliyan:2016:HFS

[BK16] Niyati Baliyan and Sandeep Kumar. A hierarchical fuzzy system for quality assessment of semantic Web application as a service. ACM SIGSOFT Software Engineering Notes, 41(1):1–7, January 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Babu:2012:IAD

[BKMJ12] P. Arun Babu, C. Senthil Kumar, N. Murali, and T. Jayakumar. An intuitive approach to determine test adequacy in safety-critical software. ACM SIGSOFT Software Engineering Notes, 37(5):1–10, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brooks:2016:CST

[BKP16] Andrew Brooks, Laura Krebs, and Brandon Paulsen. A comparison of sorting times between Java 8 and Parallel Colt: an exploratory experiment. ACM SIGSOFT Software Engineering Notes, 41(4):1–5, July 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bandyopadhyay:2010:ATD

[BL10] Anup Kumar Bandyopadhyay and Nibedita Lenka. Application of TLRO to dining philosophers problem. ACM SIGSOFT Software Engineering Notes, 35(6):1–6, November 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2010:BRM

[BM10a] Mordechai Ben-Menachem. Book review: Managing and Leading Software Projects by Richard Fairley and published by John Wiley & Sons, Inc. with IEEE Computer Society 2009, (hardback), ISBN 978-0470-29455-0, pp. 510. ACM SIGSOFT Software Engineering Notes, 35(4):34, July 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2010:PCV

[BM10b] Mordechai Ben-Menachem.

Book review: Parallel Coordinates: Visual Multidimensional Geometry and its Applications, by Alfred Inselberg, and published by Springer; 2009; ISBN 978-0-387-21507-5; pp. 580. ACM SIGSOFT Software Engineering Notes, 35(3):39, May 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Menachem:2010:BRR

[BM10c] Mordechai Ben-Menachem.

Book review: Reactive Systems: Modelling, Specification and Verification by L. Aceto, et al.; and published by Cambridge University Press; distributed by Cambridge University Press; 2007, (hardback), ISBN 978-0-521-87546-2, pp. 300. ACM SIGSOFT Software Engineering Notes, 35(4):34–35, July 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bowles:2012:SCT

[BM12] Juliana Bowles and Dulani Meedeniya. Strongly consistent transformation of partial scenarios. ACM SIGSOFT Software Engineering Notes, 37(4):1–8, July 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Benala:2018:SSD

[BM18] Tirimula Rao Benala and Rajib Mall. SEET: Software development effort estimation using ensemble techniques. ACM SIGSOFT Software Engineering Notes, 43(3):17, July 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Baresi:2012:LBS

[BMMR12] Luciano Baresi, Angelo Morzenti, Alfredo Motta, and Matteo Rossi. A logic-based semantics for the verification of multidiagram UML models. ACM SIGSOFT Software Engineering Notes, 37(4):1–8, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhat:2010:CVB

[BMRB10] Shubha Bhat, Vindhya Malagi, Krishnan Rangarajan, and Ramesh Babu. Computer vision based guidance in UAVs: software engineering challenges. ACM SIGSOFT Software Engineering Notes, 35 (6):1–6, November 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Banerjee:2010:RSS

[BP10] C. Banerjee and S. K. Pandey. Research on software security awareness: problems and prospects. ACM SIGSOFT Software Engineering Notes, 35(5):1–5, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Berliner:2010:EDC

[BR10] Brian Berliner and Nayan B.

Ruparelia. Early days of CVS. ACM SIGSOFT Software Engineering Notes, 35 (5):5–6, September 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhardwaj:2016:KSM

[BR16] Mridul Bhardwaj and Ajay Rana. Key software metrics and its impact on each other for software development projects. ACM SIG-SOFT Software Engineering Notes, 41(1):1–4, January 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Banerjee:2012:DAB

[BRD^{+}12] A. Banerjee, S. Ray, P. Dasgupta, P. P. Chakrabarti, S. Ramesh, P. Vignesh, and V. Ganesan. A dynamic assertion-based verification platform for validation of UML designs. ACM SIGSOFT Software Engineering Notes, 37(1):1–14, January 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

B:2013:SHE

[BRS13] Sugavanesh B., Hari Prasath R., and Selvakumar S. SHS-HTTPS enforcer: enforcing HTTPS and preventing MITM attacks. ACM SIG-SOFT Software Engineering Notes, 38(6):1–4, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Blech:2012:TFF

[BS12] Jan Olaf Blech and Bernhard Schätz. Towards a formal foundation of behavioral types for UML statemachines. ACM SIGSOFT Software Engineering Notes, 37(4):1–8, July 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhasin:2013:CGT

[BS13] Harsh Bhasin and Neha Singla. Cellular-genetic test data generation. ACM SIG-SOFT Software Engineering Notes, 38(5):1–9, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Banerjee:2017:RAF

[BS17] Shreya Banerjee and Anirban Sarkar. A requirements analysis framework for development of service oriented systems. ACM SIGSOFT Software Engineering Notes, 42 (3):1–12, July 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brahmasani:2012:PXA

[BSS12] Siva Brahmasani, Subramanian Selvakumar, and E. Sivasankar. Prevention of XSS attacks using STCD: Server side tagging and client side differentiation. ACM SIGSOFT Software Engineering Notes, 37(6):1–9, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Baliyan:2013:FMH

[BSS13a] Niyati Baliyan, Vidushi Sharma, and Shivani. A fuzzy model for high-level clones in software. ACM SIGSOFT Software Engineering Notes, 38(3):1–4, May 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bhasin:2013:CAB

[BSS13b] Harsh Bhasin, Neha Singla, and Shruti Sharma. Cellular automata based test data generation. ACM SIGSOFT Software Engineering Notes, 38(4):1–7, July 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brahmasani:2013:PXA

[BSS13c] Siva Brahmasani, Subramanian Selvakumar, and E. Sivasankar. Prevention of XSS attacks using STCD:

Server side tagging and client side differentiation. ACM SIGSOFT Software Engineering Notes, 38(1):46–49, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bultan:2018:SCA

[Bul18] Tevfik Bultan. Side-channel analysis via symbolic execution and model counting. ACM SIGSOFT Software Engineering Notes, 43 (4):55, October 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bures:2017:SES

[BWS^{+}17] Tomas Bures, Danny Weyns, Bradley Schmer, Eduardo Tovar, Eric Boden, Thomas Gabor, Ilias Gerostathopoulos, Pragya Gupta, Eunsuk Kang, Alessia Knauss, Pankesh Patel, Awais Rashid, Ivan Ruchkin, Roykrong Sukkerd, and Christos Tsigkanos. Software engineering for smart cyber-physical systems: Challenges and promising solutions. ACM SIGSOFT Software Engineering Notes, 42 (2):19–24, April 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bures:2019:SES

[BWS^{+}19] Tomas Bures, Danny Weyns, Bradley Schmerl, John Fitzgerald, Adina Aniculaesei, Christian Berger, João Cambeiro, Jan Carlson, Shafiul Azam Chowdhury, Marian Daun, Nianyu Li, Matthias Markthaler, Claudio Menghi, Birgit Penzenstadler, Aedan Pettit, Robert Pettit, Luca Sabatucci, Christos Tranoris, Hans Vangheluwe, Sebastian Voss, and Edith Zavala. Software engineering for smart cyber-physical systems (SEsCPS 2018) — workshop report. ACM SIG-SOFT Software Engineering Notes, 44(4):11–13, December 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3364452.3364465.

Bures:2018:SESa

[BWSF18a] Tomas Bures, Danny Weyns, Bradley Schmer, and John Fitzgerald. Software engineering for smart cyber-physical systems: Models, systemenvironment boundary, and social aspects. ACM SIG-SOFT Software Engineering Notes, 43(4):42–44, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bures:2018:SESb

[BWSF18b] Tomas Bures, Danny Weyns, Bradley Schmerl, and John Fitzgerald. Software engineering for smart cyber-physical systems: Models, systemenvironment boundary, and social aspects. ACM SIG-SOFT Software Engineering Notes, 43(4):54, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bu:2018:MBC

[BZC^{+}18] Lei Bu, Tian Zhang, Xin Chen, Linzhang Wang, Jianhua Zhao, and Xuandong Li. Model-based construction and verification of cyber-physical systems. ACM SIGSOFT Software Engineering Notes, 43(3):6–10, July 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Capretz:2010:WDW

[CA10] Luiz Fernando Capretz and Faheem Ahmed. Why do we need personality diversity in software engineering? ACM SIGSOFT Software Engineering Notes, 35(2):1–11, March 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cartaxo:2018:SKT

[Car18] Bruno Cartaxo. Supporting knowledge transfer from secondary studies to software engineering practice. ACM SIGSOFT Software Engineering Notes, 43(1):1–6, January 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Catal:2013:TEB

[Cat13]  Cagatay Catal.  Teaching

evidence-based software engineering to master students: a single lecture within a course or an entire semester-long course? ACM SIGSOFT Software Engineering Notes, 38 (2):1–2, March 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Caglayan:2012:IOA

[CB12] Bora Caglayan and Ayse Bener. Issue ownership activity in two large software projects. ACM SIGSOFT Software Engineering Notes, 37(6):1–7, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Campos:2010:IAC

[CBdRS10] Juliana Pinheiro Campos, José Luis Braga, Antânio Maria Pereira de Resende, and Carlos Henrique Osório Silva. Identification of aspect candidates by inspecting use cases descriptions. ACM SIGSOFT Software Engineering Notes, 35(4):1–9, July 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chan:2010:STR

[CBK10] Wing Kwong Chan, Christof J.

Budnik, and Gregory M. Kapfhammer. Software testing research in practice: report on the 5th International Workshop on the Automation of Software Test. ACM SIG-SOFT Software Engineering Notes, 35(5):25–26, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chaudhary:2013:ERA

[CC13] Ritika Chaudhary and Ram Chatterjee. Essence of reusability in aspect-oriented systems. ACM SIGSOFT Software Engineering Notes, 38(3):1–5, May 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Canfora:2010:IWE

[CCM^{+}10] Gerardo Canfora, Giulio Concas, Michele Marchesi, Ewan Tempero, and Hongyu Zhang. 2010 ICSE Workshop on Emerging Trends in Software Metrics. ACM SIGSOFT Software Engineering Notes, 35 (5):51–53, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Carrillo:2012:FVC

[CCM12] Oscar Carrillo, Samir Chouali, and Hassan Mountassir. Formalizing and verifying compatibility and consistency of SysML blocks. ACM SIG-SOFT Software Engineering Notes, 37(4):1–8, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chechik:2017:RWM

[CD17] Marsha Chechik and Davide Di Ruscio. Report from the 9th Workshop on Modelling in Software Engineering (MiSE 2017). ACM SIGSOFT Software Engineering Notes, 42 (4):21–24, October 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Choi:2019:SEI

[CEH19] Byoungju Choi, María-José Escalona, and Kim Herzig. Summary of the 14th edition of the IEEE/ ACM Workshop on Automation of Software Test (AST). ACM SIGSOFT Software Engineering Notes, 44(3):53, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3356773.3356808.

Chapelle:2013:BRP

[Cha13a] Gregory Chapelle. Book review: A practical guide to Linux commands, editors, and shell-programming, third edition by Mark G. Sobell. ACM SIGSOFT Software Engineering Notes, 38(4):38, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chapelle:2013:BRO

[Cha13b] Gregory Chapelle. Book review: The official Ubuntu book, seventh edition by Matthew Helmke and Amber Graner. ACM SIGSOFT Software Engineering Notes, 38 (1):54–55, January 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ciccozzi:2019:RIW

[CHMW19] Federico Ciccozzi, Nico Hochgeschwender, Ivano Malavolta, and Andreas Wortmann. Report on the 2nd International Workshop on Robotics Software Engineering (RoSE’19). ACM SIGSOFT Software Engineering Notes, 44(3):38– 40, November 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3356773.3356804.

Chodkowski:2010:RIA

[Cho10] Steven Chodkowski. Recommendations for the information architect’s book shelf. ACM SIGSOFT Software Engineering Notes, 35 (2):1–5, March 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chatterjee:2010:PAA

[CJ10] Ram Chatterjee and Kalpana Johari. A prolific approach for automated generation of test cases from informal requirements. ACM SIGSOFT Software Engineering Notes, 35(5):1–11, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chandra:2011:AST

[CK11a] Shalini Chandra and Raees Ahmad Khan. Availability state transition model. ACM SIG-SOFT Software Engineering Notes, 36(3):1–3, May 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Choudhary:2011:TSF

[CK11b] R. K. Choudhary and R. A.

Khan. Testing software fault tolerance techniques: future direction. ACM SIGSOFT Software Engineering Notes, 36(3):1–5, May 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cordeiro:2018:BJV

[CKS18] Lucas C. Cordeiro, Daniel Kroening, and Peter Schrammel. Benchmarking of Java verification tools at the Software Verification Competition (SV-COMP). ACM SIGSOFT Software Engineering Notes, 43(4):56, October 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Couto:2013:CDE

[CMGV13] Cesar Couto, Cristiano Maffort, Rogel Garcia, and Marco Tulio Valente. COMETS: a dataset for empirical research on software evolution using source code metrics and time series analysis. ACM SIGSOFT Software Engineering Notes, 38(1):1–3, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chhillar:2011:EAO

[CN11] Rajender Singh Chhillar and Nisha. Empirical analysis of object-oriented design metrics for predicting high, medium and low severity faults using Mallows Cp. ACM SIG-SOFT Software Engineering Notes, 36(6):1–9, November 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cooper:2012:BRD

[Coo12]  Greg Cooper.  Book re-

view: DTrace: dynamic tracing in Oracle Solaris, Mac OS X, and FreeBSD by Brendan Gregg and Jim Mauro. ACM SIGSOFT Software Engineering Notes, 37(1):34, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Colombo:2012:PSS

[CPG^{+}12] Regina Thienne Colombo, Marcelo Schneck Pessôa, Ana Cervigni Guerra, Amandio Balcão Filho, and Célio Caruso Gomes. Prioritization of software security intangible attributes. ACM SIGSOFT Software Engineering Notes, 37(6):1–7, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cheluvaraju:2012:QMP

[CPPC12] Bharath Cheluvaraju, Anjaneyulu Pasala, Srinivas Padmanabhuni, and Sadhana Chevireddy. A quantitative measure for preventive maintenance in software. ACM SIGSOFT Software Engineering Notes, 37(4):1–5, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ceccarello:2012:TGC

[CS12a] Matteo Ceccarello and Nastaran Shafiei. Tools to generate and check consistency of model classes for Java PathFinder. ACM SIGSOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Counsell:2012:IAR

[CS12b] Steve Counsell and Stephen Swift. Issues arising from refactoring studies: an experience report. ACM SIGSOFT Software Engineering Notes, 37(3):1–5, May 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chauhan:2013:DSM

[CSG13] Sandeep Chauhan, Arun Sharma, and P. S. Grover. Developing self managing software systems using agile modeling. ACM SIGSOFT Software Engineering Notes, 38 (6):1–3, November 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chanda:2011:SGA

[CSKB11] Jayeeta Chanda, Sabnam Sengupta, Ananya Kanjilal, and Swapan Bhattacharya. SCAG: a graphical approach to measure the complexity of the SOA application. ACM SIG-SOFT Software Engineering Notes, 36(5):1–6, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chanda:2012:TBS

[CSKB12] Jayeeta Chanda, Sabnam Sengupta, Ananya Kanjilal, and Swapan Bhattacharya. Traceability between service component and class: a model based approach. ACM SIG-SOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chanda:2013:BSE

[CSKB13] Jayeeta Chanda, Sabnam Sengupta, Ananya Kanjilal, and Swapan Bhattacharya. Behavioral and structural evolution of SOA from OO: an integrated approach. ACM SIGSOFT Software Engineering Notes, 38(5):1–9, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Calefato:2019:SIC

[CTD19] Fabio Calefato, Paolo Tell, and Alpana Dubey. Summary of the 14th International Conference on Global Software Engineering (ICGSE). ACM SIGSOFT Software Engineering Notes, 44(3):30– 33, November 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3356773.3356802.

Castelluccia:2013:TEB

[CV13] Daniela Castelluccia and Giuseppe Visaggio. Teaching evidence-based software engineering: learning by a collaborative mapping study of open source software. ACM SIGSOFT Software Engineering Notes, 38(6):1–4, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chen:2019:SVR

[CYWD19] Zhenbang Chen, Hengbiao Yu, Ji Wang, and Wei Dong. Symbolic verification of regular properties for Java programs. ACM SIGSOFT Software Engineering Notes, 44(4):17, December 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3364452.33644561.

Dahiya:2010:TSE

[Dah10] Deepak Dahiya. Teaching software engineering: a practical approach. ACM SIGSOFT Software Engineering Notes, 35(2):1–5, March 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DeSousaCoelho:2013:SDM

[DBA13] Jailton De Sousa Coelho, Jr., José Luis Braga, and Bernardo Giori Ambrósio. System dynamics model for simulation of the software inspection process. ACM SIG-SOFT Software Engineering Notes, 38(5):1–8, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dutta:2013:ERD

[DBK^{+}13] Animesh Dutta, Shrutilipi Bhattacharjee, Ananya Kanjilal, Ranjan Dasgupta, and Swapan Bhattacharya. Engineering of requirements for a distributed teleteaching system: a conceptual graphbased approach. ACM SIG-SOFT Software Engineering Notes, 38(6):1–12, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dalal:2013:ESR

[DC13] Sandeep Dalal and Rajender Singh Chhillar. Empirical study of root cause analysis of software failure. ACM SIG-SOFT Software Engineering Notes, 38(4):1–7, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

deCastro:2013:SGP

[dCBS13] Ronney Moreira de Castro, José Luis Braga, and Liziane Santos Soares. Selection of good practices for small software development teams: a knowledge-based approach. ACM SIGSOFT Software Engineering Notes, 38 (6):1–15, November 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Machado:2012:STP

[dCMMdA12] Ivan do Carmo Machado, John D. McGregor, and Eduardo Santana de Almeida. Strategies for testing products in software product lines. ACM SIGSOFT Software Engineering Notes, 37 (6):1–8, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dave:2011:CRM

[DD11] Vachik S. Dave and Kamlesh Dutta. Comparison of regression model, feed-forward neural network and radial basis neural network for software development effort estimation. ACM SIGSOFT Software Engineering Notes, 36 (5):1–5, September 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dekhane:2010:IAT

[Dek10] Sonal Dekhane. Install anywhere tutorial and reference guide by zero g team. ACM SIGSOFT Software Engineering Notes, 35(5):57, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DelRa:2011:BRU

[Del11a] William Del Ra. Book review: Usability testing essentials: ready, set… test! by Carol M. Barnum. ACM SIG-SOFT Software Engineering Notes, 36(5):49–50, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DelRa:2011:BRB

[Del11b] William Del Ra III. Book review: Brave NUI world: designing natural user interfaces for touch and gesture by Daniel Wigdor and Dennis Wixon. ACM SIGSOFT Software Engineering Notes, 36(6):29–30, November 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DelRa:2011:MBD

[Del11c] William Del Ra III. Book review: Model-based development: applications, by H. S. Lahman. ACM SIGSOFT Software Engineering Notes, 36(6):29, November 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DelRa:2012:BRA

[Del12a] William Del Ra III. Book review: ACM Ruby learning path by David A. Black. ACM SIGSOFT Software Engineering Notes, 37(3):38, May 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DelRa:2012:BRSb

[Del12b] William Del Ra III. Book review: Service design patterns: fundamental design solutions for SOAP/WSDL and RESTful web services by Robert Daigneau. ACM SIGSOFT Software Engineering Notes, 37(4):40, July 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DelRa:2012:BRSa

[Del12c] William Del Ra III. Book review: Software systems architecture, second edition, by Nick Rozanski and Eoin Woods. ACM SIGSOFT Software Engineering Notes, 37 (2):36, March 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DelRa:2012:BRB

[Del12d] William Del Ra III. Book review: The book of Ruby by Huw Collingbourne. ACM SIGSOFT Software Engineering Notes, 37(1):37, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DelRa:2013:BRJ

[Del13] William Del Ra III. Book review: Java application architecture: modularity patterns with examples using OSGi by Kirk Knoernschild. ACM SIG-SOFT Software Engineering Notes, 38(1):55, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Duarte:2017:IWC

[DJB17] Carlos Henrique C. Duarte, Andreas Jedlitschka, and Ayse Bener. 4th International Workshop on Conducting Empirical Studies in Industry (CESI 2016): Post-workshop report. ACM SIGSOFT Software Engineering Notes, 42 (2):15–18, April 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doss:2016:COA

[DK16] O. Doss and T. P. Kelly. Challenges and opportunities in agile development in safety critical systems: a survey. ACM SIGSOFT Software Engineering Notes, 41(2):30–31, March 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dingsoyr:2013:RCL

[DM13] Torgeir Dingsøyr and Nils Brede Moe. Research challenges in large-scale agile software development. ACM SIGSOFT Software Engineering Notes, 38(5):38–39, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2010:SNSa

[Doe10a] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 35 (2):5–14, March 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2010:SNSb

[Doe10b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 35 (3):14–23, May 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2010:SNSc

[Doe10c] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 35 (4):8–16, July 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2010:SNSd

[Doe10d] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 35(5):8–17, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2010:SNSe

[Doe10e] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 35 (6):9–18, November 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2011:SNSa

[Doe11a] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 36 (2):10–18, March 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2011:SNSb

[Doe11b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 36 (3):9–18, May 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2011:SNSc

[Doe11c] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 36(5):10–19, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2011:SNSd

[Doe11d] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 36 (6):9–18, November 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2012:SNSa

[Doe12a] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 37 (1):12–20, January 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2012:SNSb

[Doe12b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 37 (2):11–20, March 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2012:SNSc

[Doe12c] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 37 (3):10–17, May 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2012:SNSd

[Doe12d] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 37 (4):11–19, July 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2012:SNSe

[Doe12e] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 37(5):17–26, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2012:SNSf

[Doe12f] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 37(6):10–18, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2013:SNSa

[Doe13a] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 38 (1):10–19, January 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2013:SNSb

[Doe13b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 38 (2):11–19, March 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2013:SNSc

[Doe13c] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 38 (3):11–20, May 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2013:SNSd

[Doe13d] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 38 (4):10–18, July 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2013:SNSe

[Doe13e] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 38(5):19–27, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2013:SNSf

[Doe13f] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 38(6):11–20, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2014:SNS

[Doe14] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 39 (6):8–16, November 2014. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2016:SNSa

[Doe16a] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 41 (1):9–17, January 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2016:SNSb

[Doe16b] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 41 (2):6–15, March 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2016:SNSc

[Doe16c] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 41 (4):9–17, July 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2016:SNSd

[Doe16d] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIG-SOFT Software Engineering Notes, 41(5):11–18, September 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Doernhoefer:2018:SNS

[Doe18] Mark Doernhoefer. Surfing the net for Software Engineering Notes. ACM SIGSOFT Software Engineering Notes, 43 (2):17–25, April 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dubey:2010:AUM

[DR10] Sanjay Kumar Dubey and Ajay Rana. Assessment of usability metrics for objectoriented software system. ACM SIGSOFT Software Engineering Notes, 35(6):1– 4, November 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dubey:2011:AMM

[DR11a] Sanjay Kumar Dubey and Ajay Rana. Assessment of maintainability metrics for object-oriented software system. ACM SIGSOFT Software Engineering Notes, 36 (5):1–7, September 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dubey:2011:UES

[DR11b] Sanjay Kumar Dubey and Ajay Rana. Usability estimation of software system by using object-oriented metrics. ACM SIGSOFT Software Engineering Notes, 36 (2):1–6, March 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dwivedi:2018:TAN

[DR18] Ashish Kumar Dwivedi and Santanu Kumar Rath. Transformation of alloy notation into a semantic notation. ACM SIGSOFT Software Engineering Notes, 43 (1):1–6, January 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dubey:2012:MPO

[DRD12] Sanjay Kumar Dubey, Ajay Rana, and Yajnaseni Dash. Maintainability prediction of object-oriented software system by multilayer perceptron model. ACM SIGSOFT Software Engineering Notes, 37 (5):1–4, September 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Drori:2016:TSD

[Dro16] Offer Drori. Template for a system design file using OODPM version 2015. ACM SIGSOFT Software Engineering Notes, 41(1):1–4, January 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dabaghchian:2017:CAS

[DRO^{+}17] Maryam Dabaghchian, Zvonimir Rakamaric, Burcu K. Ozkan, Erdal Mutlu, and Serdar Tasiran. Consistencyaware scheduling for weakly consistent programs. ACM SIGSOFT Software Engineering Notes, 42(4):1–5, October 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

deSousa:2010:AAR

[dSAVP10] Thiago C. de Sousa, Jorge R. Almeida, Jr., Sidney Viana, and Judith Pavón. Automatic analysis of requirements consistency with the B method. ACM SIGSOFT Software Engineering Notes, 35 (2):1–4, March 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dittrich:2010:CHA

[DST^{+}10] Yvonne Dittrich, Helen Sharp, Heike Winshiers Theophilus, Cleidson De Souza, Mikko Korpela, and Janice Singer. Cooperative and human aspects of software engineering: CHASE 2010. ACM SIG-SOFT Software Engineering Notes, 35(5):27–29, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fazal-e-Amin:2010:PES

[eAMO10] Fazal e Amin, Ahmad Kamil Mahmood, and Alan Oxley. Proposal for evaluation of software reusability assessment approach employing a mixed method. ACM SIGSOFT Software Engineering Notes, 35 (5):1–4, September 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ebert:2013:BRS

[Ebe13] John Ebert. Book review: SOA with REST: principles, patterns & constraints for building enterprise solutions with REST by Thomas Erl, Benjamin Carlyle, Cesare Pautasso, Raj Balasubramanian. ACM SIGSOFT Software Engineering Notes, 38 (3):32–33, May 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

El-Deeb:2012:VSS

[ED12] Ahmed El-Deeb. A view on the state of software engineering practice: experience from Egypt. ACM SIGSOFT Software Engineering Notes, 37 (6):1–4, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Eck:2019:PMCa

[EH19a] Stefan Eck and Hans A. Hansson. PROMPT — master courses for professional software developers. ACM SIG-SOFT Software Engineering Notes, 44(1):29–30, January 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Eck:2019:PMCb

[EH19b] Stefan Eck and Hans A. Hansson. PROMPT — master courses for professional software developers. ACM SIG-SOFT Software Engineering Notes, 44(1):35, January 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Eisenberg:2012:MSD

[Eis12a] Robert Eisenberg. Managing software debt building for inevitable change by chris sterling. ACM SIGSOFT Software Engineering Notes, 37 (2):36–37, March 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Eisenberg:2012:TBA

[Eis12b] Robert J. Eisenberg. A threshold based approach to technical debt. ACM SIGSOFT Software Engineering Notes, 37(2):1–6, March 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Elbaum:2016:SI

[Elb16] Sebastian Elbaum. The state of ICSE. ACM SIGSOFT Software Engineering Notes, 41(5):9–10, September 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Elbaum:2017:SI

[Elb17] Sebastian Elbaum. The state of ICSE. ACM SIGSOFT Software Engineering Notes, 42(2):4–5, April 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Elbaum:2019:SI

[Elb19] Sebastian Elbaum. The state of ICSE. ACM SIGSOFT Software Engineering Notes, 44(2):4–5, April 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Exman:2016:SPG

[EPBR16] Iaakov Exman, Dewayne E.

Perry, Balbir Barn, and Paul Ralph. Separability principles for a general theory of software engineering: Report on the GTSE 2015 workshop. ACM SIGSOFT Software Engineering Notes, 41(1):25–27, January 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Epps:2011:BRE

[Epp11] Bob Epps. Book review: The elements of MATLAB style by Richard K. Johnson. ACM SIGSOFT Software Engineering Notes, 36(3):33–34, May 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fontana:2017:TDA

[FCT^{+}17] Francesca Arcelli Fontana, Alexander Chatzigeorgiou, Wolfgang Trumler, Clemente Izurieta, Paris Avgeriou, and Robert L. Nord. Technical debt in agile development: Report on the Ninth Workshop on Managing Technical Debt (MTD 2017). ACM SIGSOFT Software Engineering Notes, 42(3):18–21, July 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fraser:2018:ACPb

[FM18a] Steven Fraser and Dennis Manci. Agile culture: a panels report from XP 2017. ACM SIGSOFT Software Engineering Notes, 43(4):52, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fraser:2018:NSBb

[FM18b] Steven Fraser and Dennis Manci. No silver bullet reloaded: Report on XP 2017 panel session. ACM SIGSOFT Software Engineering Notes, 43(4):53, October 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fraser:2018:ACPa

[FM18c] Steven Fraser and Dennis Mancl. Agile culture: a panels report from XP 2017. ACM SIGSOFT Software Engineering Notes, 43(4):21–23, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fraser:2018:NSBa

[FM18d] Steven Fraser and Dennis Mancl. No silver bullet reloaded: Report on XP 2017 panel session. ACM SIG-SOFT Software Engineering Notes, 43(4):39–41, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Franky:2011:AMD

[Fra11] María Consuelo Franky. Agile management and development of software projects based on collaborative environments. ACM SIGSOFT Software Engineering Notes, 36(3):1–6, May 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fraser:2016:SER

[Fra16] Steven Fraser. Software Engineering Research and Industrial Practice IEEE ICSE 2015 Workshop Report: May 17, 2015. ACM SIGSOFT Software Engineering Notes, 41 (1):28–31, January 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Froberg:2012:BRP

[Fro12a] Scott Froberg. Book review:

Programming perl, fourth edition by Tom Christiansen, Brian D. Foy and Larry Wall with Jon Orwant. ACM SIG-SOFT Software Engineering Notes, 37(4):40, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Froberg:2012:BRS

[Fro12b]  Scott Froberg.  Book re-

view: Software Testing by Yogesh Singh. ACM SIGSOFT Software Engineering Notes, 37(3):36, May 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Froberg:2013:BRD

[Fro13a] Scott Froberg. Book review: Distributed and cloud computing from parallel processing to the Internet of Things by Kai Hwang, Geoffry C. Fox, and Jack J. Dongarra. ACM SIG-SOFT Software Engineering Notes, 38(2):34, March 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Froberg:2013:BRP

[Fro13b] Scott Froberg. Book review: Private cloud computing: consolidation, virtualization, and service-oriented infrastructure by Stephen R. Smoot, Nam K. Tan. ACM SIGSOFT Software Engineering Notes, 38 (2):35, March 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fricker:2011:IRE

[FS11] Samuel Fricker and Norbert Seyff. 1st International Requirements Engineering Efficiency Workshop: REEW 2011. ACM SIGSOFT Software Engineering Notes, 36 (3):26–28, May 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fowler:2018:BUP

[FS18] Megan Fowler and Tim Schwab. BeginToReason: Understanding the purpose of code. ACM SIGSOFT Software Engineering Notes, 43 (3):18, July 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Funes:2012:RMC

[FSK12] Diego Funes, Junaid Haroon Siddiqui, and Sarfraz Khurshid. Ranged model checking. ACM SIGSOFT Software Engineering Notes, 37 (6):1–5, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Farrell-Vinary:2011:J

[FV11] P. Farrell-Vinary. JustInMind.

ACM SIGSOFT Software Engineering Notes, 36(3):34–35, May 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Galster:2011:VSA

[GAWM11] Matthias Galster, Paris Avgeriou, Danny Weyns, and Tomi Männistö. Variability in software architecture: current practice and challenges. ACM SIGSOFT Software Engineering Notes, 36(5):30– 32, September 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gupta:2010:TFR

[GB10] Amit Gupta and Rajesh Bhatia. Testing functional requirements using b model specifications. ACM SIGSOFT Software Engineering Notes, 35 (2):1–7, March 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gandhi:2011:EGR

[GB11] Parul Gandhi and Pradeep Kumar Bhatia. Estimation of generic reusability for objectoriented software an empirical approach. ACM SIGSOFT Software Engineering Notes, 36(3):1–4, May 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Goel:2013:ARO

[GB13a] Brij Mohan Goel and Pradeep Kumar Bhatia. Analysis of reusability of object-oriented systems using object-oriented metrics. ACM SIGSOFT Software Engineering Notes, 38(4):1–5, July 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Goel:2013:IHL

[GB13b] Brij Mohan Goel and Pradeep Kumar Bhatia. Investigating of high and low impact faults in object-oriented projects. ACM SIGSOFT Software Engineering Notes, 38(6):1–6, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gulati:2016:SRB

[GBSL16] Jayati Gulati, Priya Bhardwaj, Bharti Suri, and Anu Singh Lather. A study of relationship between performance, temperament and personality of a software programmer. ACM SIGSOFT Software Engineering Notes, 41 (1):1–5, January 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gulia:2012:NAG

[GC12] Preeti Gulia and R. S. Chillar.

A new approach to generate and optimize test cases for UML state diagram using genetic algorithm. ACM SIG-SOFT Software Engineering Notes, 37(3):1–5, May 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gama:2010:SAA

[GD10] Kiev Gama and Didier Donsez. A survey on approaches for addressing dependability attributes in the OSGi service platform. ACM SIGSOFT Software Engineering Notes, 35(3):1–8, May 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gupta:2012:MCS

[GD12] Daya Gupta and Rinky Dwivedi. Method configuration from situational method engineering. ACM SIGSOFT Software Engineering Notes, 37(3):1–11, May 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garg:2013:NBG

[GDF13] Deepak Garg, Amitava Datta, and Tim French. A novel bipartite graph approach for selection and prioritisation of test cases. ACM SIGSOFT Software Engineering Notes, 38(6):1–6, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Georgieva:2010:CFS

[Geo10] Konstantina Georgieva. Conducting FMEA over the software development process. ACM SIGSOFT Software Engineering Notes, 35(3):1–5, May 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gorschek:2010:TIW

[GFBE10] Tony Gorschek, Samuel Fricker, Sjaak Brinkkemper, and Christof Ebert. Third International Workshop on Software Product Management — IWSPM’09. ACM SIGSOFT Software Engineering Notes, 35(2):25–29, March 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Geihs:2010:RAI

[GGR10] Kurt Geihs, Stefan Gruner, and Kay Römer. Report about 1st ICSE Workshop on Software Engineering for Sensor Network Applications (SESENA 2010). ACM SIG-SOFT Software Engineering Notes, 35(5):34–37, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Graziotin:2013:RDP

[GJ13] Daniel Graziotin and Andreas Jedlitschka. Recent developments in product-focused software process improvement: PROFES 2013 conference report. ACM SIGSOFT Software Engineering Notes, 38 (6):29–34, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Garbervetsky:2012:RIW

[GK12] Diego Garbervetsky and Sunghun Kim. Report from 2nd International Workshop on Developing Tools as Plug-Ins (TOPI 2012). ACM SIG-SOFT Software Engineering Notes, 37(6):24–27, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Geetha:2011:FHP

[GKK11] D. Evangelin Geetha, T. V. Suresh Kumar, and K. Rajani Kanth. Framework for hybrid performance prediction process model: use case performance engineering approach. ACM SIGSOFT Software Engineering Notes, 36(3):1–15, May 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gerostathopoulos:2019:CDD

[GKK^{+}19] Ilias Gerostathopoulos, Marco Konersmann, Stephan Krusche, David I. Mattos, Jan Bosch, Tomas Bures, Brian Fitzgerald, Michael Goedicke, Henry Muccini, Helena H. Olsson, Thomas Brand, Robert Chatley, Nikolaos Diamantopoulos, Arik Friedman, Miguel Jiménez, Jan Ole Johanssen, Putra Manggala, Masumi Koseki, Jorge Melegati, Nuthan Munaiah, Gabriel Tamura, Vasileios Theodorou, Jeffrey Wong, and Iris Figalist. Continuous data-driven software engineering — towards a research agenda: Report on the Joint 5th International Workshop on Rapid Continuous Software Engineering (RCoSE 2019) and 1st International Works. ACM SIG-SOFT Software Engineering Notes, 44(3):60–64, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3356773.3356811.

Gleirscher:2018:SRIa

[GKL18a] Mario Gleirscher, Stefan Kugele, and Sven Linker. SCAV’18: Report of the 2nd International Workshop on Safe Control of Autonomous Vehicles. ACM SIGSOFT Software Engineering Notes, 43(4):45–47, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gleirscher:2018:SRIb

[GKL18b] Mario Gleirscher, Stefan Kugele, and Sven Linker. SCAV’18: Report of the 2nd International Workshop on Safe Control of Autonomous Vehicles. ACM SIGSOFT Software Engineering Notes, 43(4):54, October 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gerard:2018:RIW

[GKMM18] Sébastien Gérard, Dimitrios S.

Kolovos, Ivano Malavolta, and Henry Muccini. Report from the 1st International Workshop on Collaborative Modelling in MDE (COMMitMDE 2016). ACM SIGSOFT Software Engineering Notes, 43 (2):26–27, April 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gleirscher:2017:SCA

[GKS17] Mario Gleirscher, Stefan Kugele, and Jonathan Sprinkle. Safe control of autonomous & connected vehicles (SCAV’17): Report from the 1st International Workshop at CPSWeek 2017. ACM SIGSOFT Software Engineering Notes, 42(3):22–23, July 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Glass:2018:ECH

[GL18] Robert L. Glass and Frank Land. Errors in computing history. ACM SIGSOFT Software Engineering Notes, 43 (4):7–8, October 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Glaves:2011:BRA

[Gla11]  Leslie Glaves.  Book re-

view: API design for C++ by Martin Reddy. ACM SIGSOFT Software Engineering Notes, 36(5):50, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Glass:2012:BRM

[Gla12] Robert L. Glass. Book review: Managing projects in trouble: achieving turnaround and success by Ralph L. Kliem. ACM SIGSOFT Software Engineering Notes, 37(1):34, January 2012. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Galster:2013:VSE

[GMCH^{+}13] Matthias Galster, Mehdi Mirakhorli, Jane Cleland-Huang, Janet E. Burge, Xavier Franch, Roshanak Roshandel, and Paris Avgeriou. Views on software engineering from the twin peaks of requirements and architecture. ACM SIG-SOFT Software Engineering Notes, 38(5):40–42, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gousios:2012:BRS

[Gou12] Georgios Gousios. Book review: Scalability rules 50 principles for scaling web sites by Martin L. Abbott and Michael T. Fisher. ACM SIGSOFT Software Engineering Notes, 37(1):35, January 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gibson:2012:IUF

[GP12] J. Paul Gibson and Isabelle Perseil. Introduction to UML and formal methods. ACM SIGSOFT Software Engineering Notes, 37(4):32–33, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ge:2012:TPD

[GPC12] Ning Ge, Marc Pantel, and Xavier Crégut. Time properties dedicated transformation from UML-MARTE activity to time transition system. ACM SIGSOFT Software Engineering Notes, 37(4):1–8, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gotz:2017:RIW

[GPW17] Sebastian Götz, Christian Piechnick, and Andreas Wortmann. Report on the 4th International Workshop on Model-driven Robot Software Engineering (MORSE). ACM SIGSOFT Software Engineering Notes, 42(4):32–34, October 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gruner:2012:FWF

[GR12] Stefan Gruner and Bernhard Rumpe. FormSERA Workshop on Formal Methods in Software Engineering Rigorous and Agile Approaches: 2nd of June 2012 at ICSE’2012 in Zürich (CH). ACM SIG-SOFT Software Engineering Notes, 37(6):28–30, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Grechanik:2012:CDD

[Gre12a] Mark Grechanik. The curse of database deadlocks: the problem with no good solution. ACM SIGSOFT Software Engineering Notes, 37 (5):13–14, September 2012.

CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Grechanik:2012:RBA

[Gre12b] Mark Grechanik. Random benchmark application generation for evaluating program analysis and testing tools. ACM SIGSOFT Software Engineering Notes, 37 (6):6–7, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2013:Pa

[Gro13a]  Alex Groce.  Passages.

ACM SIGSOFT Software Engineering Notes, 38(5):17– 18, September 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2013:Pb

[Gro13b] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 38(6):6–7, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2014:P

[Gro14] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 39(6):6–7, November 2014. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2016:Pa

[Gro16a] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 41(1):5, January 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2016:Pb

[Gro16b] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 41(2):5, March 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2016:Pc

[Gro16c] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 41(4):6–7, July 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2016:Pd

[Gro16d] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 41(5):6–5, September 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2017:Pa

[Gro17a] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 42(2):5–6, April 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2017:Pb

[Gro17b] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 42(3):9, July 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2017:Pc

[Gro17c] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 42(4):5–6, October 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2018:Pa

[Gro18a] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 43(1):1–2, January 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2018:Pb

[Gro18b] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 43(2):6–7, April 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2018:Pc

[Gro18c] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 43(3):5, July 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2018:Pd

[Gro18d] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 43(4):6–7, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2019:Pa

[Gro19a] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 44(1):7, January 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2019:Pb

[Gro19b] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 44(2):3–4, April 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2019:P

[Gro19c]  Alex Groce.  Passages.

ACM SIGSOFT Software Engineering Notes, 44(3):10– 11, November 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3356773.3356775.

Gill:2010:NCM

[GS10] Nasib S. Gill and Sunil Sikka. New complexity model for classes in object oriented system. ACM SIGSOFT Software Engineering Notes, 35 (5):1–7, September 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gaur:2012:AIM

[GS12] Vibha Gaur and Anuja Soni.

Analytical inference model for prediction and customization of inter-agent dependency requirements. ACM SIGSOFT Software Engineering Notes, 37(2):1–11, March 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gandotra:2011:LSA

[GSB11] Vandana Gandotra, Archana Archana Singhal, and Punam Bedi. Layered security architecture for threat management using multi-agent system. ACM SIGSOFT Software Engineering Notes, 36(5):1– 11, September 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gill:2010:MDP

[GT10] Nasib Singh Gill and Pradeep Tomar. Modified development process of component-based software engineering. ACM SIGSOFT Software Engineering Notes, 35(2):1–6, March 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Galster:2017:TUS

[GTK17] Matthias Galster, Damian A.

Tamburri, and Rick Kazman.

Towards understanding the social and organizational dimensions of software architecting. ACM SIGSOFT Software Engineering Notes, 42(3):24–25, July 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gupta:2011:VDC

[Gup11]  Varun Gupta.  Validation

of dynamic coupling metrics for object-oriented software. ACM SIGSOFT Software Engineering Notes, 36 (5):1–3, September 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gvero:2013:BRC

[Gve13a] Igor Gvero. Book review: Core Java volume I: fundamentals, 9th edition by Cay S. Horstmann and Gary Cornell. ACM SIGSOFT Software Engineering Notes, 38(3):33, May 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gvero:2013:BRO

[Gve13b] Igor Gvero. Book review: Observing the user experience, 2nd edition: a practitioner’s guide to user research by Elizabeth Goodman, Mike Kuniavsky, and Andrea Moed. ACM SIGSOFT Software Engineering Notes, 38 (2):35, March 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gvero:2013:CCE

[Gve13c] Igor Gvero. Computers as components, 3rd edition: principles of embedded computing system design by marilyn Wolf. ACM SIGSOFT Software Engineering Notes, 38(5):67–68, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Galster:2017:VCS

[GWG^{+}17] Matthias Galster, Danny Weyns, Michael Goedicke, Uwe Zdun, Jácome Cunha, and Jaime Chavarriaga. Variability and complexity in software design: Towards quality through modeling and testing. ACM SIGSOFT Software Engineering Notes, 42 (4):35–37, October 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hagar:2011:BRT

[Hag11] Jon D. Hagar. Book review: Testing IT: an offthe-shelf software testing process by John Watkins and Simon Mills. ACM SIGSOFT Software Engineering Notes, 36(3):33, May 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hunt:2013:RPS

[HAJW13] Johanna Hunt, Olumide Akerele, Tomi Juhola, and Michael Waterman. Report from the PhD symposium at XP2013: an adaptive experiment in feedback. ACM SIG-SOFT Software Engineering Notes, 38(5):59–62, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Haller:2013:MT

[Hal13] Klaus Haller. Mobile testing. ACM SIGSOFT Software Engineering Notes, 38 (6):1–8, November 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hooker:2019:JHT

[HAM^{+}19] Joshua Hooker, Peter Aldous, Eric Mercer, Benjamin Ogles, Kyle Storey, and S. Jacob Powell. JPF-HJ: a tool for task parallel program analysis. ACM SIG-SOFT Software Engineering Notes, 44(4):19, December 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3364452.33644563.

Hathhorn:2012:BRE

[Hat12] Chris Hathhorn. Book review: Engineering a compiler, second edition by Keith D. Cooper and Linda Torczon. ACM SIGSOFT Software Engineering Notes, 37 (1):36–37, January 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hopfner:2010:EAD

[HB10] Hagen Höpfner and Christian Bunse. Energy aware data management on AVR micro controller based systems. ACM SIGSOFT Software Engineering Notes, 35 (3):1–8, May 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hasteer:2013:PAR

[HBM13] Nitasha Hasteer, Abhay Bansal, and B. K. Murthy. Pragmatic assessment of research intensive areas in cloud: a systematic review. ACM SIGSOFT Software Engineering Notes, 38(3):1–6, May 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harrison:2012:RFI

[HdCH^{+}12] Rachel Harrison, Daniela da Cruz, Pedro Henriques, Maria João Varanda Pereira, Shih-Hsi Liu, Tim Menzies, Marjan Mernik, and Daniel Rodriguez. Report from the First International Workshop on Realizing Artificial Intelligence Synergies in Software Engineering (RAISE 2012). ACM SIGSOFT Software Engineering Notes, 37(5):34– 35, September 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Herrmann:2012:IWC

[HDDS12] Andrea Herrmann, Maya Daneva, Joerg Doerr, and Kurt Schneider. 2nd International Workshop on Creativity in Requirements Engineering: CreaRE 2012. ACM SIGSOFT Software Engineering Notes, 37(4):30–31, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hazra:2013:CAR

[HDKB13] Rumpa Hazra, Shouvik Dey, Ananya Kanjilal, and Swapan Bhattacharya. Comparative analysis of real time resource access control protocols using UML 2.0. ACM SIGSOFT Software Engineering Notes, 38(5):1–7, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hu:2016:UUC

[HJ16] Wei-Chung Hu and Hewijin Christine Jiau. UCFrame: a use case framework for crowd-centric requirement acquisition. ACM SIGSOFT Software Engineering Notes, 41(2):1–13, March 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harikrishnan:2012:SEN

[HK12] S. Harikrishnan and Rajeev Kumar. Space efficient nonconstant time multi-method dispatch in object oriented systems. ACM SIGSOFT Software Engineering Notes, 37(2):1–6, March 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Herzwurm:2012:RIW

[HKPS12] Georg Herzwurm, Benedikt Krams, Wolfram Pietsch, and Sixten Schockert. Report from the 3rd International Workshop on Requirements Prioritization for Customer Oriented Software Development: (RePriCo’12). ACM SIG-SOFT Software Engineering Notes, 37(4):32–33, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Haider:2018:AAD

[HMB18] Umaima Haider, John D. Mc-Gregor, and Rabih Bashroush. The ALI Architecture Description Language. ACM SIG-SOFT Software Engineering Notes, 43(4):52, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hanssen:2016:PWA

[HMS16] Geir K. Hanssen, Thor Myklebust, and Tor Stålhane. Proceedings on the 1st Workshop on Agile Methods Applied to Development and Certification of Safety-critical Software. ACM SIGSOFT Software Engineering Notes, 41 (2):32–33, March 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Honig:2016:LAS

[HNT16] William L. Honig, Natsuko Noda, and Shingo Takada. Lack of attention to singular (or atomic) requirements despite benefits for quality, metrics and management. ACM SIGSOFT Software Engineering Notes, 41(4):1–5, July 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Huang:2013:TOQ

[HPO^{+}13] Shihong Huang, Dragutin Petkovic, Kazunori Okada, Marc Sosnick, Shenhaochen Zhu, and Rainer Todtenhoefer. Toward objective and quantitative assessment and prediction of teamwork effectiveness in software engineering courses. ACM SIGSOFT Software Engineering Notes, 38(1):7–9, January 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hall:2010:IIW

[HRZN10] Jon G. Hall, Lucia Rapanotti, Liping Zhao, and James Naish. 2010 ICSE International Workshop on Advances and Applications of Problem Orientation (WAAPO-2010). ACM SIGSOFT Software Engineering Notes, 35(5):40– 41, September 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hughes:2016:BRA

[HSS^{+}16] Jeffrey Hughes, Cassandra Sparks, Alley Stoughton, Rinku Parikh, Albert Reuther, and Suresh Jagannathan. Building Resource Adaptive Software Systems (BRASS): Objectives and system evaluation. ACM SIGSOFT Software Engineering Notes, 41 (1):1–2, January 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hunt:2013:RDA

[HW13] Johanna Hunt and Xiaofeng Wang. Research Dojo: applying agile principles to agile research — workshop summary from XP2013. ACM SIG-SOFT Software Engineering Notes, 38(5):43–46, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hofer:2012:AWI

[HWA12] Birgit Hofer, Franz Wotawa, and Rui Abreu. AI for the win: improving spectrumbased fault localization. ACM SIGSOFT Software Engineering Notes, 37(6):1–8, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Izurieta:2018:TDR

[IOSS18] Clemente Izurieta, Ipek Ozkaya, Carolyn Seaman, and Will Snipes. Technical debt: a research roadmap report on the Eighth Workshop on Managing Technical Debt (MTD 2016). ACM SIGSOFT Software Engineering Notes, 43 (2):28–31, April 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jahns:2012:PDI

[Jah12] Veit Jahns. Principles of data integration by anhai doan, Alon halevy, zachary Ives. ACM SIGSOFT Software Engineering Notes, 37 (5):43, September 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jahns:2013:BRD

[Jah13] Veit Jahns. Book review: Data insights: new ways to visualize and make sense of data by Hunter Whitney. ACM SIG-SOFT Software Engineering Notes, 38(6):45–46, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jain:2011:ARM

[Jai11] Ajay Jain. Approach for reducing menu access time by enabling bidirectional cursor movement within nested menu(s). ACM SIGSOFT Software Engineering Notes, 36(5):1–8, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jain:2012:OFA

[Jai12] Ajay Jain. Optimizing featureaccess time through dynamic updates to application menu layout. ACM SIGSOFT Software Engineering Notes, 37 (5):1–14, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jain:2013:TTO

[Jai13] Ajay Jain. Touch target optimization technique using virtual finger-tip library. ACM SIGSOFT Software Engineering Notes, 38(3):1–9, May 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Janus:2012:TCA

[Jan12] André Janus. Towards a common agile software development model (ASDM). ACM SIGSOFT Software Engineering Notes, 37(4):1–8, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jain:2013:MSD

[JC13] Ajay Jain and Kusha Chopra. Malware signing detection system. ACM SIGSOFT Software Engineering Notes, 38 (5):1–8, September 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jeet:2013:SRE

[JD13] Kawal Jeet and Renu Dhir. Software re-engineering using imperialist competitive algorithm. ACM SIGSOFT Software Engineering Notes, 38 (6):1–5, November 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jeet:2012:CSB

[JDV12] Kawal Jeet, Renu Dhir, and Harsh Verma. A comparative study of Bayesian and fuzzy approach to assess and predict maintainability of the software using activity-based quality model. ACM SIG-SOFT Software Engineering Notes, 37(3):1–9, May 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jain:2012:CSV

[JG12] Hemant Jain and Anjana Gosain. A comprehensive study of view maintenance approaches in data warehousing evolution. ACM SIG-SOFT Software Engineering Notes, 37(5):1–8, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jain:2013:NAS

[JG13] Paritosh Jain and Nitish Garg.

A novel approach for slicing of object oriented programs. ACM SIGSOFT Software Engineering Notes, 38 (4):1–4, July 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jie:2016:ICS

[Jie16] Jason Lee Hua Jie. Industrial case study of transition from V-Model into Agile SCRUM in embedded software testing industries. ACM SIGSOFT Software Engineering Notes, 41(2):1–3, March 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jin:2018:OMB

[Jin18] Zhi Jin. Open models: Beyond the open source software development. ACM SIGSOFT Software Engineering Notes, 43(4):9–12, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Johari:2011:ESE

[JK11] Kalpana Johari and Arvinder Kaur. Effect of software evolution on software metrics: an open source case study. ACM SIGSOFT Software Engineering Notes, 36(5):1–8, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Johari:2012:VOO

[JK12] Kalpana Johari and Arvinder Kaur. Validation of object oriented metrics using open source software system: an empirical study. ACM SIGSOFT Software Engineering Notes, 37(1):1–4, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jalila:2013:EEO

[JM13] A. Jalila and D. Jeya Mala. Empirical evidence on OCL formal specification-based metrics as a predictor of faultproneness. ACM SIGSOFT Software Engineering Notes, 38(5):1–10, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jones:2013:FPU

[Jon13] Capers Jones. Function points as a universal software metric. ACM SIGSOFT Software Engineering Notes, 38(4):1–27, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Johnson:2013:RSS

[JRG^{+}13] Pontus Johnson, Paul Ralph, Michael Goedicke, Pan-Wei Ng, Klaas-Jan Stol, Kari Smolander, Iaakov Exman, and Dewayne E. Perry. Report on the Second SEMAT Workshop on General Theory of Software Engineering (GTSE 2013). ACM SIG-SOFT Software Engineering Notes, 38(5):47–50, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jeet:2012:BNB

[JRX12] Kawal Jeet, Yadvirender Rana, and Ruichi Xin. A Bayesian network based approach for software reusability prediction. ACM SIGSOFT Software Engineering Notes, 37(4):1–5, July 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jain:2012:AYL

[JS12] Ajay Jain and Chandan Singh.

“Ad you like it”: advertisement sourcing and selection technique across multiple heterogeneous applications. ACM SIGSOFT Software Engineering Notes, 37(4):1–6, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jain:2018:MDN

[JS18a] Ajay Jain and Sachin Soni.

Multi-directional navigation method for optimized consumption of user generated content through semantic mapping of features derived from the user generated content. ACM SIGSOFT Software Engineering Notes, 43 (4):52, October 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jain:2018:PMF

[JS18b] Ritu Jain and Ugrasen Suman. A project management framework for global software development. ACM SIGSOFT Software Engineering Notes, 43 (1):1–10, January 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jones:2018:MSC

[JWB^{+}18] Joseph Jones, James Wasson, Sean Brown, Seth Poulsen, Peter Aldous, and Eric Mercer. Memory safety in C by abstract interpretation. ACM SIGSOFT Software Engineering Notes, 43(4):56, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jiau:2012:FIC

[JY12] Hewijin Christine Jiau and Feng-Pu Yang. Facing up to the inequality of crowdsourced API documentation. ACM SIGSOFT Software Engineering Notes, 37(1):1–9, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jiang:2013:CBP

[JZGH13] JianMin Jiang, Shi Zhang, Ping Gong, and Zhong Hong. Configuring business process models. ACM SIGSOFT Software Engineering Notes, 38 (4):1–10, July 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jiang:2012:TDG

[JZY12] Shujuan Jiang, Yanmei Zhang, and Dandan Yi. Test data generation approach for basis path coverage. ACM SIG-SOFT Software Engineering Notes, 37(3):1–7, May 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kamei:2019:UGL

[Kam19] Fernando Kenji Kamei. The use of grey literature review as evidence for practitioners. ACM SIGSOFT Software Engineering Notes, 44(3):23, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3356773.3356797.

Katic:2013:WAC

[Kat13] Marija Katic. Why attending conferences is important for your research: ESEC/ FSE 2013 experience summary. ACM SIGSOFT Software Engineering Notes, 38 (6):5–6, November 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kayes:2011:ATI

[Kay11] Imrul Kayes. Agile testing: Introducing PRAT as a metric of testing quality in Scrum. ACM SIGSOFT Software Engineering Notes, 36(2):1–5, March 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2014:DLB

[KAZS14] Narander Kumar, Shalini Agarwal, Taskeen Zaidi, and Vipin Saxena. A distributed load-balancing scheme based on a complex network model of cloud servers. ACM SIG-SOFT Software Engineering Notes, 39(6):1–6, November 2014. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumari:2011:AOO

[KB11a] Usha Kumari and Sucheta Bhasin. Application of objectoriented metrics to C++ and Java: a comparative study. ACM SIGSOFT Software Engineering Notes, 36 (2):1–10, March 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumari:2011:CCM

[KB11b] Usha Kumari and Sucheta Bhasin. A composite complexity measure for componentbased systems. ACM SIG-SOFT Software Engineering Notes, 36(6):1–5, November 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krishnamurthy:2012:PBA

[KB12] Vallidevi Krishnamurthy and Chitra Babu. Pattern based adaptation for service oriented applications. ACM SIGSOFT Software Engineering Notes, 37(1):1–6, January 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2017:BSA

[KBRS17a] Lov Kumar, Ranjan Kumar Behera, Santanu Rath, and Ashish Sureka. A bibliometric study of ACM SIGSOFT software engineering notes from 2007 to 2016. ACM SIGSOFT Software Engineering Notes, 42(3):1–7, July 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2017:TLC

[KBRS17b] Lov Kumar, Ranjan Kumar Behera, Santanu Rath, and Ashish Sureka. Transfer learning for cross-project changeproneness prediction in objectoriented software systems: a feasibility analysis. ACM SIG-SOFT Software Engineering Notes, 42(3):1–11, July 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Khatri:2011:MBC

[KCS11] Sujata Khatri, R. S. Chhillar, and V. B. Singh. Measuring bug complexity in object oriented software system. ACM SIGSOFT Software Engineering Notes, 36(6):1–8, November 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2011:RBC

[KD11] Pankaj Kumar and Kamlesh Dutta. Relationship between crosscutting concerns and defects. ACM SIGSOFT Software Engineering Notes, 36 (2):1–7, March 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2011:QOR

[KGS11] Manoj Kumar, Anjana Gosain, and Yogesh Singh. Quality-oriented requirements engineering for a data warehouse. ACM SIGSOFT Software Engineering Notes, 36 (5):1–4, September 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kienle:2012:PDP

[Kie12] Holger M. Kienle. Personal data privacy and protection in a surveillance era: technologies and practices by christina akrivopoulou and athanasios psygkas. ACM SIGSOFT Software Engineering Notes, 37(5):45–46, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kienle:2013:BRE

[Kie13a] Holger M. Kienle. Book review: Effective JavaScript: 68 specific ways to harness the power of JavaScript by David Herman. ACM SIG-SOFT Software Engineering Notes, 38(6):40–41, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kienle:2013:BRG

[Kie13b] Holger M. Kienle. Book review: Grounded innovation: strategies for creating digital products by Lars Erik Holmquist. ACM SIGSOFT Software Engineering Notes, 38(1):55–56, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kimm:2013:BRO

[Kim13]  Larry Kimm.  Book re-

view: It’s our research: getting stakeholder buy-in for user experience research projects by Tomer Sharon. ACM SIG-SOFT Software Engineering Notes, 38(4):38–39, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krizevnik:2010:ISP

[KJ10] Marcel Krizevnik and Matjaz B. Juric. Improved SOA persistence architectural model. ACM SIGSOFT Software Engineering Notes, 35(3):1, May 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Khajaria:2011:MSR

[KK11] Krishna Khajaria and Manoj Kumar. Modeling of security requirements for decision information systems. ACM SIGSOFT Software Engineering Notes, 36(5):1–4, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Khan:2012:IQM

[KK12a] Suhel Ahmad Khan and Raees Ahmad Khan. Integrity quantification model for object oriented design. ACM SIGSOFT Software Engineering Notes, 37(2):1–3, March 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2012:PSA

[KK12b] Surender Kumar and Rajeev Kumar. Precise static analysis for generic programs in object oriented languages. ACM SIG-SOFT Software Engineering Notes, 37(3):1–6, May 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2013:RBR

[KK13] Kuldeep Kumar and Sandeep Kumar. A rule-based recommendation system for selection of software development life cycle models. ACM SIGSOFT Software Engineering Notes, 38(4):1–6, July 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaur:2014:MSI

[KK14] Navdeep Kaur and Parminder Kaur. Mitigation of SQL injection attacks using threat modeling. ACM SIGSOFT Software Engineering Notes, 39 (6):1–6, November 2014. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumari:2019:EGP

[KK19] Neha Kumari and Rajeev Kumar. Evolution of generic programming in OOPLs. ACM SIGSOFT Software Engineering Notes, 44(1):35, January 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2011:MSR

[KKK11] Ravinder Kumar, Kiran Khatter, and Arvind Kalia. Measuring software reliability: a fuzzy model. ACM SIG-SOFT Software Engineering Notes, 36(6):1–6, November 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Knutson:2010:RIW

[KKPJ10] Charles D. Knutson, Jonathan L. Krein, Lutz Prechelt, and Natalia Juristo. Report from the 1st International Workshop on Replication in Empirical Software Engineering Research (RESER 2010). ACM SIG-SOFT Software Engineering Notes, 35(5):42–44, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krein:2012:RIW

[KKPJ12] Jonathan L. Krein, Charles D. Knutson, Lutz Prechelt, and Natalia Juristo. Report from the 2nd International Workshop on Replication in Empirical Software Engineering Research (RESER 2011). ACM SIGSOFT Software Engineering Notes, 37(1):27–30, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kraemer:2018:RAC

[KL18] Eileen Kraemer and Aubrey Lawson. Reasoning about concurrency: Scenarios for activities. ACM SIGSOFT Software Engineering Notes, 43(3):18, July 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Karre:2019:RSE

[KMR^{+}19] Sai Anirudh Karre, Lalit Mohan, Y. Raghu Raghu Reddy, K. V. Raghavan, R. D. Naik, Rahul Purandare, and Amey Karkare. A report on 1st Software Engineering Research in India Update Meeting (SERI 2019). ACM SIGSOFT Software Engineering Notes, 44 (3):41–42, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3356773.3356805.

Kuhrmann:2017:SIW

[KMTD17] Marco Kuhrmann, Jürgen Münch, Paolo Tell, and Philipp Diebold. Summary of the 1st International Workshop on Hybrid Development Approaches in Software Systems Development. ACM SIGSOFT Software Engineering Notes, 42(4):18–20, October 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kruchten:2013:TDT

[KNOF13] Philippe Kruchten, Robert L.

Nord, Ipek Ozkaya, and Davide Falessi. Technical debt: towards a crisper definition report on the 4th International Workshop on Managing Technical Debt. ACM SIG-SOFT Software Engineering Notes, 38(5):51–54, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kruchten:2012:TDS

[KNOV12] Philippe Kruchten, Robert L.

Nord, Ipek Ozkaya, and Joost Visser. Technical debt in software development: from metaphor to theory report on the Third International Workshop on Managing Technical Debt. ACM SIGSOFT Software Engineering Notes, 37(5):36–38, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kuhrmann:2018:SICa

[KOH^{+}18a] Marco Kuhrmann, Rory V. O’Connor, Dan Houston, Regina Hebig, and David Raffo. Summary of the International Conference on Software and System Processes (ICSSP 2018). ACM SIG-SOFT Software Engineering Notes, 43(4):48–51, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kuhrmann:2018:SICb

[KOH^{+}18b] Marco Kuhrmann, Rory V. O’Connor, Dan Houston, Regina Hebig, and David Raffo. Summary of the International Conference on Software and System Processes (ICSSP 2018). ACM SIG-SOFT Software Engineering Notes, 43(4):54, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kuhrmann:2016:SIC

[KOPR16] Marco Kuhrmann, Rory V. O’Connor, Dewayne E. Perry, and David Raffo. Summary of the International Conference on Software and System Processes (ICSSP 2016): [co-located with ICSE 2016]. ACM SIGSOFT Software Engineering Notes, 41(5):27– 30, September 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

kotti:2016:QSA

[kP16] Jayasri kotti and Seetharamaiah Panchumarthy. The quantitative safety assessment and evaluation for safety-critical computer systems. ACM SIGSOFT Software Engineering Notes, 41(1):1–8, January 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2010:SMM

[KPA10] Krishna Kumar, Rajesh Prasad, and Suneeta Agarwal. Software maintenance by multipatterns parameterized string matching with q-gram. ACM SIGSOFT Software Engineering Notes, 35(3):1–5, May 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Khyzha:2012:AP

[KPP12] Artem Khyzha, Pavel Parízek, and Corina S. Păsăreanu. Abstract pathfinder. ACM SIG-SOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Kraemer:2018:TDC

[Kra18] Eileen Kraemer. Teaching the design-by-contract concept in a software engineering course using RESOLVE. ACM SIG-SOFT Software Engineering Notes, 43(3):18, July 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kreitz:2019:SDS

[Kre19] Mark Kreitz. Security by design in software engineering. ACM SIGSOFT Software Engineering Notes, 44 (3):23, November 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3356773.3356798.

Krishnamurthi:2013:AES

[Kri13] Shriram Krishnamurthi. Artifact evaluation for software conferences. ACM SIGSOFT Software Engineering Notes, 38(3):7–10, May 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kalaimagal:2010:QIQ

[KS10] Sivamuni Kalaimagal and Rengaramanujam Srinivasan. Q’Facto 12: an improved quality model for COTS components. ACM SIGSOFT Software Engineering Notes, 35 (2):1–4, March 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaur:2011:DMI

[KS11a] Kulwant Kaur and Hardeep Singh. Determination of maintainability index for object oriented systems. ACM SIG-SOFT Software Engineering Notes, 36(2):1–6, March 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaur:2011:MVC

[KS11b] Parminder Kaur and Hardeep Singh. A model for versioning control mechanism in component-based systems. ACM SIGSOFT Software Engineering Notes, 36 (5):1–8, September 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krishna:2011:APV

[KS11c] Raj P. M. Krishna and K. G. Srinivasa. Analysis of projects and volunteer participation in large scale free and open source software ecosystem. ACM SIGSOFT Software Engineering Notes, 36 (2):1–5, March 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaur:2012:MVC

[KS12a] Parminder Kaur and Hardeep Singh. A modified version control tool for componentbased systems. ACM SIG-SOFT Software Engineering Notes, 37(6):1–10, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2012:AST

[KS12b] Pradeep Kumar and Yogesh Singh. Assessment of software testing time using soft computing techniques. ACM SIGSOFT Software Engineering Notes, 37(1):1–6, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaur:2013:MVC

[KS13a] Parminder Kaur and Hardeep Singh. A modified version control tool for componentbased systems. ACM SIG-SOFT Software Engineering Notes, 38(1):39–41, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kumar:2013:QAE

[KS13b] Rakesh Kumar and Hardeep Singh. A qualitative analysis of effects of security risks on architecture of an information system. ACM SIGSOFT Software Engineering Notes, 38 (6):1–3, November 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kukreja:2012:AMT

[KSR12] Deepika Kukreja, Umang Singh, and B. V. R. Reddy. Analytical models for trust based routing protocols in wireless ad hoc networks. ACM SIGSOFT Software Engineering Notes, 37(4):1–16, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lago:2010:OSA

[LAK10] Patricia Lago, Paris Avgeriou, and Philippe Kruchten. Organizing a software architecture body of knowledge: summary of the 5th SHARK workshop, at ICSE 2010. ACM SIG-SOFT Software Engineering Notes, 35(5):37–40, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Langsworth:2011:USA

[Lan11] Anthony Langsworth. Using static analysis tools to detect and correct non-compliant cryptography. ACM SIG-SOFT Software Engineering Notes, 36(6):1–7, November 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liu:2017:PMB

[LAX17] Pan Liu, Jun Ai, and Zhenning (Jimmy) Xu. Probability model-based test suite reduction. ACM SIGSOFT Software Engineering Notes, 42(3):1–6, July 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lee:2010:OIP

[Lee10] Tony Tony Lee. Optimizing IT process management. ACM SIGSOFT Software Engineering Notes, 35(4):1–10, July 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lee:2018:OTC

[Lee18] Amanda Lee. One-time contributors to FLOSS: Surveys and data analysis. ACM SIGSOFT Software Engineering Notes, 43(1):1–6, January 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Levine:2013:CSS

[Lev13] Gertrude N. Levine. Computer security with service degradations. ACM SIG-SOFT Software Engineering Notes, 38(6):1–7, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Louridas:2012:NRR

[LG12] Panos Louridas and Georgios Gousios. A note on rigour and replicability. ACM SIG-SOFT Software Engineering Notes, 37(5):1–4, September 2012. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Li:2010:TER

[LGMM10] Zude Li, Mechelle Gittens, Syed Shariyar Murtaza, and Nazim H. Madhavji. A towards an extended relational algebra for software architecture. ACM SIGSOFT Software Engineering Notes, 35 (3):1–4, May 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lewis:2013:RIIa

[LIL13] Grace A. Lewis, Anca Ionita, and Marin Litoiu. Report of the 2012 IEEE 6th International Workshop on the Maintenance and Evolution of Service-Oriented and Cloud-Based Systems (MESOCA 2012). ACM SIGSOFT Software Engineering Notes, 38 (2):29–32, March 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lago:2013:EIC

[LKM^{+}13] Patricia Lago, Rick Kazman, Niklaus Meyer, Maurizio Morisio, Hausi A. Müller, and Frances Paulisch. Exploring initial challenges for green software engineering: summary of the first GREENS workshop, at ICSE 2012. ACM SIGSOFT Software Engineering Notes, 38(1):31–33, January 2013. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Lago:2012:RIW

[LLM^{+}12] Patricia Lago, Grace A. Lewis, Andreas Metzger, Vladimir Tosic, Domenico Bianculli, Antinisca Di Marco, Andrea Polini, and Pierluigi Plebani. Report of the 4th International Workshop on Principles of Engineering Service-Oriented Systems (PESOS 2012). ACM SIGSOFT Software Engineering Notes, 37 (6):1–7, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lago:2013:RIW

[LLM^{+}13] Patricia Lago, Grace A. Lewis, Andreas Metzger, Vladimir Tosic, Domenico Bianculli, Antinisca Di Marco, Andrea Polini, and Pierluigi Plebani. Report of the 4th International Workshop on Principles of Engineering Service-Oriented Systems (PESOS 2012). ACM SIGSOFT Software Engineering Notes, 38 (1):35–38, January 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Li:2012:MQG

[LLS12] Dan Li, Xiaoshan Li, and Volker Stolz. Model querying with graphical notation of QVT relations. ACM SIG-SOFT Software Engineering Notes, 37(4):1–8, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lilja:2011:PAI

[LMS11] David J. Lilja, Raffaela Mirandola, and Kai Sachs. Paper abstracts of the 2nd International Conferernce on Performance Engineering (ICPE 2011). ACM SIGSOFT Software Engineering Notes, 36 (5):36–53, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lewis:2013:RIIb

[LNG^{+}13] Grace A. Lewis, Nachiappan Nagappan, Jeff Gray, David Rosenblum, Henry Muccini, and Emad Shihab. Report of the 2013 ICSE 1st International Workshop on Engineering Mobile-Enabled Systems (MOBS 2013): 12. ACM SIG-SOFT Software Engineering Notes, 38(5):55–58, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lochmann:2012:BIA

[Loc12] Klaus Lochmann. A benchmarkinginspired approach to determine threshold values for metrics. ACM SIGSOFT Software Engineering Notes, 37 (6):1–8, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Long:2010:TDSa

[Lon10a] Brad Long. Towards the design of a set-based Java collections framework. ACM SIGSOFT Software Engineering Notes, 35(5):1–7, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Long:2010:TDSb

[Lon10b] Brad Long. Towards the design of a set-based Java collections framework. ACM SIGSOFT Software Engineering Notes, 35(6):1–7, November 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Le:2019:SAG

[LPP^{+}19] Xuan-Bach D. Le, Corina Pasareanu, Rohan Padhye, David Lo, Willem Visser, and Koushik Sen. Saffron: Adaptive grammar-based fuzzing for worst-case analysis. ACM SIGSOFT Software Engineering Notes, 44(4):14, December 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3364452.3364455.

Lutteri:2011:RIS

[LRS11] Emiliano Lutteri, Barbara Russo, and Giancarlo Succi. Report of the 4th international symposium on empirical software engineering and measurement ESEM 2010. ACM SIGSOFT Software Engineering Notes, 36(2):28–34, March 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lewis:2010:RIW

[LSM^{+}10] Grace Lewis, Dennis Smith, Andreas Metzger, Andrea Zisman, and Marco Pistore. Report of the 2nd International Workshop on Principles of Engineering Service-Oriented Systems (PESOS 2010). ACM SIGSOFT Software Engineering Notes, 35 (5):30–33, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Langdon:2019:SFG

[LWT^{+}19] William B. Langdon, Westley Weimer, Christopher Timperley, Oliver Krauss, Zhen Yu Ding, Yiwei Lyu, Nicolas Chausseau, Eric Schulte, Shin Hwei Tan, Kevin Leach, Yu Huang, and Gabin An. The state and future of genetic improvement. ACM SIG-SOFT Software Engineering Notes, 44(3):25–29, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3356773.3356801.

Leite:2018:RIW

[LZK^{+}18] Julio C. S. P. Leite, Liping Zhao, Sylwia Kopczńska, Sam Supakkul, and Lawrence Chung. Report from the 6th International Workshop on Requirements Patterns (RePa’16). ACM SIGSOFT Software Engineering Notes, 43(2):32–33, April 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

M:2013:BRL

[M.13] Srinivasan M. Book review: Learning Rails 3 by Simon St. Laurent, Edd Dumbill and Eric J. Gruber. ACM SIG-SOFT Software Engineering Notes, 38(2):34, March 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Macedo:2010:MDD

[Mac10]  Hendrik Macedo.  Model

driven development approach to natural language generation systems. ACM SIGSOFT Software Engineering Notes, 35(4):1–7, July 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Matsubara:2019:DSE

[Mat19] Patrícia Gomes Fernandes Matsubara. Dealing with software estimates distortions from the perspective of negotiation theories. ACM SIG-SOFT Software Engineering Notes, 44(3):22, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3356773.3356794.

Majumdar:2012:ICF

[MB12] Dipankar Majumdar and Swapan Bhattacharya. Interoperability of constrained finite state automata. ACM SIG-SOFT Software Engineering Notes, 37(2):1–8, March 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mattmann:2010:UAT

[MBC10] Chris A. Mattmann, Amy J.

Braverman, and Daniel J. Crichton. Understanding architectural tradeoffs necessary to increase climate model intercomparison efficiency. ACM SIGSOFT Software Engineering Notes, 35(3):1–6, May 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mala:2013:CAT

[MBN13] D. Jeya Mala, S. Balamurugan, and K. Sabari Nathan. Criticality analyzer and tester: an effective approach for critical component identification & verification using ABC. ACM SIGSOFT Software Engineering Notes, 38(6):1–12, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mohana:2012:AIP

[MD12] Rajni Mohana and Deepak Dahiya. Approach and impact of a protocol for selection of service in web service platform. ACM SIGSOFT Software Engineering Notes, 37 (1):1–6, January 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mei:2017:RSE

[Mei17] Hong Mei. A review of software engineering research in China. ACM SIGSOFT Software Engineering Notes, 42 (4):6–9, October 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mendell:2012:BRP

[Men12] Matt Mendell. Book review: A practical guide to Fedora and Red Hat enterprise Linux by Mark G. Sobell. ACM SIGSOFT Software Engineering Notes, 37(1):36, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Meng:2013:PBL

[Men13] Zhang Meng. PPCA-based license plate detection algorithm. ACM SIGSOFT Software Engineering Notes, 38 (3):1–4, May 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mandrioli:2010:SFS

[MFF^{+}10] Dino Mandrioli, Stephen Fickas, Carlo A. Furia, Mehdi Jazayeri, Matteo Rossi, and Michal Young. SCORE: the first student contest on software engineering. ACM SIG-SOFT Software Engineering Notes, 35(4):24–30, July 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

M:2012:ESG

[MG12] Krishna Raj P. M. and Srinivasa K. G. Empirical studies of global volunteer collaboration in the development of free and open source software: analysis of six top ranked projects

in sourceforge.net. ACM

SIGSOFT Software Engineering Notes, 37(2):1–11, March 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Milhau:2012:RPA

[MGLF12] Jérémy Milhau, Frédéric Gervais, Régine Laleau, and Marc Frappier. Refinement patterns for ASTD. ACM SIGSOFT Software Engineering Notes, 37(4):1–8, July 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Miranda:2011:TBP

[Mir11]  Eduardo Miranda.  Time

boxing planning: buffered Moscow rules. ACM SIG-SOFT Software Engineering Notes, 36(6):1–5, November 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mitrache:2011:BRI

[Mit11] Cristina Mitrache. Book review: IT manager’s handbook, the business edition by Bill Holtsnider and Brian D. Jaffe. ACM SIGSOFT Software Engineering Notes, 36 (5):51, September 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Miyachi:2011:ASA

[Miy11] Christine Miyachi. Agile software architecture. ACM SIG-SOFT Software Engineering Notes, 36(2):1–3, March 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Malhotra:2011:SFP

[MJ11] Ruchika Malhotra and Ankita Jain. Software fault prediction for object oriented systems: a literature review. ACM SIGSOFT Software Engineering Notes, 36(5):1–6, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Monteiro:2017:BTM

[MJCdLF17] Felipe R. Monteiro, Francisco A. P. Januário, Lucas C. Cordeiro, and Eddie B. de Lima Filho. BMCLua: a translator for model checking Lua programs. ACM SIG-SOFT Software Engineering Notes, 42(3):1–10, July 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Majumdar:2011:SSC

[MKB11] Dipankar Majumdar, Ananya Kanjilal, and Swapan Bhattacharya. Separation of scattered concerns: a graph based approach for aspect mining. ACM SIGSOFT Software Engineering Notes, 36 (2):1–11, March 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Misra:2012:JSC

[MKK^{+}12a] Janardan Misra, Annervaz KM, Vikrant Kaulgud, Shubhashis Sengupta, and Gary Titus. Java source-code clustering: Unifying syntactic and semantic features. ACM SIG-SOFT Software Engineering Notes, 37(6):1–8, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mrunalini:2012:DPM

[MKK12b] M. Mrunalini, T. V. Suresh Kumar, and K. Rajani Kanth. Dynamic process model for identifying modified data using mobile agents in real time ETL processes. ACM SIG-SOFT Software Engineering Notes, 37(6):1–9, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Misra:2013:JSC

[MKK^{+}13a] Janardan Misra, Annervaz KM, Vikrant Kaulgud, Shubhashis Sengupta, and Gary Titus. Java source-code clustering: Unifying syntactic and semantic features. ACM SIG-SOFT Software Engineering Notes, 38(1):41–43, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mrunalini:2013:DPM

[MKK13b] M. Mrunalini, T. V. Suresh Kumar, and K. Rajani Kanth. Dynamic process model for identifying modified data using mobile agents in real time ETL processes. ACM SIG-SOFT Software Engineering Notes, 38(1):43–46, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mahajan:2012:AGA

[MKP12] Manish Mahajan, Sumit Kumar, and Rabins Porwal. Applying genetic algorithm to increase the efficiency of a data flow-based test data generation approach. ACM SIG-SOFT Software Engineering Notes, 37(5):1–5, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Malhotra:2010:AML

[MKS10] Ruchika Malhotra, Arvinder Kaur, and Yogesh Singh. Application of machine learning methods for software effort prediction. ACM SIGSOFT Software Engineering Notes, 35(3):1–6, May 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mala:2010:QIO

[MM10a] D. Jeya Mala and V. Mohan.

Quality improvement and optimization of test cases: a hybrid genetic algorithm based approach. ACM SIGSOFT Software Engineering Notes, 35(3):1–14, May 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mishra:2010:PMS

[MM10b] Amit Mishra and Sanjay Misra. People management in software industry: the key to success. ACM SIGSOFT Software Engineering Notes, 35 (6):1–4, November 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mishra:2011:CAS

[MM11a] Alok Mishra and Deepti Mishra. A curriculum for agile software development methodologies. ACM SIGSOFT Software Engineering Notes, 36 (3):1–2, May 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mishra:2011:ILG

[MM11b] Alok Mishra and Deepti Mishra. Industry linked graduate software engineering curriculum. ACM SIGSOFT Software Engineering Notes, 36(6):1–4, November 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mishra:2013:CLS

[MM13a] Alok Mishra and Deepti Mishra. A curriculum for large scale software development. ACM SIGSOFT Software Engineering Notes, 38 (1):1–2, January 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mishra:2013:SPM

[MM13b] Alok Mishra and Deepti Mishra. Software project management tools: a brief comparative view. ACM SIGSOFT Software Engineering Notes, 38(3):1–4, May 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mishra:2010:RIS

[MMM10] Alok Mishra, Jürgen Münch, and Deepti Mishra. A report on the “Information Systems in Distributed Environments” (ISDE) Workshop at the OTM 2009 conferences. ACM SIGSOFT Software Engineering Notes, 35(3):33–34, May 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mishra:2011:ISD

[MMM11] Alok Mishra, Jürgen Münch, and Deepti Mishra. Information systems in distributed environments: ISDE 2010. ACM SIGSOFT Software Engineering Notes, 36(3):28–30, May 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mishra:2013:ISD

[MMM13] Alok Mishra, Jürgen Münch, and Deepti Mishra. Information systems in distributed environments: ISDE 2013. ACM SIGSOFT Software Engineering Notes, 38(6):34– 35, November 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mishra:2016:ISD

[MMM16] Alok Mishra, Jürgen Münch, and Deepti Mishra. Information systems in distributed environments 2015. ACM SIG-SOFT Software Engineering Notes, 41(4):25–26, July 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mirzaei:2012:TAA

[MMP^{+}12] Nariman Mirzaei, Sam Malek, Corina S. Păsăreanu, Naeem Esfahani, and Riyadh Mahmood. Testing Android apps through symbolic execution. ACM SIGSOFT Software Engineering Notes, 37 (6):1–5, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mala:2013:CCT

[MNB13] D. Jeya Mala, K. Sabari Nathan, and S. Balamurugan. Critical components testing using hybrid genetic algorithm. ACM SIGSOFT Software Engineering Notes, 38(5):1–13, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Misra:2011:SAM

[MO11] Sanjay Misra and Martha Omorodion. Survey on agile metrics and their interrelationship with other traditional development metrics. ACM SIGSOFT Software Engineering Notes, 36 (6):1–3, November 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Moreland:2013:BRP

[Mor13] Doug Moreland. Book review: A practical approach to large-scale agile development: how HP transformed Laser-Jet FutureSmart firmware by Gary Gruver, Mike Young and Pat Fulghm. ACM SIG-SOFT Software Engineering Notes, 38(6):41–42, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Motta:2019:EBF

[Mot19]  Rebeca C. Motta.  An

evidence-based framework for supporting the engineering of IoT software systems. ACM SIGSOFT Software Engineering Notes, 44(3):22– 23, November 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3356773.3356795.

Milewicz:2017:SPM

[MP17] Reed M. Milewicz and Simon Poulding. Scalable parallel model checking via Monte-Carlo tree search. ACM SIGSOFT Software Engineering Notes, 42(4):1–5, October 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mercer:2012:CVI

[MPR12] Eric Mercer, Suzette Person, and Neha Rungta. Computing and visualizing the impact of change with Java PathFinder extensions. ACM SIGSOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mohalik:2017:WDA

[MRJD17] Swarup Kumar Mohalik, Badrinath Ramamurthy, Mahesh Babu Jayaraman, and Meenakshi D’Souza. Workshop on Developmental aspects of Intelligent Adaptive Systems (DIAS): Co-located with 10th Innovations in Software Engineering Conference (ISEC), Jaipur, India. ACM SIGSOFT Software Engineering Notes, 42(4):25–27, October 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mary:2013:PSA

[MRN13] S. Roselin Mary, Paul Rodrigues, and E. R.. Naganathan. Patterns of software architecture in Vastu: a new revelation. ACM SIG-SOFT Software Engineering Notes, 38(6):1–6, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marijan:2019:GPA

[MS19] Dusica Marijan and Sagar Sen. Good practices in aligning software engineering research and industry practice. ACM SIGSOFT Software Engineering Notes, 44 (3):65–67, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3356773.3356812.

Majumdar:2010:MRM

[MSK^{+}10] Dipankar Majumdar, Sabnam Sengupta, Ananya Kanjilal, Swagata Kundu, and Swapan Bhattacharya. A mathematical reusability model for quantifying the reduction in development effort. ACM SIG-SOFT Software Engineering Notes, 35(4):1–8, July 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mukherjee:2018:PSI

[MSM18] Debashis Mukherjee, Dibyanshu Shekhar, and Rajib Mall. Proposal for a structural integration test coverage metric for object-oriented programs. ACM SIGSOFT Software Engineering Notes, 43 (1):1–4, January 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Marijan:2019:IRA

[MSS19] Dusica Marijan, Weiyi Shang, and Rakesh Shukla. Implications of resurgence in artificial intelligence for research collaborations in software engineering. ACM SIGSOFT Software Engineering Notes, 44(3):68–70, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3356773.3356813.

Malhotra:2013:DFT

[MT13] Ruchika Malhotra and Divya Tiwari. Development of a framework for test case prioritization using genetic algorithm. ACM SIGSOFT Software Engineering Notes, 38 (3):1–6, May 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Muna:2019:APL

[Mun19] Altherwi Muna. Assessing programming language impact on software development productivity based on mining OSS repositories. ACM SIG-SOFT Software Engineering Notes, 44(1):36–37, January 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mohanta:2010:AEP

[MVGM10] Sirsendu Mohanta, Gopika Vinod, A. K. Ghosh, and Rajib Mall. An approach for early prediction of software reliability. ACM SIGSOFT Software Engineering Notes, 35 (6):1–9, November 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Meneely:2012:ICM

[MW12] Andrew Meneely and Oluyinka Williams. Interactive churn metrics: socio-technical variants of code churn. ACM SIGSOFT Software Engineering Notes, 37(6):1–6, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mascardi:2019:EMAa

[MWR19a] Viviana Mascardi, Danny Weyns, and Alessandro Ricci. Engineering multi-agent systems: State of affairs and the road ahead. ACM SIG-SOFT Software Engineering Notes, 44(1):18–28, January 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mascardi:2019:EMAb

[MWR19b] Viviana Mascardi, Danny Weyns, and Alessandro Ricci. Engineering multi-agent systems: State of affairs and the road ahead. ACM SIGSOFT Software Engineering Notes, 44(1):35, January 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nair:2010:DPM

[NAS10] T. R. Gopalakrishnan Nair, Sri Aravindh, and R. Selvarani. Design property metrics to maintainability estimation: a virtual method using functional relationship mapping. ACM SIGSOFT Software Engineering Notes, 35 (6):1–6, November 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neogi:2010:EEV

[NB10] Madhumita S. Neogi and Vandana Bhattacherjee. Evaluating the effectiveness of VOSDM: a vision oriented approach. ACM SIGSOFT Software Engineering Notes, 35 (2):1–8, March 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neto:2019:SSR

[Net19] Amadeu Anderlin Neto. A strategy to support replications of controlled experiments in software engineering. ACM SIGSOFT Software Engineering Notes, 44(3):23, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3356773.3356796.

Neumann:2010:RPa

[Neu10a] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 35(2):15–24, March 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2010:RPb

[Neu10b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 35(3):24–32, May 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2010:RPc

[Neu10c] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 35(4):17–23, July 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2010:RPd

[Neu10d] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 35(5):18–24, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2010:RPe

[Neu10e] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 35(6):19–26, November 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2011:RPa

[Neu11a] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 36(2):19–27, March 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2011:RPb

[Neu11b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 36(3):19–25, May 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2011:RPc

[Neu11c] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 36(5):20–23, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2011:RPd

[Neu11d] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 36(6):19–23, November 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2012:RPa

[Neu12a] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 37(1):21–26, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2012:RPb

[Neu12b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 37(2):21–29, March 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2012:RPc

[Neu12c] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 37(3):18–25, May 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2012:RPd

[Neu12d] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 37(4):20–29, July 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2012:RPe

[Neu12e] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 37(5):27–33, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2012:RPf

[Neu12f] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 37(6):19–23, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2013:RPa

[Neu13a] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 38(1):20–26, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2013:RPb

[Neu13b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 38(2):20–25, March 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2013:RPc

[Neu13c] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 38(3):21–28, May 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2013:RPd

[Neu13d] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 38(4):19–24, July 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2013:RPe

[Neu13e] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 38(5):28–33, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2013:RPf

[Neu13f] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 38(6):21–28, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2014:RP

[Neu14] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 39(6):17–22, November 2014. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2016:RPa

[Neu16a] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 41(1):18–24, January 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2016:RPb

[Neu16b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 41(2):16–22, March 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2016:RPc

[Neu16c] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 41(4):18–24, July 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2016:RPd

[Neu16d] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 41(5):19–26, September 2016. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2017:RPa

[Neu17a] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 42(2):7–14, April 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2017:RPb

[Neu17b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 42(3):10–17, July 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2017:RPc

[Neu17c] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 42(4):10–17, October 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2018:RPa

[Neu18a] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 43(1):1–6, January 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2018:RPb

[Neu18b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 43(2):8–11, April 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2018:RPc

[Neu18c] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 43(2):9–16, April 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2018:RPd

[Neu18d] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 43(3):11–16, July 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2018:RPe

[Neu18e] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 43(4):13–20, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2019:RPa

[Neu19a] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 44(1):11–17, January 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2019:RPb

[Neu19b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 44(2):6–12, April 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Neumann:2019:RPc

[Neu19c] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 44(3):16–21, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3356773.3372312.

Neumann:2019:RPd

[Neu19d] Peter G. Neumann. Risks to the public. ACM SIG-SOFT Software Engineering Notes, 44(4):5–10, December 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3364452.3364453.

Nautiyal:2014:MRC

[NGD14a] Lata Nautiyal, Neena Gupta, and Sushil Chandra Dimri. Measurement of the reliability of a component-based development using a path-based approach. ACM SIGSOFT Software Engineering Notes, 39 (6):1–5, November 2014. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nautiyal:2014:NAC

[NGD14b] Lata Nautiyal, Neena Gupta, and Sushil Chandra Dimri. A novel approach to componentbased software testing. ACM SIGSOFT Software Engineering Notes, 39(6):1–4, November 2014. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ngo:2011:BRD

[Ngo11] Terry Ngo. Book review:

Data mining: practical machine learning tools and technique, third edition by Ian H. Witten, Eibe Frank, Mark A. Hell. ACM SIGSOFT Software Engineering Notes, 36(5):51–52, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ngo:2012:BRE

[Ngo12] Terry Ngo. Book review: Essential app engine: building high-performance Java apps with Google app engine by Adriaan de Jonge. ACM SIG-SOFT Software Engineering Notes, 37(2):37, March 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nierstrasz:2012:ASA

[Nie12] Oscar Nierstrasz. Agile software assessment with moose. ACM SIGSOFT Software Engineering Notes, 37(3):1–5, May 2012. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

N:2013:MER

[NK13] Parimala N. and Ranjeet Kumar. Mapping extended rationale diagrams to OLAP queries. ACM SIGSOFT Software Engineering Notes, 38 (3):1–6, May 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nerurkar:2010:ARA

[NKS10] N. W. Nerurkar, Avadhesh Kumar, and Pallavi Shrivastava. Assessment of reusability in aspect-oriented systems using fuzzy logic. ACM SIG-SOFT Software Engineering Notes, 35(5):1–5, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Namakonov:2019:SDR

[NMPS19] Egor Namakonov, Eric Mercer, Pavel Parizek, and Kyle Storey. Symbolic data race detection for Habanero programs. ACM SIGSOFT Software Engineering Notes, 44(4):18, December 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3364452.33644562.

Nuthakki:2011:UUG

[NMVS11] Murali K. Nuthakki, Mutlu Mete, Cihan Varol, and Sang C. Suh. UXSOM: UML generated XML to software metrics. ACM SIGSOFT Software Engineering Notes, 36(3):1–6, May 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Noller:2019:CSS

[NNT^{+}19] Yannic Noller, Hoang Lam Nguyen, Minxing Tang, Timo Kehrer, and Lars Grunske. Complete shadow symbolic execution with Java PathFinder. ACM SIGSOFT Software Engineering Notes, 44(4):15– 16, December 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3364452.33644558.

Noller:2017:SSE

[NNTK17] Yannic Noller, Hoang Lam Nguyen, Minxing Tang, and Timo Kehrer. Shadow symbolic execution with Java PathFinder. ACM SIGSOFT Software Engineering Notes, 42(4):1–5, October 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Northrop:2018:DSSb

[NOFK18a] Lina Northrop, Ipek Ozkaya, George Fairbanks, and Michael Keeling. Designing the software systems of the future. ACM SIGSOFT Software Engineering Notes, 43 (4):53, October 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Northrop:2018:DSSa

[NOFK18b] Linda Northrop, Ipek Ozkaya, George Fairbanks, and Michael Keeling. Designing the software systems of the future. ACM SIGSOFT Software Engineering Notes, 43(4):28– 30, October 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Noorwali:2018:SCD

[Noo18] Ibtehal Noorwali. Stakeholder concern-driven requirements analytics. ACM SIGSOFT Software Engineering Notes, 43(1):1–6, January 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:2010:ATF

[Not10] David Notkin. ACM TOSEM: FAQs and figures. ACM SIG-SOFT Software Engineering Notes, 35(3):5–6, May 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Notkin:2012:TN

[Not12] David Notkin. TOSEM news. ACM SIGSOFT Software Engineering Notes, 37(4):7–8, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nautiyal:2016:NAE

[NP16] Lata Nautiyal and Preeti. A novel approach of equivalence class partitioning for numerical input. ACM SIGSOFT Software Engineering Notes, 41(1):1–5, January 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nair:2010:PMB

[NS10a] T. R. Gopalakrishnan Nair and V. Suma. A paradigm for metric based inspection process for enhancing defect management. ACM SIGSOFT Software Engineering Notes, 35(3):1, May 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nami:2010:SNN

[NS10b] Mohammad Reza Nami and Jila Saneipour. Self-* enursing: a new idea in nursing. ACM SIGSOFT Software Engineering Notes, 35 (2):1–4, March 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nagpal:2013:ASE

[NUK13] Geeta Nagpal, Moin Uddin, and Arvinder Kaur. Analyzing software effort estimation using k means clustered regression approach. ACM SIGSOFT Software Engineering Notes, 38(1):1–9, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ozkaya:2011:MTD

[OKNB11] Ipek Ozkaya, Philippe Kruchten, Robert L. Nord, and Nanette Brown. Managing technical debt in software development: report on the 2nd International Workshop on Managing Technical Debt, held at ICSE 2011. ACM SIG-SOFT Software Engineering Notes, 36(5):33–35, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

OSullivan:2011:DMM

[O’S11] Michael O’Sullivan. Designing with the mind in mind: simple guide to understanding user interface design rules by jeff Johnson. ACM SIG-SOFT Software Engineering Notes, 36(5):52, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Osterweil:2016:P

[Ost16a] Leon J. Osterweil. Be prepared. ACM SIGSOFT Software Engineering Notes, 41 (5):4–5, September 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Osterweil:2016:PEM

[Ost16b] Leon J. Osterweil. Preview: Ethical and moral issues for software engineers. ACM SIG-SOFT Software Engineering Notes, 41(4):5, July 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Osterweil:2017:R

[Ost17] Leon J. Osterweil. Be responsible. ACM SIGSOFT Software Engineering Notes, 42(3):5–8, July 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Osterweil:2018:G

[Ost18a] Leon J. Osterweil. Be gracious. ACM SIGSOFT Software Engineering Notes, 43 (2):4–6, April 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Osterweil:2018:YSD

[Ost18b] Leon J. Osterweil. Your software dwells in the house of tomorrow, too. ACM SIGSOFT Software Engineering Notes, 43(2):7–8, April 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ozturk:2016:IGTa

[ÖZ16a] Muhammed Maruf Öztürk and Ahmet Zengin. Improved GUI testing using task parallel library. ACM SIGSOFT Software Engineering Notes, 41(1):1–8, January 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ozturk:2016:IGTb

[ÖZ16b] Muhammed Maruf Öztürk and Ahmet Zengin. Improved GUI testing using task parallel library. ACM SIGSOFT Software Engineering Notes, 41(2):1–8, March 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Prasad:2010:PSM

[PA10] Rajesh Prasad and Suneeta Agarwal. Parameterized string matching: an application to software maintenance. ACM SIGSOFT Software Engineering Notes, 35(3):1–5, May 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Paige:2013:BRS

[Pai13a] Michael Paige. Book review: Sams teach yourself jQuery mobile in 24 hours by Phil Dutson. ACM SIGSOFT Software Engineering Notes, 38 (1):56–57, January 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Paige:2013:BRT

[Pai13b] Michael Paige. Book review: The tangled web: a guide to securing modern web applications by Michal Zalewski. ACM SIGSOFT Software Engineering Notes, 38(4):39–40, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pandey:2010:ADL

[Pan10] R. K. Pandey. Architectural description languages (ADLs) vs UML: a review. ACM SIG-SOFT Software Engineering Notes, 35(3):1–5, May 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pastor:2019:DSP

[Pas19] Oscar Pastor. Design science for PhD research in the software engineering domain. ACM SIGSOFT Software Engineering Notes, 44 (3):22, November 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3356773.3372314.

Pandey:2013:SEA

[PASS13] Adesh Kr. Pandey, C. P.

Agrawal, Arun Sharma, and P. Sasikala. Study of empirical approaches to analyze the software metrics. ACM SIGSOFT Software Engineering Notes, 38(4):1–5, July 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Payton:2013:BRD

[Pay13]  Ryan Payton.  Book re-

view: DevOps Troubleshooting by Kyle Rankin. ACM SIGSOFT Software Engineering Notes, 38(6):42, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parashar:2014:MCR

[PC14] Anshu Parashar and Jitender Kumar Chhabra. Measuring change-readiness of classes by mining changehistory. ACM SIGSOFT Software Engineering Notes, 39 (6):1–5, November 2014. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Priyanka:2012:EEC

[PCR12] Priyanka, Inderveer Chana, and Ajay Rana. Empirical evaluation of cloud-based testing techniques: a systematic review. ACM SIGSOFT Software Engineering Notes, 37 (3):1–9, May 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Panizo:2012:EJP

[PdMG12] Laura Panizo and María del Mar Gallardo. An extension of Java PathFinder for hybrid systems. ACM SIG-SOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Prikladnicki:2013:CHA

[PDS^{+}13] Rafael Prikladnicki, Yvonne Dittrich, Helen Sharp, Cleidson De Souza, Marcelo Cataldo, and Rashina Hoda. Cooperative and human aspects of software engineering: CHASE 2013. ACM SIG-SOFT Software Engineering Notes, 38(5):34–37, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pande:2013:OCS

[PGP13] Jeetendra Pande, Christopher J. Garcia, and Durgesh Pant. Optimal component selection for component based software development using pliability metric. ACM SIG-SOFT Software Engineering Notes, 38(1):1–6, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Phan:2018:TIG

[Pha18] Quoc-Sang Phan. Test input generation using separation logic. ACM SIGSOFT Software Engineering Notes, 43(4):55, October 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Panigrahi:2010:MBR

[PM10] Chhabi Rani Panigrahi and Rajib Mall. Model-based regression test case prioritization. ACM SIGSOFT Software Engineering Notes, 35 (6):1–7, November 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Patwa:2012:RME

[PM12] Sanjeev Patwa and Anil Kumar Malviya. Reusability metrics and effect of reusability on testing of object oriented systems. ACM SIGSOFT Software Engineering Notes, 37 (5):1–4, September 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Park:2016:SPE

[PMM16] June Sung Park, Paul E. McMahon, and Barry Myburgh. Scrum powered by essence. ACM SIGSOFT Software Engineering Notes, 41 (1):1–8, January 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Phan:2012:SQI

[PMTP12] Quoc-Sang Phan, Pasquale Malacaria, Oksana Tkachuk, and Corina S. Păsăreanu. Symbolic quantitative information flow. ACM SIGSOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Paquin:2018:AAS

[PSJ18] Maria Paquin, Elena Sherman, and Amit Jain. Assessing the adequacy of synthetic programs for learning SPF’s configurations. ACM SIGSOFT Software Engineering Notes, 43(4):55, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raina:2013:ATR

[RA13] Shikha Raina and Arun Prakash Agarwal. An automated tool for regression testing in web applications. ACM SIGSOFT Software Engineering Notes, 38(4):1–4, July 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rajaram:2010:ESV

[RB10] Kanchana Rajaram and Chitra Babu. Evolution of a simple vehicle registration system to an SOA based egovernance application: a case study. ACM SIGSOFT Software Engineering Notes, 35 (3):1–7, May 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rathee:2017:ROO

[RC17] Amit Rathee and Jitender Kumar Chhabra. Restructuring of object-oriented software through cohesion improvement using frequent usage patterns. ACM SIGSOFT Software Engineering Notes, 42(3):1–8, July 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raccoon:2013:U

[RD13]  Raccoon and Dog.  Un-

knownness. ACM SIGSOFT Software Engineering Notes, 38(5):8–17, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ralph:2014:HDG

[REN^{+}14] Paul Ralph, Iaakov Exman, Pan-Wei Ng, Pontus Johnson, Michael Goedicke, Alper Tolga Kocata, and Kate Liu Yan. How to develop a general theory of software engineering: Report on the GTSE 2014 workshop. ACM SIG-SOFT Software Engineering Notes, 39(6):23–25, November 2014. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raibulet:2018:RIWb

[RFD^{+}18a] Claudia Raibulet, Mariagrazia Fugini, Khalil Drira, Patrizio Pelliccione, Ilias Gerostathopoulos, Christian Prefoher, and Klaus Moessne. Report of the 1st International Workshop on Contextaware Autonomous and Smart Architectures (CASA@ECSA 2017). ACM SIGSOFT Software Engineering Notes, 43 (4):52–53, October 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raibulet:2018:RIWa

[RFD^{+}18b] Claudia Raibulet, Mariagrazia Fugini, Khalil Drira, Patrizio Pelliccione, Ilias Gerostathopoulos, Christian Prehofer, and Klaus Moessner. Report of the 1st International Workshop on Contextaware Autonomous and Smart Architectures (CASA@ECSA 2017). ACM SIGSOFT Software Engineering Notes, 43 (4):24–27, October 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Radhakishan:2010:CDI

[RFS10] V. Radhakishan, Yaser Farook, and S. Selvakumar. CRAYSE: design and implementation of efficient text search algorithm in a web crawler. ACM SIGSOFT Software Engineering Notes, 35(4):1–8, July 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Robles:2014:FRT

[RGBR14] Gregorio Robles, Jesús M.

González-Barahona, and Dirk Riehle. FLOSS Research Track at the 10th International Symposium on Open Collaboration (OpenSym 2014). ACM SIGSOFT Software Engineering Notes, 39(6):26– 27, November 2014. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rimlinger:2012:TGS

[Rim12] Frank Rimlinger. Test generation via symbolic simulation. ACM SIGSOFT Software Engineering Notes, 37 (6):1–5, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ripon:2010:PAS

[Rip10] Shamim H. Ripon. Process algebraic support for web service composition. ACM SIGSOFT Software Engineering Notes, 35(2):1–7, March 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ripon:2012:UTM

[Rip12] Shamim H. Ripon. A unified tabular method for modeling variants of software product line. ACM SIGSOFT Software Engineering Notes, 37(3):1–7, May 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ralph:2013:RFS

[RJJ13] Paul Ralph, Pontus Johnson, and Howell Jordan. Report on the First SEMAT Workshop on General Theory of Software Engineering (GTSE 2012). ACM SIGSOFT Software Engineering Notes, 38 (2):26–28, March 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rathore:2016:DTR

[RK16] Santosh Singh Rathore and Sandeep Kumar. A decision tree regression based approach for the number of software faults prediction. ACM SIGSOFT Software Engineering Notes, 41(1):1–6, January 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rabelo:2013:ACG

[RMFO13] Ricardo A. C. Rabelo, Hendrik T. Macedo, Eduardo O. Freire, and Rodrigo M. Oliveira. Automatic code generation of SIMUROSOT game strategies: an approach based on finite state machines. ACM SIGSOFT Software Engineering Notes, 38(2):1–8, March 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rogers:2010:BRG

[Rog10] David S. Rogers. Book review: Geometric Spanner Networks by Giri Narasimhan and Michael Smid, and published by Cambridge University Press, 2007, Hardback ISBN 978-0-521-81513-0, 500 pp., $105. ACM SIGSOFT Software Engineering Notes, 35(4):35, July 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Romanovsky:2012:DFM

[Rom12] Alexander Romanovsky. Deployment of formal methods in industry: the legacy of the FP7 ICT DEPLOY integrated project. ACM SIGSOFT Software Engineering Notes, 37 (5):1–4, September 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rosenblum:2012:LCb

[Ros12a] David S. Rosenblum. Letter from chair. ACM SIG-SOFT Software Engineering Notes, 37(4):1, July 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rosenblum:2012:LCa

[Ros12b] David S. Rosenblum. Letter from the chair. ACM SIG-SOFT Software Engineering Notes, 37(3):1, May 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Roy:2019:EIC

[Roy19] Subhajit Roy. Envisioning an intelligent collaborative integrated development environment. ACM SIGSOFT Software Engineering Notes, 44 (1):35, January 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rashid:2012:SAM

[RPB12] Ekbal Rashid, Srikanta Patnayak, and Vandana Bhattacherjee. A survey in the area of machine learning and its application for software quality prediction. ACM SIG-SOFT Software Engineering Notes, 37(5):1–7, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rech:2011:AEE

[RR11] Jörg Rech and Eric Ras. Aggregation of experiences in experience factories into software patterns. ACM SIGSOFT Software Engineering Notes, 36(2):1–4, March 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rao:2013:CPS

[RRK13] V. Bhujanga Rao, P. Seetha Ramaiah, and K. Raja Kumar. Clinical programming software to manage patient’s data with a cochlear implant. ACM SIGSOFT Software Engineering Notes, 38 (3):1–7, May 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rao:2013:OST

[RRN13] K. Koteswara Rao, GSVP Raju, and Srinivasan Nagaraj. Optimizing the software testing efficiency by using a genetic algorithm: a design methodology. ACM SIGSOFT Software Engineering Notes, 38(3):1–5, May 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raghunath:2013:DRB

[RRSV13] Amritha Raghunath, Sadhana Ramachandran, Selvakumar Subramanian, and Sreevatsan Vaidyanathan. Data rate based adaptive thread assignment solution for combating the SlowPOST denial of service attack. ACM SIGSOFT Software Engineering Notes, 38(5):1–5, September 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ratneshwer:2010:DAS

[RT10] Ratneshwer and A. K. Tripathi. Dependence analysis of software component. ACM SIGSOFT Software Engineering Notes, 35(4):1–9, July 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rai:2013:BIO

[RT13] Deepak Rai and Kirti Tyagi. Bio-inspired optimization techniques: a critical comparative study. ACM SIGSOFT Software Engineering Notes, 38(4):1–7, July 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ruparelia:2010:SDL

[Rup10] Nayan B. Ruparelia. Software development lifecycle models. ACM SIGSOFT Software Engineering Notes, 35(3):8–13, May 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Russo:2011:BRM

[Rus11] Aryldo G. Russo, Jr. Book review: Modeling in Event-B — system and software engineering by Jean-Raymond Abrial. ACM SIGSOFT Software Engineering Notes, 36 (2):38–39, March 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reddy:2012:PWS

[RVB12] K. Sudheer Reddy, G. Partha Saradhi Varma, and I. Ramesh Babu. Preprocessing the web server logs: an illustrative approach for effective usage mining. ACM SIGSOFT Software Engineering Notes, 37(3):1–5, May 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reddy:2012:DWU

[RVR12] K. Sudheer Reddy, G. Partha Saradhi Varma, and M. Kantha Reddy. Discovering web usage patterns by using an innovative practical approach. ACM SIG-SOFT Software Engineering Notes, 37(3):1–4, May 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sabharwal:2014:IIC

[SA14] Sangeeta Sabharwal and Manuj Aggarwal. Identifying interactions for combinatorial testing using data flow techniques. ACM SIGSOFT Software Engineering Notes, 39 (6):1–4, November 2014. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Soujanya:2016:GFC

[SA16] K. L. S. Soujanya and A. AnandaRao. A generic framework for configuration management of SPL and controlling evolution of complex software products. ACM SIG-SOFT Software Engineering Notes, 41(1):1–10, January 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sarkar:2017:HEI

[SA17] Santonu Sarkar and Gargi Alavani. How easy it is to write software for heterogeneous systems? ACM SIGSOFT Software Engineering Notes, 42(4):1–7, October 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saarimaki:2019:MIO

[Saa19] Nyyti Saarimäki. Methodological issues in observational studies. ACM SIG-SOFT Software Engineering Notes, 44(3):24, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3356773.3356799.

Safonov:2010:BRM

[Saf10] Vladimir O. Safonov. Book review: Microsoft Windows communication foundation: hands-on!, by Craig Mc-Murtry, Marc Mercuri and Nigel Watling, and published by Sams Publishing, 2006 (paperback), ISBN 0-672-32877-1, 539 pp., $39.99. ACM SIG-SOFT Software Engineering Notes, 35(3):40, May 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sutton:2019:SIC

[SAHC19] Stanley M. Sutton, Ove Armbrust, Regina Hebig, and Paul Clarke. Summary of the 2019 International Conference on Software and System Processes (ICSSP 2019). ACM SIGSOFT Software Engineering Notes, 44(3):34– 37, November 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3356773.3356803.

Sampaio:2013:BRQ

[Sam13a] Alberto Sampaio. Book review: Quantifying the user experience: practical statistics for user research by Jeff Sauro and James R. Lewis. ACM SIGSOFT Software Engineering Notes, 38(1):57–58, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saxena:2013:UML

[SAM13b] Vipin Saxena, Deepak Arora, and Nimesh Mishra. UML modeling of load optimization for distributed computer systems based on genetic algorithm. ACM SIGSOFT Software Engineering Notes, 38 (1):1–7, January 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Seth:2011:DSB

[SAS11] Ashish Seth, Himanshu Agarwal, and Ashim Raj Singla. Designing a SOA based model. ACM SIGSOFT Software Engineering Notes, 36(5):1– 7, September 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sharma:2016:IME

[SAS16] Richa Sharma, Peeyush Aggarwal, and Ashish Sureka. Insights from mining eleven years of scholarly paper publications in requirements engineering (RE) series of conferences. ACM SIGSOFT Software Engineering Notes, 41 (2):1–6, March 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2010:BRS

[Sau10]  Joe Saur.  Book review:

Simulation-Based Engineering of Complex Systems, Second Edition, by John R. Clymer, and published by John Wiley and Sons, 2009, hardcover, ISBN 978-0-470-40129-3, 503 pp. ACM SIGSOFT Software Engineering Notes, 35 (2):33–34, March 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2011:NSM

[Sau11]  Joe Saur.  Network sci-

ence for military coalition operations: information exchange and interaction by dinesh Verma. ACM SIGSOFT Software Engineering Notes, 36(6):30, November 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2013:BRD

[Sau13a] Joe Saur. Book review: Design for Emotion by Trevor van Gorp and Edie Adams. ACM SIGSOFT Software Engineering Notes, 38(4):40, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2013:TLP

[Sau13b] Joe Saur. Think like a programmer: an introduction to creative problem solving by v. Anton spraul. ACM SIGSOFT Software Engineering Notes, 38(1):58, January 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2012:ADC

[SBB12] Vinay Singh, Vandana Bhattacherjee, and Sandeep Bhattacharjee. An analysis of dependency of coupling on software defects. ACM SIGSOFT Software Engineering Notes, 37(1):1–6, January 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singhal:2013:CRV

[SBK13] Abhishek Singhal, Abhay Bansal, and Avadhesh Kumar. A critical review of various testing techniques in aspect-oriented software systems. ACM SIGSOFT Software Engineering Notes, 38 (4):1–9, July 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Siegmund:2019:SSE

[SBP19] Janet Siegmund, Andrew Begel, and Norman Peitek. Summary of the Sixth Edition of the International Workshop on Eye Movements in Programming. ACM SIG-SOFT Software Engineering Notes, 44(3):54–55, November 2019. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3356773.3356809.

Sangwan:2011:RBF

[SBS11] Om Prakash Sangwan, Pradeep Kumar Bhatia, and Yogesh Singh. Radial basis function neural network based approach to test oracle. ACM SIGSOFT Software Engineering Notes, 36(5):1–5, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2010:LSM

[Sch10] Robert Schaefer. The limits of systems-making organizations. ACM SIGSOFT Software Engineering Notes, 35 (6):1–20, November 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2011:LVP

[Sch11a] Robert Schaefer. On the limits of visual programming languages. ACM SIGSOFT Software Engineering Notes, 36 (2):7–8, March 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2011:SPH

[Sch11b] Robert Schaefer. Software perspectives on history, culture and efficiency. ACM SIGSOFT Software Engineering Notes, 36(6):6–7, November 2011. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2012:BRF

[Sch12a] Robert Schaefer. Book review: Fundamentals of multicore software development by Victor Pankratius, Ali-Reza Adi-Tabatabai, Walter Tichy. ACM SIGSOFT Software Engineering Notes, 37(3):37, May 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2012:LWD

[Sch12b]  Robert Schaefer.  Limita-

tions of web design. ACM SIGSOFT Software Engineering Notes, 37(6):7–8, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2012:HSJ

[Sch12c] Robert Schaefer. On having a second job as an adjunct. ACM SIGSOFT Software Engineering Notes, 37(4):9–10, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2012:SLA

[Sch12d] Robert Schaefer. The secret life of academic papers. ACM SIGSOFT Software Engineering Notes, 37(2):7–8, March 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2013:BRD

[Sch13a] Robert Schaefer. Book review: Dynamic reconfigurable network-on-chip design: innovations for computational processing and communication by Jih-Sheng Shen and Pao-Ann Hsuing. ACM SIG-SOFT Software Engineering Notes, 38(6):42–43, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2013:BRM

[Sch13b] Robert Schaefer. Book review: Managing the unmanageable: rules, tools, and insights for managing software people and teams by Micky W. Mantle and Ron Lichty. ACM SIGSOFT Software Engineering Notes, 38(3):34–35, May 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2013:BES

[Sch13c] Robert Schaefer. Businessefficient software development processes. ACM SIGSOFT Software Engineering Notes, 38(4):7–8, July 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2013:RLL

[Sch13d] Robert Schaefer. Relay ladder logic considered harmful. ACM SIGSOFT Software Engineering Notes, 38 (2):8–9, March 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2013:WLM

[Sch13e]  Robert Schaefer.  Why

Lehman matters: mediocracy. ACM SIGSOFT Software Engineering Notes, 38 (6):8–9, November 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2016:SSS

[Sch16a] Robert Schaefer. Science, society, and software engineering. Part 2: If science is outlawed, will only outlaws do science? ACM SIGSOFT Software Engineering Notes, 41 (5):7–8, September 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2016:WWE

[Sch16b] Robert Schaefer. The world will not end with a bang but with Http Error 404. ACM SIGSOFT Software Engineering Notes, 41(1):5–7, January 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2018:SSS

[Sch18] Robert Schaefer. Science, society, and software engineering: Part 3 — the catch. ACM SIGSOFT Software Engineering Notes, 43(2):4–6, April 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaefer:2019:WDW

[Sch19] Robert Schaefer. What do we mean when we talk about artificial intelligence?: (part 1). ACM SIGSOFT Software Engineering Notes, 44 (1):7–10, January 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sorkin:2011:LPG

[SD11] Arthur Sorkin and Peter Donovan. LR(1) parser generation system: LR(1) error recovery, oracles, and generic tokens. ACM SIGSOFT Software Engineering Notes, 36 (2):1–5, March 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saha:2019:IAS

[SEK^{+}19] Seemanta Saha, William Eiers, Ismet Burak Kadron, Lucas Bang, and Tevfik Bultan. Incremental attack synthesis. ACM SIGSOFT Software Engineering Notes, 44(4):16, December 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3364452.336445759.

Singh:2018:ERW

[SFTS18] Paramvir Singh, Sheikh Umar Farooq, Saurabh Tiwari, and Ashish Sureka. An experience report on Workshop on Emerging Software Engineering Education. ACM SIGSOFT Software Engineering Notes, 43(2):12–23, April 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sharma:2012:DRS

[SGM12] Neeraj Kumar Sharma, Vibha Gaur, and S. K. Muttoo. A dynamic reputation system with built-in attack resilience to safeguard buyers in e-market. ACM SIGSOFT Software Engineering Notes, 37(4):1–19, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sekar:2012:ASB

[SGS12a] Lakshmi Priya Sekar, Vinitha Reddy Gankidi, and Selvakumar Subramanian. Avoidance of security breach through selective permissions in Android operating system. ACM SIG-SOFT Software Engineering Notes, 37(5):1–9, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sharma:2012:SOO

[SGS12b] Meenakshi Sharma, Nasib S. Gill, and Sunil Sikka. Survey of object-oriented metrics: focusing on validation and formal specification. ACM SIGSOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shah:2016:ESD

[Sha16] Unnati S. Shah. An excursion to software development life cycle models: an old to ever-growing models. ACM SIGSOFT Software Engineering Notes, 41(1):1–6, January 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2013:MMQ

[Sin13]  Gagandeep Singh.  Met-

rics for measuring the quality of object-oriented software. ACM SIGSOFT Software Engineering Notes, 38 (5):1–5, September 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2019:UNL

[Sin19] Maninder Singh. Using natural language processing and graph mining to explore interrelated requirements in software artefacts. ACM SIG-SOFT Software Engineering Notes, 44(1):37, January 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sharma:2010:EES

[SK10a] Ashish Sharma and Dharmender Singh Kushwaha. Early estimation of software complexity using requirement engineering document. ACM SIGSOFT Software Engineering Notes, 35(5):1–7, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2010:AFF

[SK10b] Yogesh Singh and Pradeep Kumar. Application of feedforward neural networks for software reliability prediction. ACM SIGSOFT Software Engineering Notes, 35 (5):1–6, September 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2011:EEO

[SK11] Satwinder Singh and K. S. Kahlon. Effectiveness of encapsulation and objectoriented metrics to refactor code and identify error prone classes using bad smells. ACM SIGSOFT Software Engineering Notes, 36 (5):1–10, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sharma:2012:ARB

[SK12a] Ashish Sharma and D. S. Kushwaha. Applying requirement based complexity for the estimation of software development and testing effort. ACM SIGSOFT Software Engineering Notes, 37 (1):1–11, January 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shirole:2012:TCU

[SK12b] Mahesh Shirole and Rajeev Kumar. Testing for concurrency in UML diagrams. ACM SIGSOFT Software Engineering Notes, 37(5):1–8, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2012:PMD

[SK12c] Indu Singh and Manoj Kumar. A proposed model for data warehouse user behaviour using intrusion detection system. ACM SIGSOFT Software Engineering Notes, 37 (6):1–7, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2012:ERM

[SK12d] Satwinder Singh and K. S.

Kahlon. Effectiveness of refactoring metrics model to identify smelly and error prone classes in open source software. ACM SIGSOFT Software Engineering Notes, 37 (2):1–11, March 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shirole:2013:UBM

[SK13] Mahesh Shirole and Rajeev Kumar. UML behavioral model based test case generation: a survey. ACM SIG-SOFT Software Engineering Notes, 38(4):1–13, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saha:2018:ASS

[SKE^{+}18] Seemanta Saha, Ismet Burak Kadron, William Eiers, Lucas Bang, and Tevfik Bultan. Attack synthesis for strings using meta-heuristics. ACM SIGSOFT Software Engineering Notes, 43(4):56, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

S:2013:NSO

[SKJ^{+}13] Indu S., Srinivas N. K., Harish P. J., GangaPrasad R., Nobby Varghese, N. S. Sreekanth, and Supriya N. Pal. [NLP@Desktop]: a service oriented architecture for integrating NLP services in desktop clients. ACM SIGSOFT Software Engineering Notes, 38(4):1–4, July 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2010:TCP

[SKS10] Yogesh Singh, Arvinder Kaur, and Bharti Suri. Test case prioritization using ant colony optimization. ACM SIGSOFT Software Engineering Notes, 35(4):1–7, July 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2010:MRG

[SKT10] V. B. Singh, P. K. Kapur, and Abhishek Tandon. Measuring reliability growth of software by considering fault dependency, debugging time lag functions and irregular fluctuation. ACM SIGSOFT Software Engineering Notes, 35 (3):1–11, May 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Santos:2012:ICC

[SM12a] Bruno F. L. Santos and Hendrik T. Macedo. Improving CUDA^{T M} C/C++ encoding readability to foster parallel application development. ACM SIGSOFT Software Engineering Notes, 37 (1):1–5, January 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shafiei:2012:MCL

[SM12b] Nastaran Shafiei and Peter Mehlitz. Modeling class loaders in Java PathFinder version 7. ACM SIGSOFT Software Engineering Notes, 37 (6):1–5, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Staalhane:2016:ASA

[SM16] Tor Stålhane and Thor Myklebust. Agile safety analysis. ACM SIGSOFT Software Engineering Notes, 41 (2):27–29, March 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sahu:2017:CDS

[SM17] Madhusmita Sahu and Durga Prasad Mohapatra. Computing dynamic slices of feature– oriented programs using execution trace file. ACM SIG-SOFT Software Engineering Notes, 42(2):1–16, April 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Storey:2019:SDP

[SMP19] Kyle Storey, Eric Mercer, and Pavel Parizek. A sound dynamic partial order reduction engine for Java Pathfinder. ACM SIGSOFT Software Engineering Notes, 44(4):15, December 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3364452.3364457.

Steghofer:2019:SSS

[SNGM19] Jan-Philipp Steghöfer, Nan Niu, Jin L. C. Guo, and Anas Mahmoud. SST’19 — software and systems traceability:

Summary of the 10th International Workshop at the 41st International Conference on Software Engineering (ICSE), May 27, 2019. ACM SIG-SOFT Software Engineering Notes, 44(3):43–47, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3356773.3356806.

Satish:2017:TPA

[SNR17] Preeti Satish, Peri Nikhil, and Krishnan Rangarajan. A test prioritization algorithm that cares for “Don’t Care” values and higher order combinatorial coverage. ACM SIGSOFT Software Engineering Notes, 42(4):1–9, October 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sagar:2010:SCBa

[SNS10a] Shrddha Sagar, N. W. Nerurkar, and Arun Sharma. A soft computing based approach to estimate reusability of software components. ACM SIGSOFT Software Engineering Notes, 35(4):1–4, July 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sagar:2010:SCBb

[SNS10b] Shrddha Sagar, N. W. Nerurkar, and Arun Sharma. A soft computing based approach to estimate reusability of software components. ACM SIG-SOFT Software Engineering Notes, 35(5):1–5, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sengupta:2010:EME

[SNS10c] Abhik Sengupta, Vivek Nandey, and Sabnam Sengupta. ETD-SOA: a model for event and time driven service oriented architecture. ACM SIGSOFT Software Engineering Notes, 35(6):1–9, November 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Solomon:2012:NML

[Sol12] A. D. Solomon. On a new masters level program in software engineering. ACM SIG-SOFT Software Engineering Notes, 37(5):1–3, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Soldani:2019:GLS

[Sol19] Jacopo Soldani. Grey literature: a safe bridge between academy and industry? ACM SIGSOFT Software Engineering Notes, 44 (3):11–12, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3356773.3356776.

Shareef:2013:CCA

[SP13] Jawwad Wasat Shareef and R. K. Pandey. CAME: Component Assembly Metrics Extraction using UML. ACM SIGSOFT Software Engineering Notes, 38(4):1–12, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2016:GBD

[SPKM16] Jagannath Singh, Subhrakanta Panda, P. M. Khilar, and D. P. Mohapatra. A graphbased dynamic slicing of distributed aspect-oriented software. ACM SIGSOFT Software Engineering Notes, 41 (2):1–8, March 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sharma:2012:SLT

[SRS12] Vineet Sharma, S. A. M.

Rizvi, and Arun Sharma. Software licenses — a tool to control distribution of software. ACM SIGSOFT Software Engineering Notes, 37 (6):1–4, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sharma:2013:SLT

[SRS13] Vineet Sharma, S. A. M.

Rizvi, and Arun Sharma. Software licenses — a tool to control distribution of software. ACM SIGSOFT Software Engineering Notes, 38 (1):49–51, January 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Suri:2010:AGD

[SS10a] P. K. Suri and Gurdev Singh. Automatic generation of design search keywords from software specifications to improve design search results. ACM SIGSOFT Software Engineering Notes, 35 (6):1–8, November 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Suri:2010:DMF

[SS10b] P. K. Suri and Gurdev Singh. DG-metrics formulization for DGML-based software design. ACM SIGSOFT Software Engineering Notes, 35(3):1–8, May 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Suri:2011:ATC

[SS11] Bharti Suri and Shweta Singhal. Analyzing test case selection and prioritization using ACO. ACM SIGSOFT Software Engineering Notes, 36 (6):1–5, November 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2013:ESE

[SS13] Gagandeep Singh and Hardeep Singh. Effect of software evolution on metrics and applicability of Lehman’s laws of software evolution. ACM SIGSOFT Software Engineering Notes, 38(1):1–7, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sharma:2016:NYS

[SS16] Richa Sharma and Ashish Sureka. A nine year story of the India software engineering conference from 2008 to 2016. ACM SIGSOFT Software Engineering Notes, 41(5):31–44, September 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shukla:2012:RSE

[SSJM12] Rakesh Shukla, Ashish Sureka, Rushikesh Joshi, and Rajib Mall. A report on Software Engineering Education Workshop. ACM SIGSOFT Software Engineering Notes, 37 (3):26–31, May 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2013:TGC

[SSK13] Pavitdeep Singh, Satwinder Singh, and Jatinder Kaur. Tool for generating code metrics for C# source code using abstract syntax tree technique. ACM SIGSOFT Software Engineering Notes, 38 (5):1–6, September 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

StPierre:2012:BRU

[St.12] Richard St. Pierre. Book review: The UX book, process and guidelines for ensuring a quality user experience by Rex Hartson and Pardha S. Pyla. ACM SIG-SOFT Software Engineering Notes, 37(5):43–44, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stobie:2013:BRX

[Sto13] Keith Stobie. Book review: OS X Mountain Lion Pocket Guide by Chris Seibold. ACM SIGSOFT Software Engineering Notes, 38(1):59, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sun:2018:RAR

[Sun18]  Yu-Shan Sun.  Reasoning

about reference behavior with RESOLVE. ACM SIGSOFT Software Engineering Notes, 43(3):18–19, July 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sinhal:2013:NFB

[SV13] Amit Sinhal and Bhupendra Verma. A novel fuzzy based approach for effort estimation in software development. ACM SIGSOFT Software Engineering Notes, 38 (5):1–6, September 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:2013:RPT

[SVT13] Lalit Kumar Singh, Gopika Vinod, and A. K. Tripathi. Reliability prediction through system modeling. ACM SIG-SOFT Software Engineering Notes, 38(6):1–10, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sitaraman:2018:STF

[SW18] Murali Sitaraman and Bruce W.

Weide. A synopsis of twenty five years of RESOLVE PhD research efforts. ACM SIG-SOFT Software Engineering Notes, 43(3):17, July 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Swamy:2012:BRSa

[Swa12a] Harisankar Krishna Swamy.

Book review: Space based technologies and commercialized development: economic implications and benefits by Stella Tkatchova. ACM SIGSOFT Software Engineering Notes, 37(3):37–38, May 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Swamy:2012:BRSb

[Swa12b] Harisankar Krishna Swamy.

Book review: Structured parallel programming patterns for efficient computation by Michael McCool, Arch D. Robison and James Reinders. ACM SIGSOFT Software Engineering Notes, 37 (6):43, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sharma:2017:VCS

[SWMV17] Vaibhav Sharma, Michael W. Whalen, Stephen McCamant, and Willem Visser. Veritesting challenges in symbolic execution of Java. ACM SIGSOFT Software Engineering Notes, 42(4):1–5, October 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Taibi:2013:ROS

[Tai13] Fathi Taibi. Reusability of open-source program code: a conceptual model and empirical investigation. ACM SIG-SOFT Software Engineering Notes, 38(4):1–5, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tan:2012:BRH

[Tan12] Kwee Heong Tan. Book review: How Google tests software by James A. Whittaker, Jason Arbon and Jeff Carollo. ACM SIGSOFT Software Engineering Notes, 37(5):44– 45, September 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tekinerdogan:2012:AGS

[TCB^{+}12] Bedir Tekinerdogan, Semih Cetin, Muhammad Ali Babar, Patricia Lago, and Juho Mäkiö. Architecting in global software engineering. ACM SIGSOFT Software Engineering Notes, 37(1):1–7, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Theisen:2017:SER

[TDWV17] Christopher Theisen, Marcel Dunaiski, Laurie Williams, and Willem Visser. Software engineering research at the International Conference on Software Engineering in 2016. ACM SIGSOFT Software Engineering Notes, 42 (4):1–7, October 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tee:2010:MCC

[Tee10a] Sim-Hui Tee. Measuring class cohesion using mutant methods. ACM SIGSOFT Software Engineering Notes, 35 (6):1–4, November 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tee:2010:MEM

[Tee10b] Sim-Hui Tee. Method efficiency model based on value relevancy. ACM SIGSOFT Software Engineering Notes, 35(4):1–3, July 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tee:2011:ISS

[Tee11]  Sim-Hui Tee.  Identifying

structural similarity of methods using isomorphic graphs. ACM SIGSOFT Software Engineering Notes, 36(6):1– 3, November 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teixeira:2018:IQC

[Tei18] Eudis Teixeira. Improving the quality of controlled experiments in software engineering. ACM SIGSOFT Software Engineering Notes, 43 (1):1–6, January 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tekinerdogan:2012:FTS

[Tek12a] Bedir Tekinerdogan. First Turkish Software Product Line Engineering Workshop summary. ACM SIGSOFT Software Engineering Notes, 37(6):30–34, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tekinerdogan:2012:RTA

[Tek12b]  Bedir Tekinerdogan.  Re-

flection on Turkish aspectoriented software development workshop series. ACM SIG-SOFT Software Engineering Notes, 37(2):30–33, March 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2011:BRI

[Teo11] Vasile G. Teodorovici. Book review: Integrating and extending BIRT 3rd edition by J. Weathersby, T. Bondurand, and I. Chatalbasheva. ACM SIGSOFT Software Engineering Notes, 36(6):31, November 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2012:BRC

[Teo12a] Vasile G. Teodorovici. Book review: Continuous testing with Ruby, Rails and JavaScript by Ben Rady and Rod Coffin. ACM SIGSOFT Software Engineering Notes, 37(1):36, January 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2012:BRE

[Teo12b] Vasile G. Teodorovici. Book review: The Eclipse graphical editing framework (GEF) by D. Rubel, J. Wren, and E. Clayberg. ACM SIGSOFT Software Engineering Notes, 37(1):35, January 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2012:BRT

[Teo12c] Vasile G. Teodorovici. Book review: Theory of conditional games by Wynn C. Stirling. ACM SIGSOFT Software Engineering Notes, 37(4):41, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2012:BRW

[Teo12d] Vasile G. Teodorovici. Book review: Work item management with IBM rational Clear-Quest and Jazz: a customization guide by Shmuel Bashan and David E. Bellagio. ACM SIGSOFT Software Engineering Notes, 37(2):37–38, March 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2013:BRA

[Teo13a] Vasile G. Teodorovici. Book review: Advanced programming in the UNIX environment, third edition by W. Richard Stevens and Stephen A. Rago. ACM SIGSOFT Software Engineering Notes, 38(6):45, November 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2013:BRD

[Teo13b] Vasile G. Teodorovici. Book review: Drupal For Designers by Dani Nordin. ACM SIGSOFT Software Engineering Notes, 38(5):68, September 2013. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2013:BRJ

[Teo13c] Vasile G. Teodorovici. Book review: jQuery, jQuery UI and jQuery Mobile: recipes and examples by Adriaan de Jonge and Phil Dutson. ACM SIGSOFT Software Engineering Notes, 38(5):68– 69, September 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2013:BRL

[Teo13d] Vasile G. Teodorovici. Book review: Learning JavaScript: a hands-on guide to the fundamentals of modern JavaScript by Tim Wright. ACM SIG-SOFT Software Engineering Notes, 38(3):35–36, May 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2013:BRM

[Teo13e] Vasile G. Teodorovici. Book review: Modern embedded computing: designing connected, pervasive, media-rich systems by Peter Barry and Patrick Crowley. ACM SIG-SOFT Software Engineering Notes, 38(1):59–60, January 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2013:BRS

[Teo13f] Vasile G. Teodorovici. Book review: Security and privacy for Microsoft Office 2010 users by Mitch Tulloch. ACM SIG-SOFT Software Engineering Notes, 38(4):40–41, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2013:BRO

[Teo13g] Vasile G. Teodorovici. Book review: The object-oriented thought process, fourth edition by Matt Weisfeld. ACM SIG-SOFT Software Engineering Notes, 38(6):44–45, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2013:BRU

[Teo13h] Vasile G. Teodorovici. Book review: Ubuntu unleashed, 2013 edition by Matthew Helmke with Andrew Hudson and Paul Hudson. ACM SIGSOFT Software Engineering Notes, 38(4):41, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2013:XML

[Teo13i] Vasile G. Teodorovici. OS x mountain lion: the missing manual by David pogue. ACM SIGSOFT Software Engineering Notes, 38(6):44, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Teodorovici:2013:TTD

[Teo13j] Vasile G. Teodorovici. Tcl/ Tk: a developer’s guide, third edition by clif flynt. ACM SIGSOFT Software Engineering Notes, 38(3):36, May 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Thakur:2011:DRB

[TG11] Garima Thakur and Anjana Gosain. DWEVOLVE: a requirement based framework for data warehouse evolution. ACM SIGSOFT Software Engineering Notes, 36 (6):1–8, November 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tiwari:2013:RRT

[TG13] Rajeev Tiwari and Noopur Goel. Reuse: reducing test effort. ACM SIGSOFT Software Engineering Notes, 38 (2):1–11, March 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tiwari:2012:MCA

[TJ12] Manisha Tiwari and Padmaja Joshi. Method cohesion analysis through concept lattices. ACM SIGSOFT Software Engineering Notes, 37(3):1–4, May 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Torre:2016:IWU

[TLG^{+}16] Damiano Torre, Yvan Labiche, Marcela Genero, Maged Elaasar, Tuhin Kanti Das, Bernhard Hoisl, and Matthias Kowal. 1st International Workshop on UML Consistency Rules (WU-COR 2015): Post workshop report. ACM SIGSOFT Software Engineering Notes, 41 (2):34–37, March 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Terra:2013:QCC

[TMVB13] Ricardo Terra, Luis Fernando Miranda, Marco Tulio Valente, and Roberto S. Bigonha. Qualitas.class Corpus: a compiled version of the Qualitas Corpus. ACM SIGSOFT Software Engineering Notes, 38 (5):1–4, September 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tonelli:2019:WSI

[TODM19] Roberto Tonelli, Marco Ortu, Stephane Ducasse, and Michele Marchesi. Workshop summary: 2019 IEEE /ACM Second International Workshop on Emerging Trends in Software Engineering for Blockchain (WETSEB 2019). ACM SIGSOFT Software Engineering Notes, 44(3):48– 52, November 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3356773.3356807.

Tracz:2010:BRM

[Tra10a] Will Tracz. Book review:

Making it big in software: get the job, work the org, become great by Sam Lightstone. ACM SIGSOFT Software Engineering Notes, 35(5):58, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2010:BRR

[Tra10b] Will Tracz. Book review:

Remarkable Engineers: From Riquet to Shannon by Ioan James. ACM SIGSOFT Software Engineering Notes, 35 (5):58, October 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL http://

dl.acm.org/citation.cfm?
doid=1838687.1862454.

Tracz:2010:BRD

[Tra10c] Will Tracz. Book review: The design of design: essays of a computer scientist by Frederick R. Brooks, Jr. ACM SIGSOFT Software Engineering Notes, 35(5):58, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2011:BRG

[Tra11a] Will Tracz. Book review:

Glitch: the hidden impact of faulty software by Jeff Papows. ACM SIGSOFT Software Engineering Notes, 36 (2):41, March 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2011:LFE

[Tra11b] Will Tracz. Lord of the files: essays on the social aspects of software engineering by russel ovans. ACM SIGSOFT Software Engineering Notes, 36 (6):31, November 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2012:BHD

[Tra12] Will Tracz. A bug hunter’s diary: a guided tour through the wilds of software security by tobias Klein. ACM SIGSOFT Software Engineering Notes, 37(1):37–38, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tracz:2013:SEC

[Tra13] Will Tracz. Software engineering: the current practice by vaclav rajlich. ACM SIGSOFT Software Engineering Notes, 38(6):46, November 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tribbey:2010:BRF

[Tri10a]  Will Tribbey.  Book re-

view: F# for Scientists, by Jon Harrop, and published by Wiley-Interscience, 2008, hardback, ISBN 0-470-24211-6, 368pp. ACM SIGSOFT Software Engineering Notes, 35(2):34–35, March 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tribbey:2010:BRN

[Tri10b] Will Tribbey. Book review:

Numerical Recipes: The Art of Scientific Computing (3rd Edition) by William H. Press, Saul A. Teukolsky, William T. Vetterling, and Brian P. Flannery, and published by Cambridge University Press, 2007, hardback, ISBN 978-0-521-88068-8, 1235 pp. ACM SIG-SOFT Software Engineering Notes, 35(6):30–31, November 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tyagi:2011:RCB

[TS11] Kirti Tyagi and Arun Sharma.

Reliability of component based systems: a critical survey. ACM SIGSOFT Software Engineering Notes, 36 (6):1–6, November 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Treude:2010:WRW

[TSEvD10] Christoph Treude, Margaret-Anne Storey, Kate Ehrlich, and Arie van Deursen. Workshop report from Web2SE: First Workshop on Web 2.0 for Software Engineering. ACM SIGSOFT Software Engineering Notes, 35(5):45– 50, September 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Treude:2011:WRW

[TSvD^{+}11] Christoph Treude, Margaret-Anne Storey, Arie van Deursen, Andrew Begel, and Sue Black. Workshop report from Web2SE 2011: 2nd International Workshop on Web 2.0 for Software Engineering. ACM SIGSOFT Software Engineering Notes, 36(5):24– 29, September 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Turlea:2019:MLT

[Tur19] Ana Turlea. Model-in-the-loop testing for cyber physical systems. ACM SIGSOFT Software Engineering Notes, 44 (1):37, January 2019. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Upadhyay:2010:DMI

[UDA10] Nitin Upadhyay, Bharat M. Deshapande, and Vishnu P. Agrawal. Developing maintainability index of a software component: a digraph and matrix approach. ACM SIG-SOFT Software Engineering Notes, 35(5):1–11, September 2010. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Unterkalmsteiner:2019:SIW

[UYG^{+}19]  Michael  Unterkalmsteiner,

Tingting Yu, Gregory Gay, Elizabeth Bjarnason, Markus Borg, and Michael Felderer. Summary of the 5th International Workshop on Requirements Engineering and Testing (RET 2018). ACM SIGSOFT Software Engineering Notes, 44(1):31–34, January 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vieira:2017:CPO

[VA17] Ianegitz Vieira and Alexandre Alvaro. A centralized platform of open government data as support to applications in the smart cities context. ACM SIGSOFT Software Engineering Notes, 42(4):1–13, October 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Voola:2013:CRP

[VB13] Persis Voola and A. Vinaya Babu. Comparison of requirements prioritization techniques employing different scales of measurement. ACM SIGSOFT Software Engineering Notes, 38(4):1–10, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Varona:2012:ESE

[VCPR12] Daniel Varona, Luiz Fernando Capretz, Yadenis Piñero, and Arif Raza. Evolution of software engineers’ personality profile. ACM SIGSOFT Software Engineering Notes, 37 (1):1–5, January 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

vanderLinden:2018:ESWa

[vdLR18a] Dirk van der Linden and Awais Rashid. The effect of software warranties on cybersecurity. ACM SIGSOFT Software Engineering Notes, 43 (4):31–35, October 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

vanderLinden:2018:ESWb

[vdLR18b] Dirk van der Linden and Awais Rashid. The effect of software warranties on cybersecurity. ACM SIGSOFT Software Engineering Notes, 43 (4):53, October 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

vanderMerwe:2012:VAA

[vdMvdMV12] Heila van der Merwe, Brink van der Merwe, and Willem Visser. Verifying Android applications using Java PathFinder. ACM SIGSOFT Software Engineering Notes, 37(6):1– 5, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Varshney:2013:SBS

[VM13] Sapna Varshney and Monica Mehrotra. Search based software test data generation for structural testing: a perspective. ACM SIGSOFT Software Engineering Notes, 38(4):1–6, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

V:2011:BBI

[VS11a] Sharath Chandra V. and S. Selvakumar. BIXSAN: browser independent XSS sanitizer for prevention of XSS attacks. ACM SIGSOFT Software Engineering Notes, 36 (5):1–7, September 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vembuselvi:2011:LLL

[VS11b] C. Vembuselvi and S. Selvakumar. LISISAP: link level signature based secure anonymous protocol for prevention of traffic analysis attacks. ACM SIGSOFT Software Engineering Notes, 36 (2):1–10, March 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Vu:2011:BRA

[Vu11] Jodat Vu. Book review: The art of multiprocessor programming by Maurice Herlihy and Nir Shavit. ACM SIGSOFT Software Engineering Notes, 36(5):52–53, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:2018:HCW

[Wan18] Huaimin Wang. Harnessing the crowd wisdom for software trustworthiness. ACM SIGSOFT Software Engineering Notes, 43(1):1–6, January 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wright:2010:ISD

[WC10] M. Keith Wright and Charles J. Capps III. Information systems development project performance in the 21st century. ACM SIGSOFT Software Engineering Notes, 35 (2):1–10, March 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:2018:PBJ

[WCG^{+}18] Kaiyuan Wang, Hayes Converse, Milos Gligoric, Sasa Misailovic, and Sarfraz Khurshid. A progress bar for the JPF search using program executions. ACM SIGSOFT Software Engineering Notes, 43 (4):55, October 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weide:2018:RCD

[Wei18] Alan Weide. Reasoning challenges of data abstraction and aliasing in concurrent programs. ACM SIGSOFT Software Engineering Notes, 43 (3):18, July 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Welch:2018:FID

[Wel18] Daniel Welch. Formalization integrated development environments: The current landscape. ACM SIGSOFT Software Engineering Notes, 43(3):17, July 2018. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Werden:2010:BRT

[Wer10] David A. Werden. Book review: TCP/IP architecture, design, and implementation in Linux by Sameer Seth and M. Ajaykumar Venkatesulu. ACM SIGSOFT Software Engineering Notes, 35 (5):57, September 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

White:2011:BRR

[Whi11] Randall L. White. Book review: Reconfigurable embedded control systems: applications for flexibility and agility by Mohamed Khalgui and Han-Michale Hanisch. ACM SIGSOFT Software Engineering Notes, 36(5):53, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2010:BDM

[Win10a] Mike Wing. Burt doesn’t manage. ACM SIGSOFT Software Engineering Notes, 35(4):4–6, July 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2010:M

[Win10b] Mike Wing. The making of. ACM SIGSOFT Software Engineering Notes, 35 (6):5–7, November 2010. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2010:PC

[Win10c] Mike Wing. Poetry in code. ACM SIGSOFT Software Engineering Notes, 35(3):7–8, May 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2011:ESL

[Win11a] Michael Wing. Everything in SE is a lie. ACM SIG-SOFT Software Engineering Notes, 36(5):8–9, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2011:HWT

[Win11b] Michael Wing. How we talk about basics. ACM SIGSOFT Software Engineering Notes, 36(3):4–6, May 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2012:FT

[Win12a]  Michael Wing.  Food for

thought. ACM SIGSOFT Software Engineering Notes, 37(1):9–10, January 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wing:2012:TY

[Win12b] Mike Wing. Thank you. ACM SIGSOFT Software Engineering Notes, 37(3):7–8, May 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wu:2012:MMS

[WJ12] Ye-Chi Wu and Hewijin Christine Jiau. A monitoring mechanism to support agility in service-based application evolution. ACM SIGSOFT Software Engineering Notes, 37(5):1–10, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:2017:JRJ

[WKG17] Kaiyuan Wang, Sarfraz Khurshid, and Milos Gligoric. JPR: Replaying JPF traces using standard JVM. ACM SIGSOFT Software Engineering Notes, 42(4):1–5, October 2017. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:2013:RFA

[WL13] Qianxiang Wang and Lin Liu. Report on the Fourth Asia-Pacific Symposium on Internetware (Internetware 2012). ACM SIGSOFT Software Engineering Notes, 38(4):25–26, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:2012:BPS

[WZ12] Wenyuan Wang and Zheng Zhang. Balanced partition scheme for distributed caching systems to solve load imbalance problems. ACM SIG-SOFT Software Engineering Notes, 37(6):1–6, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xie:2016:OTO

[Xie16a] Tao Xie. Outward thinking for our research community. ACM SIGSOFT Software Engineering Notes, 41(4):7–8, July 2016. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xie:2016:PPI

[Xie16b] Tao Xie. The pursuit of practice-impactful research. ACM SIGSOFT Software Engineering Notes, 41(1): 7–8, January 2016. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xie:2019:WSI

[XPP19] Xiaoyuan Xie, Pak-Lok Poon, and Laura L. Pullum. Workshop summary: 2019 IEEE /ACM Fourth International Workshop on Metamorphic Testing (MET 2019). ACM SIGSOFT Software Engineering Notes, 44(3):56–59, November 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3356773.3356810.

Xie:2013:RIS

[XZM13] Tao Xie, Lu Zhang, and Hong Mei. Report on the International Symposium on High Confidence Software (ISHCS 2011/2012). ACM SIGSOFT Software Engineering Notes, 38(4):27–33, July 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yatake:2012:SBE

[YA12] Kenro Yatake and Toshiaki Aoki. SMT-based enumeration of object graphs from UML class diagrams. ACM SIGSOFT Software Engineering Notes, 37(4):1–8, July 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yaman:2018:UIC

[Yam18] Sezin Gizem Yaman. User involvement in continuous experimentation. ACM SIG-SOFT Software Engineering Notes, 43(1):1–4, January 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yadav:2011:FSM

[YAS11] Seema Yadav, Khaleel Ahamd, and Jayant Shekhar. Finite state machine based approach to prevent format string attacks. ACM SIGSOFT Software Engineering Notes, 36 (5):1–7, September 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yazbek:2010:CQA

[Yaz10] Hashem Yazbek. A concept of quality assurance for metrics in CASE-tools. ACM SIGSOFT Software Engineering Notes, 35(5):1–8, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yamagata:2012:FSE

[YKF^{+}12] Yoriyuki Yamagata, Weiqiang Kong, Akira Fukuda, Van Tang Nguyen, Hitoshi Ohsaki, and Kenji Taguchi. Formal semantics of extended hierarchical state transition matrix by CSP. ACM SIGSOFT Software Engineering Notes, 37(4):1–8, July 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yilmaz:2011:SPE

[YO11] Murat Yilmaz and Rory V.

O’Connor. A software process engineering approach to improving software team productivity using socioeconomic mechanism design. ACM SIGSOFT Software Engineering Notes, 36(5):1–5, September 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yu:2011:CIE

[Yu11] Liguo Yu. Coevolution of information ecosystems: a study of the statistical relations among the growth rates of hardware, system software, and application software. ACM SIGSOFT Software Engineering Notes, 36 (6):1–5, November 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zage:2013:SSE

[Zag13] Wayne Zage. The security and software engineering research center (S2ERC). ACM SIGSOFT Software Engineering Notes, 38(2):6–7, March 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zhang:2012:SSS

[ZCW12] Yufeng Zhang, Zhenbang Chen, and Ji Wang. S2PF: speculative symbolic PathFinder. ACM SIGSOFT Software Engineering Notes, 37(6):1– 5, November 2012. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zheng:2018:ADS

[ZLNP18] Guolong Zheng, Quang Loc Le, ThanhVu Nguyen, and Quoc-Sang Phan. Automatic data structure repair using separation logic. ACM SIG-SOFT Software Engineering Notes, 43(4):66, October 2018. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zaidi:2013:MVE

[ZS13] Taskeen Zaidi and Vipin Saxena. Modeling and validation of execution of tasks on high speed network under distributed environment. ACM SIGSOFT Software Engineering Notes, 38(3):1–6, May 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Zaidi:2014:PES

[ZS14] Taskeen Zaidi and Vipin Saxena. Performance estimation of static step topology across distributed networks through simulation tool. ACM SIG-SOFT Software Engineering Notes, 39(6):1–4, November 2014. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).